A Complete Bibliography of Publications in International Aquatic Research

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

13 April 2024 Version 1.06

Title word cross-reference

+ [56]. ^{2+} [43]. ^{241} [61]. _{1} [252]. _{3} [350]. α [237]. β [308]. n − 6 [49]. × [246, 219].

-ATPase [56]. -glucan [308]. -Melanocyte-stimulating [237].

/K [56].

1 [214, 364]. 16S [189]. 1D1 [268].

3-related [96].

4-nonylphenol [235].

70 [57].

9-cis-carotene [369].

abalone [239, 229]. Abramis [205]. Abundance [342, 106, 245]. according [30]. accrual [54]. accumulation [369, 186]. accuracy [161]. acetylcholinesterase [81]. Acheta [223]. acid [323, 211, 82, 218, 248, 186]. acidic [222]. acids [99, 6, 24, 49, 209]. Acipenser [9, 94, 4, 40, 43, 45, 20, 15, 185, 238, 12, 5, 42, 144, 313, 161, 34, 46]. across [326, 261]. actinobacteria [189]. activated [298]. activation [322, 72]. active [290]. activities [309, 94, 36, 286, 218, 115, 120, 266, 337]. activity [253, 264, 296, 281, 97, 56, 343, 81, 260, 86, 347, 102, 79]. aculeatus [296, 18, 327]. aculeatus [?]SchistocephalusTarallo:2021:TER. Acute [312, 143, 63, 304]. Adams [243]. additional [341]. additive [204]. additives [213]. address [360]. Adenosine [186]. adherent [228]. Adriatic [106]. adult [96]. Aegean [53]. aeration [259]. aerobic [261]. Aeromonas [314, 226, 110, 339, 16, 173]. aff [333]. affect [329, 263]. affected [168, 242, 129, 74, 153]. affects [286]. affinis [195]. aflatoxin [252]. African [230, 117, 251, 125, 278, 92, 64]. after [63, 303]. against [253, 90, 339, 308, 306, 347, 332, 355]. agar [357]. Age [10, 86, 138]. Age-related [86]. aggressive [345]. agricultural [196, 117, 233]. air [126, 19, 321]. air-breathing [321]. Aitoliko [26]. akajei [267]. Alaskan [307]. albacares [29, 7, 17]. Alcalase [2, 17]. alga [198, 343]. algae [283, 27, 130, 140, 158]. Algal [54]. alginate [298]. alginate-immobilized [298]. alien [18]. alive [82]. Allium [110]. Aloe [251]. along [150, 54, 112]. alpinus [101]. Alteration [214]. Alterations [81]. alternate [284]. Altiplano [159]. altitude [159]. Am [61]. Amazon [190, 119]. amazonicus [312]. America [157]. American [51]. amino [211, 218]. Ammonia [265, 146]. ammonium [283]. among [127]. Amphiprion [261]. Amur [213, 46]. amyloliquefaciens [89]. Anacanthorus [119]. Anadara [243, 112]. anaesthetised [335]. analysis [152, 370, 162, 28, 103, 336, 165, 52]. Anambra [35]. anatomic [4]. anchovy [297]. androgen [325]. androgenesis [70]. anemone [253]. anesthetic [269, 247]. anesthetics [301]. angel [6]. Anguilla [51]. anguillarum [166]. animal [180]. animals [142]. Ankistrodesmus [256]. Annual [205, 315, 66]. anomalies [134]. Antalya [211]. Antibacterial [177, 253, 264, 281, 343, 347]. antibiofilm [253]. antibiotic [253]. antibiotic-resistant [253]. antibody [332]. antimicrobial [228, 224, 120, 260]. antioxidant [257, 290, 371, 241, 205, 343, 74, 337]. Antioxidative [129, 79, 120, 102]. aponinum [158]. apoptosis [126]. apoptosis-related [126]. Apostichopus [236]. apple [257]. application [1]. applications [295]. aqua [353]. Aquaculture [336, 367, 345, 311, 1, 334, 275, 287, 107, 47, 289, 318, 169, 261, 112]. aquatic [176, 360, 276]. aqueous [217, 177, 298]. Arabia [150]. Arabian [156, 162]. arabinogalactan [286]. arable [140]. Arctic [101]. area [157, 297, 114, 243]. areas [150, 25, 30]. Argentina [157]. Ariake [243]. arid [169, 140]. arjuna [306]. arNOX [86]. arsenic [266, 321]. artemia [6, 362]. Artificial [59, 197, 238]. Asia [336]. Asian [105, 118, 234]. asotus [213]. aspects [232]. assays [164]. assemblages [242]. assess [176]. assessed [101]. Assessing [242, 28]. Assessment [334, 25, 161, 236, 326, 372, 58, 270, 123, 143]. associated [40, 361, 100, 260, 221]. Astacus [285, 182]. Astaxanthin [153]. Atlantic [83, 126, 297, 181, 42]. Atmospheric [145]. atomic [360]. ATPase [36, 56]. Atrazine [19]. Atrina [243, 215]. atrogenes [214]. Attempts [238]. aurata [174, 106, 24, 49, 175]. auratus [39, 368, 206]. Aurelia [131]. Australian [152]. australis [91]. autochthonous [339]. autolysis [36, 129]. availability [54]. available [178]. Avicennia [273, 272]. axis [201]. axoneme [72]. Azolla [298]. azurea [221].

B [252]. B. [89]. bacteria [228, 334, 253, 326, 107, 7]. bacteria-inoculated [334]. Bacterial [64, 188, 209, 236]. bacterially [286]. badionotus [333]. baeri [4]. baerii [185, 313, 161, 34]. Bandon [165]. Bangladesh [326]. Barbus [108]. barracuda [78]. barramundi [210, 363]. based [213, 312, 323, 178, 240, 293, 48, 185, 313, 278, 189]. Baseline [206]. basilicum [335]. basin [159, 190]. Basque [191, 212]. bass [174, 219, 75, 24, 49, 262, 234]. bathing [45]. bathymetric [53]. batrachus [103]. Bay [191, 212, 165, 143]. be [227]. beach [341]. beach-stranded [341]. beaches [341]. before [303]. behaviour [105]. behavioural [164]. belcheri [163]. Belostomatidae [195]. beluga [21, 82, 41, 2, 207]. benthic [227, 108]. benzo [268]. benzo-a-pyrene [268]. benzocaine [301]. betaine [164]. Betta [316, 300]. between [174, 25, 131, 276, 195]. bicarbonate [269]. Bidyanus [192]. binding [76]. bio [116]. bio-indicator [116]. bioaccumulation [168, 14]. bioactive [253, 260]. Bioactivity [217]. Biochemical [68, 77, 168, 304, 358, 251, 244, 371, 313, 349, 116]. biochemistry [309, 340]. Bioconversion [284]. biofilm [228, 337]. biofloc [149, 250, 275, 184, 348, 351]. biofuel [346]. biogrouts [169]. biological [178]. biomarker [303]. biomarkers [176, 258]. biomass [141, 280, 26, 199, 54, 204, 284]. biosensor [279]. Biosorption [14, 298]. biota [152]. Biscay [191, 212]. bivalve [370]. bivalves [243, 128]. Bivalvia [215]. black [326, 223, 55, 218]. bladder [249]. Bleeker [211, 202]. Bloch [105, 19, 33, 234]. blood [233, 358, 15, 185, 161, 112]. bloom [303, 132]. bluefin [187]. bodies [287]. body [40, 250, 82, 130, 172, 218, 313, 234]. Boone [348]. boost [346]. booster [22]. born [77, 168]. Borodin [94]. bovine [62]. brachiosoma [93]. brackish [21]. brain [81]. brains [24]. brama [205]. branchial [56]. Brandt [4, 217, 185, 34]. Bratislava [114]. Brazil [85, 176, 341, 303]. Brazilian [190, 302, 288]. breakdown [15]. bream [174, 62, 24, 49, 205, 175]. breathing [19, 321]. breeding [197]. brief [128]. broodstock [175]. brown [323, 120]. bruneus [325]. Brycon [312]. budget [117]. Burchell [117, 92]. Burman [89]. by-product [83, 7].

C [6, 143]. cadmium [33]. caged [243]. cages [178, 118, 331]. calcarifer [363, 210, 234]. calcitonin [237, 267, 368]. calcium [344, 237, 368, 317]. calmodulin [187]. Can [341, 329, 227]. canadensis [61]. Candida [217]. candidate [103]. canola [83]. Capability [196]. capacity [371, 328]. Capitella [143]. captivity [13]. Capture [366]. captured [160]. Carassius [368, 206]. carbohydrate [179, 98, 266]. carbon [149, 145]. Carbosulfan [19]. carcass [38, 3]. Carcharhinus [37]. Carcinoscorpius [88]. carnitine [225]. carotene [369]. carotenoid [188, 369]. carp [77, 235, 322, 65, 373, 264, 32, 173, 64]. carpio [77, 235, 373, 16, 80, 340, 32, 31, 173, 64, 366]. carps [366]. cartilaginous [267]. case [341, 329, 366]. cashew [194]. Caspian [10, 245, 98, 18, 5, 52]. Cassiopea [347]. catalase [32]. Catfish [117, 309, 213, 105, 70, 178, 251, 118, 125, 103, 81, 278, 92, 317, 266, 321, 64]. Catla [65]. caudatus [53]. caught [181]. caviar [144]. cDNA [268]. Cellana [179]. cells [239, 229, 364]. cellular [346]. Central [25, 297, 195, 124, 97, 115]. cephalothorax [68]. cephalus [279]. Cerastoderma [370]. Ceratium [132]. cerevisiae [3]. Ceriodaphnia [356]. Chaetomorpha [137, 124]. chain [209]. chains [326]. challenged [16]. Chamelea [370]. Changes [77, 255, 193, 171, 248, 42, 156]. Channa [58, 19, 33]. characterisation [208]. characteristics [69, 216, 274, 66]. Characterization [290, 75, 188, 46, 279, 253, 344, 65, 359, 307, 76]. Characterizing [195]. charr [101]. Chemical [109, 9, 37, 360, 67]. Cherax [152]. Chetumal [143]. chicken [48]. Chile [159]. China [195]. Chinese [40, 43, 66]. Chinook [274]. Chironomidae [221]. chitosan [166, 16]. chloride [350, 298, 369, 317]. chloride-activated [298]. chlorophyll [220]. chromium [317]. Chromosome [138]. Chronic [63, 161]. chrysops [219]. chuii [204]. Cinnamon [358]. Cirrhinus [116]. cis [369]. City [220]. Cladoceran [231]. clam [91]. Clarias [117, 194, 180, 251, 125, 103, 278, 92, 64]. clear [315]. clinical [309]. cloning [268]. clownfish [261]. Co [137, 57]. Co-culture [137]. co-infected [57]. coast [230, 297, 112]. Coastal [372, 150, 160, 212, 342, 107, 220, 202]. cockle [277, 112]. cod [126, 38]. coelomic [86]. cold [324]. Colisa [99]. colonisation [206]. colonization [276, 221]. color [203]. coloration [316]. colors [341]. Colossoma [252, 338, 119]. colour [357]. column [216]. combating [47]. combination [196, 228, 130, 62, 28]. combined [273, 272, 308]. commercial [63, 191, 83, 219, 244, 175, 143]. commerson [162]. common [77, 373, 205, 173, 64, 366]. communis [340]. communities [148, 227, 134, 236]. community [315, 113, 329, 221, 348]. Comparative [102]. compared [191]. Comparison [294, 24, 194, 92, 115]. Complete [180]. compliance [240]. components [61]. composition [160, 211, 250, 82, 130, 172, 218, 38, 343, 109, 333, 52, 260]. compositions [349, 67]. compost [273, 272]. concentrate [83]. concentrates [163]. concentration [94, 280, 350, 19, 108]. concern [66]. Concomitant [60]. concrete [125]. conditions [40, 37, 56, 153]. conducted [231]. conserved [362]. construction [169]. consumption [43, 88]. containing [223, 6, 271]. contaminated [170]. contaminating [135]. content [346, 82, 223, 49, 313, 266]. Continuous [104]. contribution [152]. conversion [137]. cooking [9, 87]. copperoxychloride [90]. Coptodon [345]. coral [100, 188, 134]. Corbicula [280]. corn [274]. cornuta [356]. Correction [273, 239, 302]. Cortisol [126, 39]. costs [232]. counterparts [191]. Country [191, 212]. coupled [145]. covered [171]. Crab [265, 56, 88]. cracker [171]. Crassostrea [163]. Crayfish [8, 152, 305, 182]. creek [113]. cricket [223]. Critical [40, 365]. croaker [290, 197]. crocea [197]. cross [145]. cross-domain [145]. crossing [142]. cruentum [350]. cryopreservation [323]. Cu [43]. cucullata [14]. cucumber [217, 244, 333, 236]. cultivar [307]. Cultivation [203]. culture [149, 250, 125, 91, 151, 136, 50, 163, 137, 30, 351]. cultured [21, 191, 178, 313, 236, 348]. cultures [228, 104, 167, 140, 158]. current [318]. Cuttlefish [48, 208]. Cuvier [261]. cyanobacteria [115]. Cyanobacterium [158]. cyanuric [248]. cycle [315, 212, 328, 18, 329, 88]. Cypermethrin [193]. Cyprinidae [292]. Cyprinus [77, 235, 373, 16, 80, 340, 32, 31, 173, 64, 366]. cytochrome [268].

D [162]. D-loop [162]. Dactylogiridae [119]. daily [97]. dalli [183]. damage [233, 319, 340]. Danio [233, 282]. Danube [114]. daphnia [45]. dark [88]. Dasyatis [267]. data [202]. de-stratification [146]. Decadal [156]. decomposition [221]. decontaminating [121]. decussatus [370]. defenses [205]. deficiencies [218]. deforestation [190]. deformities [175]. degradation [186, 153]. degrading [296]. delta [113]. demersal [25]. Demonstration [135]. denitrifying [107]. density [71, 95, 150, 118, 348]. density-dependent [118]. dependent [118, 115]. deposition [144, 145]. depressus [56]. deprivation [201]. derived [199]. Description [97, 4]. desert [169, 140]. design [89]. detected [227]. detecting [361]. Detection [158]. detergents [143]. Determination [211, 282, 144, 349]. determine [152]. Determining [202]. detritus [221]. develop [240]. Development [361, 267, 189, 246, 279, 183, 305, 12, 208, 67]. devil [133]. dewatering [78]. DHA [323]. Dicentrarchus [174, 166, 75, 24, 49, 241, 284]. Diclofenac [305]. diet [194, 83, 179, 152, 278, 31]. Dietary [358, 179, 286, 185, 209, 248, 71, 314, 213, 356, 274, 251, 224, 90, 172, 241, 16, 98, 218, 38, 161, 207, 184]. diets [101, 178, 219, 293, 6, 91, 32, 284]. Difference [227]. differences [131]. Different [356, 77, 213, 40, 177, 148, 78, 313]. Differential [57]. differentiation [51, 262]. digestive [328, 65, 218, 42]. dihydroquercetin [286]. dinoflagellates [132]. diploid [225, 262]. Diplonychus [195]. Diplopsalis [132]. directive [240]. directly [237]. discriminate [174]. disease [213, 224, 75, 47]. disinfection [182]. disrupting [235]. dissolved [146]. distinct [187]. distribution [230, 8, 25, 147, 53, 366]. diurnal [92]. diverse [198]. diversicolor [239, 229]. Diversity [107, 157, 242, 245, 100]. DMRT1 [96]. DNA [233, 340]. DOC [145]. docosahexaenoic [323]. domain [145]. domesticus [223]. domingensis [203]. Donax [370]. dorsal [133]. dose [33]. doublesex [96]. doublesex-mab [96]. drivers [315]. drugs [295]. dry [27]. due [51, 372]. Dunaliella [369]. durans [60]. during [315, 155, 75, 87, 316, 23, 171, 42, 299, 329, 208, 192, 121]. dussumieri [37]. dyeing [337]. dynamic [11]. Dynamics [352, 150, 315, 354, 148, 296, 92, 335, 221].

E. [246]. early [118, 330, 285, 259, 338, 192]. earthen [354, 92]. East [220]. eastern [106]. Echinometra [177]. ecological [345]. ecology [1]. ecosystem [156, 263]. ecosystems [176, 360]. ECP [22]. Ecto [299]. Ecto- [299]. Editorial [73]. edule [370]. edulis [357]. eel [51]. Effect [95, 363, 272, 257, 323, 350, 251, 20, 82, 224, 293, 339, 22, 55, 62, 325, 357, 6, 78, 125, 241, 87, 16, 316, 300, 38, 32, 3, 310, 369, 278, 207, 173, 317, 146, 266, 321, 67, 348, 63, 166, 105, 110, 45, 225, 90, 371, 37, 144, 88, 273]. effectiveness [48]. effector [309]. Effects [71, 314, 9, 280, 229, 183, 274, 43, 291, 282, 99, 328, 172, 305, 285, 218, 252, 84, 66, 235, 356, 312, 155, 358, 250, 138, 177, 306, 289, 276, 175, 254, 239]. Efficacy [210, 121, 298]. efficiency [337]. effluent [121]. effluents [116]. egg [277]. eggs [182]. Eichhornia [221]. eight [349]. electrolytes [84]. elemental [370]. elements [174, 365]. elevation [227]. ELISA [279]. Elodea [61]. elvers [51]. embryo [258]. embryonic [246, 305]. embryos [126]. Emerging [47, 367]. emulsions [74]. enables [124]. encrasicolus [297]. Endocrine [235]. endophytic [334]. energy [117, 225, 232]. Engraulis [297]. enhanced [334, 209, 137, 355]. enhancer [93]. enhances [308]. enriched [6]. enrichment [45]. enterocin [60]. Enterococcus [339, 60]. environment [24, 306, 258, 263]. environmental [341, 28, 56]. enzymatic [122, 41, 29, 2, 310]. enzyme [94, 218, 89, 260]. enzymes [346, 87, 278, 266]. EPIC [142, 187]. Epilithic [283]. Epinephelus [246, 325]. episodes [30]. equipment [232]. ERK [321]. Erratum [151]. esakii [195]. esocinus [108]. essential [218, 247, 102, 335]. estuarine [212]. estuary [329, 108, 206, 132]. European [174, 241, 262, 284]. eutrophic [206]. Evaluation [213, 83, 275, 130, 264, 164, 91, 50, 163, 184, 240, 281, 258, 161, 327]. event [112]. evidence [206]. Examining [116]. existence [165]. Exogenous [31]. exon [142]. exon-priming [142]. Experimental [320]. experimentally [243]. export [135, 145]. exposed [77, 81]. Exposure [117, 63, 168, 126, 304, 268, 248]. expression [95, 126, 224, 324, 268, 286, 103, 214, 57, 167, 321, 33]. extender [323]. extension [48]. extracellular [22]. extract [314, 257, 282, 224, 93, 173, 355]. Extraction [120, 281]. extracts [217, 177, 306, 120]. Extremophiles [159]. extruded [223].

F2 [246]. Fabricius [55]. facility [216]. factor [39, 96]. factor-I [39]. factors [95]. faecalis [339]. faecium [60]. families [197]. family [292]. farm [119]. Farmed [255, 226, 174, 82, 324, 152, 75, 49, 34, 46]. farming [44, 11, 329]. farms [85, 28]. fasciatus [99, 172]. fasting [214]. fat [371, 144, 313]. fatty [211, 82, 99, 6, 24, 49]. fauna [198, 206]. faunal [150]. features [81]. FeCl [350]. fecundity [297]. fed [101, 178, 219, 179, 130]. feed [117, 178, 251, 353, 152, 204, 256, 137]. Feeding [345, 105, 155, 20, 164, 316, 18, 35]. feeds [85, 223, 287]. feedstuffs [271]. female [328, 294, 368, 313]. females [34]. ferric [350]. fertility [323]. fertilization [197]. fields [139]. fighting [316]. fillet [9, 234]. film [48, 171]. films [102]. finfish [1]. fingerling [32]. fingerlings [117, 363, 194, 180, 358, 251, 90, 252, 310, 31]. finished [85]. fipronil [338]. Fish [17, 101, 230, 226, 194, 160, 113, 85, 154, 290, 255, 1, 180, 240, 282, 354, 268, 223, 62, 371, 138, 210, 172, 170, 6, 139, 281, 49, 193, 28, 316, 80, 205, 238, 171, 19, 232, 151, 136, 60, 258, 284, 108, 123, 256, 119, 84, 267, 102, 206, 184, 72]. fish-based [240]. fish-wastes [284]. fisheries [181, 7]. fisherman [97]. fishery [366]. fishes [365, 342, 294]. fishing [211, 25, 97]. fishmeal [83, 274]. fishy [67]. flavescens [249]. flesh [274, 293, 49, 59]. floating [334]. flocculants [141]. flocculated [163]. Flocculation [344]. flora [64]. floral [150]. flounder [320]. flour [149]. flow [66]. Fluctuating [113]. fluctuations [21]. fluid [15, 86]. fluminea [280]. fly [223]. flying [170]. following [323, 126, 268]. Food [35, 201, 353, 123, 222]. Food-deprivation-induced [201]. force [360]. formaldehyde [182]. formation [320, 170]. formosana [79]. formulated [91]. fossilis [70, 81, 84, 317, 266, 321]. four [187, 293, 26]. fractions [223]. framework [240]. free [79]. frequency [202]. fresh [94, 27]. Freshwater [128, 356, 312, 141, 149, 275, 152, 287, 328, 210, 305, 193, 209, 19, 151, 136, 84, 266, 33, 66]. freshwaters [240]. fries [118]. frisii [10, 13, 23]. frissi [98]. frozen [94, 135]. fry [3, 299]. fucata [361]. fucoidan [224]. fucoidan-rich [224]. fuels [63]. full [371]. Fulvia [277]. functional [242, 263]. fungi [228, 260]. fungicide [90]. furca [132]. fuscoguttatus [246]. future [360, 318].

Gadus [126]. Gaetice [56]. Galician [191]. gallina [370]. galloprovincialis [191, 212, 293]. Gambusia [195]. gamitana [338]. Ganga [145]. gariepinus [117, 194, 180, 251, 125, 278, 92, 64]. garlic [110]. Gasterosteus [296, 18, 327]. gastrointestinal [1]. gel [62, 93]. gelatin [290, 48, 62, 129, 37, 171, 102, 46]. gelatin-based [48]. gelatine [122]. gelatinolytic [89]. Gelling [69]. gene [154, 224, 268, 286, 189]. gene-based [189]. general [269]. generating [197]. genes [126, 142, 187, 103, 135]. Genetic [5, 51]. genotoxic [58]. genus [131]. geographic [51]. Geographical [53, 127]. Germinal [15]. GF [364]. GF-1 [364]. GH [39]. giant [246, 275, 328, 209, 137]. GIFT [50]. gill [363]. gills [32]. Gilthead [174, 39, 106, 49, 175]. glance [44]. glandiforme [198]. globiferous [76]. glucan [308]. glucose [95, 323, 321]. glutathione [32]. gluten [274]. glycogen [266]. glyphosate [312, 19]. glyphosate-based [312]. Gnathonemus [35]. GnRH [167]. goatfish [211]. goby [164, 222]. goldband [211]. Goldfish [206, 237, 368]. gonad [325, 12, 313]. gonadal [99, 267]. good [111]. gourami [99]. Gracilaria [357, 203]. gracilis [256]. gradient [113, 227, 54]. granosa [112]. grapsid [56]. gravel [114]. Great [365]. Greece [26, 97, 25]. Greek [30]. green [259, 343]. greenwater [107]. grey [279]. ground [112]. groundwater [372]. grouper [246, 325, 364]. grow [316]. grow-out [316]. grown [120, 140]. Growth [168, 178, 219, 331, 98, 71, 77, 314, 117, 213, 273, 272, 356, 10, 39, 105, 239, 229, 180, 274, 358, 251, 282, 224, 11, 293, 339, 130, 118, 371, 286, 172, 373, 357, 125, 241, 193, 16, 259, 185, 218, 277, 248, 7, 3, 369, 31, 192, 134, 137, 124, 202, 348, 351]. grypus [108]. Guaribas [254]. guide [240]. Gulf [97, 14, 211, 100, 162, 134, 202]. Gut [123, 314, 342, 4, 339, 330].

H11 [89]. Habitat [276, 244, 260]. Habitat-specific [276]. habits [18, 35]. haddoni [253]. haematology [38]. hairpin [364]. Haliotis [239, 229]. Halosaccion [198]. Hamilton [309, 65]. handling [328]. haploid [70]. harbor [198]. harvest [141]. harvesting [30]. harveyi [355]. hasta [218]. hatchering [155]. hatchery [259, 359, 175, 234]. hatchery-reared [259]. hatching [323]. Hawaiian [179]. Hazardous [105]. head [17]. headwater [227]. health [176, 49, 252, 161]. heat [55, 319, 161]. heavy [196, 14, 33]. helminths [340]. hemato [251]. hemato-biochemical [251]. Hematological [119, 358, 371, 373, 351]. Hemibagrus [118]. Hemolymph [56]. hemolytic [177]. Henna [173]. hepatic [205]. hepatopancreas [248, 200]. hepatopancreatic [209]. herbicide [312]. herbivorous [123]. hermaphroditic [238]. Hermetia [223]. heterogeneity [296]. Heteropneustes [70, 81, 84, 317, 266, 321]. Heteroptera [195]. High [49, 124, 159, 55, 220, 222]. high-altitude [159]. high-nutrient [220]. highly [206]. Histamine [170]. Histological [96]. histology [235, 363, 371, 207]. histopathological [80, 81]. Historical [318]. history [296, 66]. Holothuria [217]. Horizontal [243]. hormonal [310]. hormone [39, 237]. hormones [21, 282, 98, 12, 23, 313]. horseshoe [88]. host [327]. host-parasite [327]. housefly [32]. HSP [57]. HSP-70 [57]. human [49]. Huso [21, 82, 41, 2, 207]. hybrid [246, 219, 319]. hybrids [269]. hydrocortisone [45]. hydrolysate [129, 67]. hydrolysates [290, 7, 17, 232]. hydrolysis [122, 41, 29, 2]. hydrophila [314, 110, 339, 16, 173]. hypersaline [303]. hypophthalmus [178]. Hypopthalmichthys [322]. hyposaline [215]. hypothalamic [167]. hypoxic [103].

Identification [279, 103, 326, 238]. IGF [39, 324, 214]. IGF-1 [214]. IGF-I [39, 324]. IHHNV [57]. II [298, 368, 151, 136]. illucens [223]. illumination [43]. image [144]. immature [167]. immersion [22]. immobilized [298]. Immune [309, 241, 16]. immunity [213, 110, 224, 173]. Immunoadjuvant [110]. immunoassays [279]. immunological [358, 373, 308, 355]. immunomodulatory [289]. immunostimulants [308, 47]. IMNV [57]. impact [225, 28]. impacted [231]. impairs [117]. implications [49, 335]. improve [204]. improved [50, 355]. Improvement [373]. improves [308]. in-situ [372]. in-vitro [281]. incorporated [171, 102]. incorporation [74]. increase [106]. increases [237]. increasing [363, 113]. incubated [182]. index [240]. indexes [28]. India [145, 115]. Indian [103]. indica [314]. indicator [116]. indicators [161, 34]. indices [178, 358, 250, 282, 310]. indigenous [287, 140]. Indonesia [260]. indoor [216, 330]. induce [70]. induced [193, 319, 214, 248, 201]. Induction [233, 238, 57]. industrial [116]. industry [287]. inermis [173]. inexpensive [269]. infected [166, 286, 340, 57, 173]. infection [314, 110, 308, 306, 209, 299, 355]. inferred [66]. infestation [291]. inflation [249]. Influence [259, 351, 244, 320, 33]. influenced [12]. information [341]. ingestion [222]. ingredient [256]. Inhibition [228, 364]. inhibitors [115]. inhibitory [115]. Inimicus [133]. initial [20, 125, 249]. ink [79]. inland [292]. innate [213]. inner [243]. inoculated [334]. inosinic [186]. insect [210, 289]. insecticide [338]. insight [244]. Insights [167, 101, 372]. insulin [39]. insulin-like [39]. integrated [203]. integrity [155, 209]. intensive [27, 250, 163]. inter [324]. inter-specific [324]. interstitial [276]. intertidal [242, 56]. intestinal [339]. intestinalis [123]. intestine [371, 207]. Intra [324]. Intra- [324]. intracellular [286]. introducing [144]. intron [142]. intron-crossing [142]. introns [154]. intrusion [372]. invasive [195]. invertebrate [198, 227]. invertebrates [226]. investigate [139]. involved [266]. involvement [368, 267]. involves [237]. ionic [56]. Iran [324, 359, 108, 202]. Iranian [292]. Island [100, 134]. isolated [226, 75, 359, 115, 60, 347]. isolates [127, 135]. Isolation [287, 133, 76]. isomerization [369]. Isostichopus [333]. isotope [152, 165]. isotopes [139]. Istiophorus [181]. Italy [8].

Janeiro [85]. Japan [294, 243]. Japanese [342, 277]. japonicus [36, 133, 186, 236]. Java [188]. javus [165]. jellyfish [131, 347]. JF [4]. juvenile [213, 269, 43, 45, 179, 210, 172, 325, 373, 243, 215, 164, 163, 222, 184]. juveniles [291, 185, 98, 262, 214, 66].

kagoshimensis [243]. Kamensky [10]. Kapi [109]. karst [54]. Kenya [331]. kernel [194]. key [346]. kinetic [33]. kinetics [153]. King [265]. kisspeptin [167]. kisspeptin- [167]. Klebsiella [135]. koi [235]. koji [307]. Korinthiakos [97]. Kuibyshev [148]. kuruma [36]. kutum [10, 98, 13, 23, 10, 13, 23].

L [49, 225, 76]. l-carnitine [225]. L-rhamnose-binding [76]. L. [63, 71, 77, 168, 196, 314, 95, 174, 90, 331, 24, 205, 165, 270, 175]. Labeo [110, 224, 306]. labrax [174, 166, 75, 24, 49, 241, 284]. lagoon [26, 303]. lake [315, 95, 331]. lakes [231, 365]. lanceolatus [246]. land [140]. large [197]. Larimichthys [197]. larvae [166, 265, 20, 223, 316, 88, 347, 221, 348]. larval [356, 246, 183, 27, 362, 130, 138, 210, 249, 192]. late [238]. Lates [363, 210, 234]. Latreille [88]. Laurentian [365]. Lawsonia [173]. lead [84, 266, 321]. leaf [221]. leatherjacket [129]. leaves [314, 273, 272]. lectin [133, 76, 332]. leg [204]. Length [292, 202]. lenticula [132]. Lepidopus [53]. leptin [167]. leptin-signalling [167]. leptolepis [69]. lesser [93]. Lethal [19]. Leuciscinae [292]. leucospilota [217]. level [282, 237, 360, 321]. levels [71, 113, 126, 15, 172, 98, 369, 317]. Life [18, 48, 330, 296, 285, 338, 192, 66]. life-history [296]. light [357, 88]. like [39]. limitation [54]. limpet [179]. linking [145]. Linnaeus [297, 285, 103, 18, 92]. lipase [60]. lipid [257, 205, 38, 153]. lipogenesis [346]. lipophilic [30]. Lippia [247]. liquid [273, 272]. Listeria [135]. Lithodes [265]. Litopenaeus [130, 68, 57, 204, 329, 271, 200, 184, 348]. live [287, 298]. lived [86]. liver [101, 319, 38, 32]. living [24]. liza [303]. Llancanelo [157]. Lmk. [191, 212]. local [345]. localization [61, 96]. locally [178]. Locus [187]. Locus-specific [187]. Loligo [79]. long [28, 86]. long-lived [86]. longitudinal [54]. longline [181]. longtail [202]. longtooth [325]. loop [162]. lotic [366]. low [251, 220, 348]. low-chlorophyll [220]. low-salinity [348]. lucioperca [167]. LvLTLC1 [332]. Lyko [285].

M. [219]. MA3 [361]. mab [96]. macroalgae [349]. macrobenthic [242]. Macrobrachium [328, 209]. macrolides [295]. macrophyte [61]. macropomum [252, 338, 119]. macrostigma [323]. maculatus [256]. maggot [32]. magmeal [32]. magnetic [101]. mahseer [339]. Main [315, 302, 288]. maindroni [208]. major [232, 330]. Malaysia [112]. Malaysian [339]. male [294, 331, 299]. Management [30, 8, 26, 259]. mangrove [260, 88, 132]. mannan [308]. marble [164, 222]. marbled [320]. marina [273, 272]. marine [228, 142, 295, 287, 370, 28, 343, 258, 72]. markers [5]. market [199]. marmorata [222]. marmoratus [164]. martensii [361]. masculinization [282]. Mass [112, 361]. maternally [182]. mathaei [177]. matrinxã [312]. maturation [105, 99, 238]. maturity [297]. mazut [258]. meal [101, 83, 180, 274, 210, 32, 284, 207, 184]. meals [289]. Measurement [39]. measurements [234]. measuring [267]. meat [255, 48]. Mechanical [370]. mechanism [72]. media [7]. mediated [306]. Mediterranean [293, 175]. melamine [248]. Melanin [271, 79]. Melanin-containing [271]. melanin-free [79]. Melanocyte [237]. melanosis [87]. Mercury [108]. mesocosm [216]. Mesoglein [131]. Mesolonghi [26]. mesopotamicus [247, 335]. Mesozooplankton [147]. Metabolic [101, 262, 200]. metabolism [255, 225, 237, 368, 266]. metabolites [225, 253]. metadata [28]. metal [230, 298, 33]. Metallothionein [33]. metals [196, 14, 121]. Metapenaeus [183]. method [361, 144, 313]. methodology [89]. methods [9, 281, 232, 302, 288]. methyl [108]. Mexico [143, 299]. microalga [278]. microalga-based [278]. Microalgae [256, 346, 293, 353, 344, 287, 199, 337]. microalgal [141, 91, 303, 163]. microarray [189]. Microbial [1, 62]. microbiota [330]. microcrustaceans [263]. microcystins [80]. microphylla [298]. microsatellite [5]. microscopy [360]. migrating [66]. mince [257, 67]. mine [231]. mineral [110]. Minerals [223]. mini [346, 353]. mini-review [346]. minimizing [232]. minireview [311]. Mitchell [192]. mitigations [151, 136]. mixed [331, 125]. mixed-sex [125]. model [11, 169]. models [26]. Molecular [100, 72, 326, 359, 349, 208]. Molina [265]. molitrix [322]. mollusc [370]. moluccensis [211]. monitoring [341, 189]. monocrotophos [233]. monoculture [275, 92]. monodon [273, 272, 326, 55, 137, 355]. Monogenea [119]. monophosphate [186]. monosex [331, 50]. morhua [126]. Mormyridae [35]. Moroccan [297]. Morocco [230]. Morone [219]. Morphological [359, 42, 270]. morphology [339]. mortalities [361]. mortality [202, 112]. mortem [255]. mosquito [316]. mossambicus [201]. Motility [322]. moulting [328]. movement [139]. mrigala [116]. mRNA [167, 33]. mtDNA [162]. mucous [239, 229]. mucus [264, 281]. Mugil [279, 303]. Mugilidae [26]. mullet [279]. multi [366]. multi-year [366]. multienzyme [207]. multiple [216, 165]. Multivariate [52]. municipal [116]. Musa [108]. muscle [36, 255, 55, 286]. muscle-specific [286]. muscles [116]. mussel [212, 293, 259, 30]. mussels [191]. mutagenic [58]. mutica [277]. Mycobiota [85]. mycotoxins [85]. mykiss [257, 358, 225, 11, 22, 371, 286, 359, 3, 310]. Mystus [105, 58]. Mytilus [191, 212, 293].

NADH [86]. nanoparticles [166, 105, 344, 373]. nanoscale [360]. nasutus [356]. native [195]. Natural [246, 257, 24]. naturally [152, 340, 57, 119]. NCIM5363 [60]. NCIM5427 [60]. necrosis [364]. negatively [329]. Neotropical [329, 247]. nervous [364]. net [178]. Neutrase [29]. newly [366]. newly-stocked [366]. Niger [113]. Nigeria [113, 35]. Nile [63, 71, 168, 196, 314, 95, 304, 155, 291, 250, 301, 90, 268, 275, 331, 125, 214, 38, 50, 299, 92, 67, 351]. niloticus [63, 71, 168, 196, 314, 95, 304, 250, 90, 268, 331, 125, 193, 214, 38, 50, 299, 92, 67]. nitrate [283, 285]. nitrogen [11, 7]. NMR [101]. noble [182]. nodules [121]. non [16, 50, 140]. non-arable [140]. non-improved [50]. non-specific [16]. nonglutinous [307]. nonylphenol [235]. normal [309]. north [230, 25, 188, 53]. north-west [230]. northern [100, 202, 146]. Norway [160]. novel [268, 133, 195]. NUAC09 [369]. nuclear [101]. nucleoside [164]. nucleotide [164]. number [66]. nursery [163]. nursing [118]. nutrient [314, 280, 220, 199, 115, 54]. Nutritional [353, 9, 191, 144].

observation [96]. Observations [106]. obtained [290]. occurrence [30]. occurring [152]. ocean [191]. ocellaris [261]. Ocimum [335]. odor [67]. off [100]. oil [194, 110, 346, 371, 171, 38, 247, 74, 153, 335]. oil-in-water [74]. oils [102]. oligosaccharide [308]. Oman [226, 162, 202]. Omani [147]. Oncorhynchus [257, 274, 358, 225, 11, 22, 371, 286, 359, 3, 310]. oocytes [15]. Open [140, 191, 158]. Oplegnathus [172]. opportunities [106, 360]. Optimization [141, 70, 122, 41, 29, 2, 89, 344]. orange [364]. orange-spotted [364]. Oreochromis [63, 168, 196, 314, 95, 149, 304, 250, 90, 268, 331, 125, 193, 122, 319, 214, 38, 201, 50, 299, 92, 67, 71]. organ [167]. organic [291, 217, 177, 209, 145, 203]. organisms [276]. organogenesis [42]. organophosphate [233]. organosomatic [250]. organs [155]. orientalis [187]. Origin [175]. ornamental [256]. ornata [120]. oscillations [241]. Oscillatoria [115]. osmotic [56]. Osphronemidae [300]. Osteichthyes [35]. other [95, 180]. otolith [320]. otoliths [174]. Outbreak [134, 47]. outbreaks [75]. ovarian [15, 144]. ovary [144]. Overexpressing [346]. overview [231, 199, 302, 288, 128]. oxidase [86]. oxidation [257, 153]. Oxidative [225, 340, 74, 319, 248, 204, 278]. oxide [344]. Oxyeleotris [164, 222]. Oxygen [88, 43, 146]. oxyrinchus [42]. oyster [14, 361, 163].

P450 [268]. Pacific [187, 362, 308, 87, 68, 248, 204, 153]. pacu [335]. Paddlefish [318]. paddy [139]. Pagrus [330, 59]. Pakistan [116]. Pallas [5]. palm [194, 171]. PAM [158]. Pampean [315]. pangas [309]. Pangasianodon [178]. Pangasius [309, 356]. papain [31]. Paphies [91]. parahaemolyticus [75, 308]. parameters [155, 358, 251, 371, 373, 185, 3, 313, 52, 116, 92, 119, 202, 351]. parasite [291, 327]. parasitic [306, 299]. parasitism [155]. parasitized [119]. parrot [172]. Partial [27, 279, 268, 371, 76]. particles [276]. partition [327]. paste [109]. pathogenic [217, 177]. pathogens [347]. pathological [248]. patterns [345, 366]. Pb [298, 266]. PCR [142, 154, 187, 361, 103]. pearl [361]. Pease [179]. pectinata [215]. pedicellariae [76]. peel [257]. pelagic [230, 216, 108]. pellets [341]. pen [215]. Penaeus [273, 272, 36, 362, 326, 55, 308, 367, 186, 137, 355]. Peninsular [112]. peptide [224]. Perca [249]. Perch [192, 219, 249]. Perciformes [363]. percolating [169]. performance [71, 314, 213, 178, 155, 180, 219, 358, 251, 250, 362, 11, 293, 339, 371, 328, 286, 331, 172, 125, 241, 185, 98, 252, 369, 137, 348]. Perils [151, 136]. period [168, 297]. periods [77]. peripheral [233]. Perna [259]. peroxidation [205]. Peroxiredoxin [208]. Persian [100, 134, 9, 94, 14, 45, 20, 15, 162, 12, 202]. persicus [9, 94, 45, 20, 15, 12]. perspective [263]. perspectives [302, 288]. pertusa [343]. Perugia [300]. Peruvian [119]. pesticide [233]. pesticides [117]. petersii [35]. petroleum [63]. pfluegeri [181]. pH [251, 88]. pharaonis [48]. pharmacokinetic [247]. philippinarum [243]. Philippine [343]. phosphate [317]. phosphorus [369]. photoperiod [259]. photosynthetic [159]. phototaxis [249]. phycoremediation [337]. Phylogenetic [162]. physical [109]. physico [9, 37, 360]. physico-chemical [9, 37, 360]. Physiological [304, 13, 247, 63, 95, 289, 300, 161, 335]. physiology [314, 363, 368]. Phytochemical [343]. Phytoplankton [277, 104, 280, 245, 148, 52]. phytoremediation [334, 121]. Piaractus [247, 335]. pigment [350]. pigmentation [59]. pigmented [347]. pigments [188]. pike [219]. pikeperch [167]. pileolus [76]. Pinctada [361]. pink [307]. Pinnidae [215]. pipi [91]. Pisces [363, 300]. pits [114]. pituitary [201]. Plackett [89]. plankton [228, 348]. planktonic [263, 189]. Plasma [12, 235, 237, 84]. plastic [341]. platypterus [181]. Poeciliidae [195]. Pollutants [254]. polluted [196, 58, 81]. pollution [113, 116]. polychaete [143]. polyclonal [332]. polyculture [275, 92]. polycultured [64]. polycystum [120, 355]. polymeric [141]. polymetallic [121]. Polyodon [318]. polyphenoloxidase [68]. polysaccharide [355]. polysaccharides [251]. polyunsaturated [99]. pond [178, 152, 169, 163, 140, 158]. ponds [354, 92, 146, 189]. populations [212, 342, 131, 5, 165, 352, 234]. porgy [59, 218]. Porphyridium [350]. Port [341]. Possible [368, 267]. post [323, 255, 348]. post-larvae [348]. post-mortem [255]. post-thaw [323]. postlarvae [204]. potato [149]. Potential [295, 356, 345, 1, 44, 287, 58, 164, 204, 347, 261]. poultry [83, 274, 371]. powder [48]. practical [179]. prawn [36, 183, 275, 328, 209, 137]. prawns [273, 272]. pre [87]. pre-cooking [87]. prebiotics [1]. predator [195]. predatory [125]. predict [261]. Prediction [234]. Predominant [330]. preference [261]. preferentially [283]. Preliminary [194, 123, 372]. Preparation [307, 138]. presence [80]. Present [360, 85]. pressure [55]. pretreatments [67]. prey [195]. primers [142, 154, 187]. priming [142]. prior [129]. PRL [39]. probiotic [3]. probiotics [1]. problems [106]. Procambarus [285]. process [105, 344]. processing [37, 144, 60]. produced [109]. product [83, 7]. Production [232, 332, 83, 149, 27, 350, 346, 17, 115, 60, 89, 302, 288, 30]. productive [250, 329]. products [22]. profiles [301, 349]. profiling [260]. programme [197]. prolactin [39]. Prominent [319]. promoter [31]. propagation [6]. properties [9, 290, 244, 62, 37, 78, 87, 360, 68, 38, 109, 102]. property [370]. Propofol [301]. Proportions [181]. Proposal [11]. prospects [318]. Protamex [41, 17]. protease [286, 115, 60]. protect [271]. protected [157]. protection [339]. Protective [166, 90, 306]. Protein [94, 71, 95, 257, 83, 142, 154, 180, 179, 55, 127, 172, 296, 218, 17, 232, 67]. protein-degrading [296]. proteins [36, 41, 2]. prototype [189]. provide [341]. proximate [333]. Pseudoalteromonas [347]. Pseudopleuronectes [320]. Pterophyllum [6]. PUFAs [99]. puffer [352]. punctata [33]. punctatus [58, 19]. Purification [65, 332]. Purus [190]. put [366]. pyrene [268]. pyrethroids [151, 136].

Qeshm [100, 134]. quadricarinatus [152]. Quality [326, 171, 323, 255, 114, 354, 293, 59, 135, 144, 3, 313, 190, 329, 52, 92, 348]. quantitative [361, 103]. Quintana [143].

rainbow [257, 358, 225, 11, 22, 371, 359, 3]. raised [330]. RAS [367]. rate [274, 277, 327, 366]. rates [323, 15, 243]. ratio [222, 137]. Ravi [116]. real [103]. real-time [103]. reared [210, 331, 24, 259, 184]. rearing [71, 356, 40, 241, 352, 128]. recirculating [50, 367]. recirculation [275]. recovery [121]. recruit [26]. recruits [125]. recycling [199]. red [269, 198, 118, 330, 59, 319, 150]. red-tailed [118]. redclaw [152]. Reduced [307, 86, 137]. Reduced-salt [307]. reduction [248]. reef [100]. reef-associated [100]. Reeve [239, 229, 243]. reference [103]. refrigerated [87]. regimes [315]. region [162]. regions [261]. regulating [72]. related [126, 255, 96, 86]. relation [54]. relationship [292, 195]. relationships [162]. release [11]. released [121]. Remarkably [222]. removal [298]. removing [196]. replace [210]. replacement [83, 27, 180, 371]. replacer [184]. replacing [38]. replication [364]. reproduction [356, 345, 34]. reproductive [235, 212, 328, 294, 368]. requirement [179, 185]. rerio [233, 282]. research [336]. Reservoir [148]. reservoirs [280]. residing [148]. resistance [213, 110, 224, 308, 209, 355]. resistant [253]. resonance [101]. resources [178]. respiration [327]. respond [263]. response [71, 16, 98, 57, 89, 208, 337, 200]. responses [304, 301, 241, 308, 13, 300, 303, 247, 355]. restraining [347]. resulting [337]. Reuse [250]. reversal [282, 299]. review [176, 346, 353, 289, 72]. rhamnose [76]. Rías [191]. ribosomal [142, 154]. rice [139, 307]. rich [224]. Rio [85]. risk [143]. River [220, 35, 145, 283, 58, 81, 190, 254, 61, 114, 116]. riverine [54]. RNAs [364]. rock [14]. rohita [110, 224, 306]. rohu [110]. role [47, 249]. Romsdalsfjord [160]. Roo [143]. rosenbergii [328, 209]. rostrata [51]. rotundicauda [88]. Roundup [312]. rRNA [189]. rubra [300]. rubripes [352]. ruckeri [22]. Ruditapes [370, 243]. Russia [148, 318]. ruthenus [238]. Rutilus [10, 98, 13, 23].

S [32]. S-transferase [32]. Saccate [198]. Saccharomyces [3]. Saccostrea [14]. sailfish [181]. salar [83]. salina [369]. saline [159]. Salinity [351, 113, 239, 229, 183, 43, 45, 241, 310, 88, 124, 348]. salinization [372]. Salmo [323, 83, 270]. salmon [83, 274]. salmonicida [226]. salt [307]. salted [109]. salts [78]. saltwater [372]. Salvelinus [101]. same [24]. sand [169]. Sander [167]. sandwicensis [179]. santolla [265]. Santos [341]. Saprolegnia [359]. Sarcophyton [188]. sardine [93]. Sardinella [93]. Sargassum [224, 120, 93, 355]. Sars [356]. sativum [110]. sauces [307]. Saudi [150]. saxatilis [219]. scabbardfish [53]. scalare [6]. scale [122, 97]. Schistocephalus [327]. Schlegel [187]. school [183]. schrenckii [46]. scientific [111]. scientometric [336]. Scomberomorus [162]. scope [261]. screening [260]. SE [191, 212]. Sea [150, 202, 174, 295, 217, 244, 253, 177, 75, 24, 49, 262, 333, 270, 76, 86, 175, 121, 236, 234, 10, 106, 245, 243, 188, 162, 18, 156, 52, 53]. seabass [241, 284]. seabream [39, 27, 106, 330]. Sealing [169]. season [12, 23]. Seasonal [21, 150, 245, 148]. seasonality [147]. seaward [66]. Seawater [270]. Seaweed [93, 241, 120]. secondary [302, 288]. sediment [108]. seepage [315]. Selangor [112]. Selaroides [69]. selected [150, 126, 28, 306, 266]. selective [197]. selenium [291]. Selenka [236]. semen [94]. sensitively [361]. sensory [191, 38]. Sepia [48]. Sepiella [208]. sequence [162]. sequences [127]. Serum [23, 371, 340, 317]. settlement [259, 347]. sewage [242]. sewage-affected [242]. sex [282, 331, 125, 12]. sexual [238]. Shadegan [345]. shallow [315, 148]. shape [40]. shark [37]. sharp [92]. sharptooth [125]. shelf [48, 147]. shelf-life [48]. shell [215]. shells [370]. shores [242]. short [126, 322, 227, 209, 364]. short-chain [209]. short-term [126, 322]. Shrimp [308, 311, 334, 362, 326, 130, 55, 87, 307, 171, 68, 57, 367, 248, 109, 204, 329, 186, 74, 153, 163, 332, 189, 184]. shrimps [135]. Siberian [4, 185, 313, 161, 34]. sidoides [247]. Siganus [165]. signalling [167]. significantly [86]. Silurus [213]. Silver [192, 166, 322, 53]. silvery [218]. silvery-black [218]. simulate [216, 11]. sinensis [40, 43]. situ [372, 58]. Size [297, 51, 125, 313]. skin [290, 48, 129, 37, 264, 281, 102, 46, 236]. slaughterhouse [199]. Slovakia [114]. small [113, 97]. small-scale [97]. sodium [269, 369]. soft [188]. soft-coral [188]. soldier [223]. Soleoidei [294]. solidus [327]. solution [298]. solvent [306]. solvents [281]. some [196, 150]. sonogrphy [313]. source [149, 353, 7]. sources [180, 102]. South [157]. southeast [18, 336]. southeastern [303]. Southern [265, 165, 10, 157, 52, 326]. southwestern [245]. soybean [38, 207, 184]. sp [149, 362, 361, 243, 188, 333, 89, 137, 124, 143]. space [276]. Spain [191]. Sparidentex [218]. Sparus [174, 39, 106, 24, 49, 175]. spat [259]. spathula [318]. spathulatus [119]. Spatial [296, 147]. spawning [246, 297, 99, 26, 23]. spearfish [181]. species [230, 160, 345, 292, 244, 324, 245, 287, 138, 26, 264, 359, 18, 115, 108, 349, 260]. Specific [347, 187, 324, 286, 16, 276]. specifically [361]. spectroscopy [101]. speed [40]. sperm [323, 322]. spermatogenesis [238]. spinefoot [165]. spines [133]. splendens [316]. splendid [79]. Sponge [260]. Sponge-associated [260]. spot [127, 96, 271, 200]. spotted [364]. spp [122, 319]. spraying [169]. squid [79]. squilla [166]. stability [48, 74, 153]. Stable [165, 152, 139]. stage [238]. stages [330, 285, 12, 338]. stagnant [124]. Staphylococcus [170]. State [85]. statistically [240]. status [63, 77, 150, 318, 56, 204, 161]. stellate [5]. stellatus [5]. step [232]. sterlet [238]. steroid [21, 282, 15, 12, 23, 313]. Stichodactyla [253]. stickleback [296, 18]. stimulant [164]. stimulating [237]. stinger [133]. stinging [70, 81, 317]. stingray [267]. stock [26]. stocked [196, 366]. Stocking [118, 95, 125, 348]. storage [322, 48, 87, 171, 153]. strain [361, 347]. strains [50, 203]. strait [220]. stranded [341]. strategic [155]. stratification [146]. streaked [165]. stream [276]. streams [227]. streptococci [177, 217]. Stress [34, 71, 126, 255, 45, 90, 328, 193, 103, 319, 340, 13, 248, 19, 204, 278, 161, 208]. stress- [126]. stress-induced [319]. stress-related [255]. stripe [69]. striped [105, 219, 99]. structural [61, 127, 370, 209]. structural-elemental [370]. structure [315, 113, 5, 329, 221, 348]. studies [231, 302, 288, 102]. Study [313, 320, 341, 80, 329, 153, 33, 366, 200]. sturgeon [9, 21, 94, 4, 40, 43, 44, 45, 20, 82, 15, 185, 12, 5, 42, 2, 144, 313, 161, 207, 34, 46, 66]. sub [292]. sub-family [292]. Subcellular [61]. subject [161]. subjected [241, 13]. submerged [61]. substitute [163]. substitution [274]. substrate [276]. suggests [165]. sulfate [373]. supplement [286]. supplementary [152]. Supplementation [308, 251, 241, 161, 31]. supplementations [323]. supply [326]. suppression [201]. Surface [190, 290, 89]. surimi [69, 62, 78, 93]. survey [206]. Survival [215, 273, 272, 183, 251, 130, 118, 243, 125, 259, 3, 204, 173, 124]. susceptibility [314]. suspension [110]. sustainable [353]. sweet [149]. swim [249]. swimming [40]. symbionts [188]. syndrome [127, 271, 200]. Synthetic [136, 325, 151]. System [367, 235, 157, 275, 107, 65, 220, 125, 54, 267, 327, 203, 184, 348, 351]. systems [159].

Tabuk [150]. tailed [118]. take [366]. Takifugu [352]. tamarind [314]. Tamarindus [314]. tambaqui [252]. tambroides [339]. tank [125]. tanks [330, 50]. tannery [337]. taurine [164]. taxonomy [147]. technical [240]. technological [232]. Technology [365, 250, 275, 121]. Technology-critical [365]. teleost [312, 33]. Teleostei [195]. Temminck [187]. temperate [205]. temperature [363, 183, 43, 20, 320, 241, 300, 88, 124]. temporal [296]. ten [292]. tenerrimum [93]. Terengganu [287]. term [126, 322, 28]. Terminalia [306]. terrestrial [145]. terrestris [282]. testicular [201]. testicular-axis [201]. testis [96]. Tetrapturus [181]. Tetraselmis [204]. textile [337]. Thai [369]. Thailand [165, 109, 120, 146, 124, 132]. thallus [198]. thaw [323]. thawed [94]. their [191, 281, 232, 254]. theoretical [327]. Thermal [261, 335, 208]. threadfin [62]. threats [8]. Three [155, 273, 272, 244, 293, 130, 96, 264, 115, 203, 143]. three-spot [96]. threespine [296]. threespined [18]. thrombocytes [291]. throughout [315, 34]. Thunnus [187, 29, 7, 17, 202]. thyroid [98]. tidal [220]. tiger [273, 272, 246, 326, 55, 352, 137]. tilapia [63, 71, 168, 196, 314, 95, 269, 149, 304, 155, 291, 250, 301, 90, 268, 275, 331, 125, 122, 319, 214, 38, 201, 50, 299, 92, 146, 67, 351]. time [215, 103]. times [87]. timing [20]. tissues [177, 80]. Tokunaga [243]. tolerance [270, 124, 261]. tonggol [202]. tool [11]. tooth [92]. Tor [339]. toxic [121, 30]. Toxicity [338, 77, 168, 312, 265, 19, 258, 143]. toxicogenic [254]. Toxopneustes [76]. toxopneustid [76]. Tra [178]. Trace [230, 174]. tract [1, 42]. trait [263]. traits [294, 296]. transcription [96]. transfer [145]. transferase [32]. transglutaminase [62]. transmontanus [144]. Transorb(R) [312]. transporter [95]. traps [160]. treated [373]. treatment [70, 334, 45, 325]. treatments [55]. trevally [69]. Tribullus [282]. trichlorfon [304]. trilobite [88]. triploid [225, 262]. Tris [323]. Tris-glucose [323]. trophic [150, 113]. tropical [280, 261]. trout [257, 323, 219, 358, 225, 11, 22, 371, 359, 3, 270]. trunculus [370]. trutta [323, 270]. trypsin [65]. tshawytscha [274]. Tsuyahime [307]. tuna [187, 181, 29, 7, 17, 202]. turbid [315]. Turbinaria [120]. Turkey [211, 44, 349]. Turkish [240]. two [365, 25, 131, 359, 132]. Type [151, 136]. types [273, 272, 130, 148]. ulceration [236]. Ulva [343, 123]. unicorn [129]. Universal [142, 154]. unsaturated [6]. Update [311]. Upeneus [211]. uptake [277]. urban [283]. urchin [177, 76]. urchins [86]. Use [269, 7, 258, 93, 194, 356, 283, 311, 279, 176, 362, 360, 182]. used [232]. uses [287]. Using [174, 152, 139, 101, 228, 141, 361, 344, 162, 28, 41, 29, 227, 5, 2, 17, 3, 169, 143, 140, 158, 202, 234]. utilisation [178, 199]. utilization [314, 117, 251, 192]. UV [70].

vaccination [22]. Valenciennes [303]. valid [240]. validation [191]. vannamei [362, 130, 308, 68, 57, 367, 204, 329, 271, 200, 184, 348]. Variability [212, 324]. variables [168]. variation [51]. Variations [160, 127, 333, 243, 205]. various [281, 102]. varying [56]. vera [251]. Veracruz [299]. Verbenaceae [247]. vertical [243]. vesicle [15]. Viability [149]. Vibrio [166, 342, 361, 75, 308, 352, 355]. Victoria [331, 95]. virginalis [285]. viridis [259]. virulent [135]. virus [127, 271, 364, 200]. viscera [29]. visceral [41, 2]. vitamin [90, 6, 373]. vitellogenin [235]. vitellogenins [279]. vitro [281]. vittatus [105, 58]. vulnerability [319]. vulnerable [157].

W [26]. warm [324]. warm-water [324]. waste [41, 2, 60]. wastes [196, 284]. wastewater [334, 199, 329, 337, 140]. Water [354, 77, 168, 196, 21, 363, 150, 315, 216, 240, 114, 250, 20, 324, 320, 287, 215, 58, 231, 259, 300, 81, 310, 190, 352, 329, 52, 116, 92, 146, 74, 163, 124, 348, 254]. water-born [77, 168]. water-column [216]. waters [292, 148, 124, 202]. weed [123, 158]. weight [292, 82, 349, 234]. welfare [255, 161]. west [230]. Western [160, 183, 156]. wet [199]. wetland [345, 157]. wetlands [334]. white [362, 130, 127, 87, 68, 57, 248, 144, 204, 153, 271, 200]. white-leg [204]. whitecheek [37]. Whiteleg [308, 367, 184]. wightii [224]. wild [226, 174, 106, 24]. within [135]. without [354, 171]. wrasse [96]. writing [111]. wyckioides [118].

xamachana [347]. Xiphophorus [256]. xylosus [170].

year [366]. yearling [270]. yeast [311, 3]. yellow [69, 249, 197]. yellowfin [29, 7, 17]. yellowtail [78]. Yenisei [61]. Yersinia [22]. Yield [26, 125, 144, 234]. yokohamae [320]. Yolk [192]. York [220, 220].

Zealand [91]. Zebra [282]. zebrafish [233]. zillii [345]. zinc [77, 168, 373, 185]. ZnS [105]. zoanthids [100]. Zooplankton [157, 302, 288, 315, 356, 329]. zygomycetes [101].

References

Denev:2009:MEG

[1] Stefan Denev, Georgi Beev, Yordan Staykov, and Rumyana Moutafchieva. Microbial ecology of the gastrointestinal tract of fish and the potential application of probiotics and prebiotics in finfish aquaculture. International Aquatic Research, 1(1):1–29, November 2009. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_673235.html;   https:/
/intelaquares.tonekabon.iau.ir/article_673235_fc3c20524391059180cdbf20d0193190.
pdf.

Ovissipour:2009:OEH

[2] Mahmoudreza Ovissipour, Majid Taghiof, Ali Motamedzadegan, Barbara Rasco, and Abbas Esmaeili Molla. Optimization of enzymatic hydrolysis of visceral waste proteins of beluga sturgeon Huso huso using Alcalase. International Aquatic Research, 1(1):31–38, November 2009. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_673236.html;   https:/
/intelaquares.tonekabon.iau.ir/article_673236_c26c62b230887aceaadf650e7b987966.
pdf.

Pooramini:2009:EUY

[3] Molook Pooramini, Abolghasem Kamali, Abdolmajid Hajimoradloo, Morteza Alizadeh, and Rasoul Ghorbani. Effect of using yeast (Saccharomyces cerevisiae) as probiotic on growth parameters, survival and carcass quality in rainbow trout Oncorhynchus mykiss fry. International Aquatic Research, 1(1):39–44, November 2009. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_673238.html;   https:/
/intelaquares.tonekabon.iau.ir/article_673238_980686a01d8931cd125fbeebd8945a65.
pdf.

Dapra:2009:GSS

[4] Franco Daprà, Francesco Gai, Giovanni B. Palmegiano, Benedetto Sicuro, Mimmo Falzone, Karine Cabiale, and Marco Galloni. Siberian sturgeon (Acipenser baeri, Brandt JF 1869) gut: anatomic description. International Aquatic Research, 1(1):45–60, November 2009. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_673239.html;   https:/
/intelaquares.tonekabon.iau.ir/article_673239_d7239ba016999921bd87d110a6289765.
pdf.

Norouzi:2009:GSG

[5] Mehrnoush Norouzi and Mohammad Pourkazemi. Genetic structure of Caspian populations of stellate sturgeon, Acipenser stellatus (Pallas, 1771), using microsatellite markers. International Aquatic Research, 1(1): 61–65, November 2009. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_673240.html;   https://intelaquares.tonekabon.iau.ir/
article_673240_b57803288aae51a1141bdbda6c7631da.pdf.

Langroudi:2009:EDC

[6] Hadi Ershad Langroudi, Seyed Hamed Mousavi, Bahram Falahatkar, and Zeynab Moradkhani. Effect of diets containing artemia enriched with unsaturated fatty acids and vitamin C on angel fish Pterophyllum scalare propagation. International Aquatic Research, 1(1):67–72, November 2009. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_673241.html;
https://intelaquares.tonekabon.iau.ir/article_673241_e8731756acec8b7b9cb8d4b72df002ac.
pdf.

Ovissipour:2009:UHY

[7] Mahmoudreza Ovissipour, Reza Safari, Ali Motamedzadegan, Barbara Rasco, Reza Pourgholam, Elaheh Mohagheghi, and Abbas Esmaeili Molla. Use of hydrolysates from yellowfin tuna Thunnus albacares fisheries by-product as a nitrogen source for bacteria growth media. International Aquatic Research, 1(1):73–77, November 2009. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_673243.html;   https:/
/intelaquares.tonekabon.iau.ir/article_673243_f5d23c75eef9a0d913302a08868a7601.
pdf.

Aquiloni:2010:CGI

[8] Laura Aquiloni, Elena Tricarico, and Francesca Gherardi. Crayfish in Italy: distribution, threats and management. International Aquatic Research, 2(1):1–14, April 2010. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_672550.html;   https://intelaquares.tonekabon.iau.ir/
article_672550_de2f00dbefd0d7d531c4d35065f7072f.pdf.

Alipour:2010:ECM

[9] Hakimeh Jannat Alipour, Bahareh Shabanpoor, Ali Shabani, and Alireza Sadeghi Mahoonak. Effects of cooking methods on physicochemical and nutritional properties of Persian sturgeon Acipenser persicus fillet. International Aquatic Research, 2(1):15–23, April 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677678.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677678_f7aef1f49ba0a7169a0883ca9b6789e2.
pdf.

Bandpei:2010:AGK

[10] Mohammad Ali Afraei Bandpei, Mashhor Mansor, Shahram Abdolmalaki, Farhad Keymaram, Mansor Mohamad Isa, and Ali Asghar Janbaz. Age and growth of kutum (Rutilus frisii kutum, Kamensky, 1901) in southern Caspian Sea. International Aquatic Research, 2(1):25–33, April 2010. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677681.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677681_dbfad2584bed187cdc2f248a1b335d27.
pdf.

Guo:2010:PDM

[11] Kaijun Guo, Laura Gasco, Ivo Zoccarato, Giovanni B. Palmegiano, and Benedetto Sicuro. Proposal of a dynamic model as a tool to simulate growth performance and nitrogen release in rainbow trout Oncorhynchus mykiss farming. International Aquatic Research, 2(1):35–47, April 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_677682.html;
https://intelaquares.tonekabon.iau.ir/article_677682_f025255d99b0d418e1b2ccb70eb1236f.
pdf.

Nazari:2010:PSS

[12] Rajab Mohammad Nazari. Plasma sex steroid hormones of Persian sturgeon Acipenser persicus as influenced by gonad development stages and season. International Aquatic Research, 2(1):49–54, April 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677683.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677683_9e0a318382efae3342b40b6ec7cec934.
pdf.

Nikoo:2010:PSR

[13] Mehdi Nikoo, Bahram Falahatkar, Masood Alekhorshid, Behzad Nematdost Haghi, Ali Asadollahpour, Mosa Zarei dangsareki, and Hamid Faghani. Physiological stress responses in kutum Rutilus frisii kutum subjected to captivity. International Aquatic Research, 2(1):55–60, April 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_677685.html;

https://intelaquares.tonekabon.iau.ir/article_677685_f6590dfc3d2f831c4c962234b08cadcf.
pdf.

Azarbad:2010:BBH

[14] Hamed Azarbad, Arash Javanshir Khoi, Alireza Mirvaghefi, Afshin Danekar, and Maryam Shapoori. Biosorption and bioaccumulation of heavy metals by rock oyster Saccostrea cucullata in the Persian Gulf. International Aquatic Research, 2(1):61–69, April 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677686.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677686_79a91a923eec5d0b5f8da30c82d68137.
pdf.

Hajirezaee:2010:GVB

[15] Saeed Hajirezaee, Gholam Reza Rafiee, Ramtin Hushangi, Ruhollah Rahimi, Hamid Niksirat, and Rezvan Kazemi. Germinal vesicle breakdown rates in oocytes and steroid levels in blood and ovarian fluid of the Persian sturgeon Acipenser persicus. International Aquatic Research, 2(1):71–75, April 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677687.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677687_df0152df8184b68686e01da122d04a4b.
pdf.

Maqsood:2010:EDC

[16] S Maqsood, Prabjeet Singh, Munir Hassan Samoon, and Amjad Khansaheb Balange. Effect of dietary chitosan on non-specific immune response and growth of Cyprinus carpio challenged with Aeromonas hydrophila. International Aquatic Research, 2(2):77–85, August 2010. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677688.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677688_9b9b9a868e972ce5a480660c395129c3.
pdf.

Ovissipour:2010:FPH

[17] Mahmoudreza Ovissipour, Soottawat Benjakul, Reza Safari, and Ali Motamedzadegan. Fish protein hydrolysates production from yellowfin tuna Thunnus albacares head using Alcalase and Protamex. International Aquatic Research, 2(2):87–95, August 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677689.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677689_ae12b80c46816c080333dfe515dcabed.
pdf.

Niksirat:2010:LCF

[18] Hamid Niksirat, Azadeh Hatef, and Asghar Abdoli. Life cycle and feeding habits of the threespined stickleback Gasterosteus aculeatus (Linnaeus, 1758): an alien species in the southeast Caspian Sea. International Aquatic Research, 2(2):97–104, August 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677690.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677690_0f3cad072c2e0fb3a54a37b4452cc9a3.
pdf.

Nwani:2010:LCT

[19] Christopher Didigwu Nwani, Naresh Sahebrao Nagpure, Ravindra Kumar, Basdeo Kushwaha, Pavan Kumar, and Wazir Singh Lakra. Lethal concentration and toxicity stress of Carbosulfan, Glyphosate and Atrazine to freshwater air breathing fish Channa punctatus (Bloch). International Aquatic Research, 2(2):105–111, August 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677692.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677692_56d5d90894a38996526c1816d179b1ae.
pdf.

Ghelichi:2010:EWT

[20] Afshin Ghelichi, Nourmohammad Makhdoomi, Sarah Jorjani, and Ali Taheri. Effect of water temperature on the timing of initial feeding of Persian sturgeon Acipenser persicus larvae. International Aquatic Research, 2(2):113–119, August 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677693.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677693_02fbd75e19ce6dd42dddbf9caae40ad0.
pdf.

Alizadeh:2010:SFS

[21] Morteza Alizadeh, Ali akbar Hedayati, Mahmoud Bahmani, and Hamid Reza Jamalzadeh. Seasonal fluctuations of steroid hormones in beluga sturgeon Huso huso cultured in brackish water. International Aquatic Research, 2(2):121–126, August 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677694.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677694_e805ca8b1e87b264e26dcbc14f5c326a.
pdf.

Ispir:2010:EIB

[22] Unal Ispir and Mustafa Dorucu. Effect of immersion booster vaccination with Yersinia ruckeri extracellular products (ECP) on rainbow trout Oncorhynchus mykiss. International Aquatic Research, 2(2):127–130, August 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_677695.html;
https://intelaquares.tonekabon.iau.ir/article_677695_566cda2d8f7f226bf15410eaef1ac774.
pdf.

Nikoo:2010:SSH

[23] Mehdi Nikoo, Hamid Faghani, and Abbas Esmaili Molla. Serum steroid hormones in kutum Rutilus frisii kutum during spawning season. International Aquatic Research, 2(2):131–133, August 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677696.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677696_b54b34cef7b46dc3e630f56df851602e.
pdf.

Lenas:2010:CFA

[24] Dimitrios Lenas, Panaghiotis Loghothetis, Grigorios Kanlis, and Cosmas Nathanailides. Comparison of fatty acids in the brains of wild and reared sea bass Dicentrarchus labrax L. and sea bream Sparus aurata L., and living in the same natural environment. International Aquatic Research, 2(3):135–141, December 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677697.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677697_d833ca5af76eff9dd743726d28529b8b.
pdf.

Klaoudatos:2010:ADD

[25] Dimitris Klaoudatos, Costas Kapiris, Alexios Conides, Costas Papaconstantinou, and Sotirios Glykokokkalos. Assessment of demersal distribution between two fishing areas in North and Central Greece. International Aquatic Research, 2(3):143–154, December 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677698.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677698_3a0081230e17535535e5f224667f0ebc.
pdf.

Katselis:2010:YPR

[26] George Katselis, Katerina Koukou, and Dimitrios Moutopoulos. Yield per recruit and spawning stock biomass models for the management of four Mugilidae species in Mesolonghi–Aitoliko lagoon (W. Greece). International Aquatic Research, 2(3):155–162, December 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677699.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677699_8077c753f638ca9b5745ba401c56a6c4.
pdf.

Dimitrios:2010:PRF

[27] Dimopoulos Dimitrios, Panagiotaki Panagiota, Golomazou Eleni, Malandrakis Emmanuil, Dadali Olga, Martsikalis Petros, Exadactylos Athanassios, Klaoudatos Spyridon, and Neofytou Christos. Partial replacement of fresh algae with dry, in intensive seabream larval production. International Aquatic Research, 2(3):163–166, December 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677700.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677700_82765cf69cda110470b2dd3a155a0113.
pdf.

Mavraganis:2010:CSI

[28] Theo Mavraganis, Trevor Telfer, and Cosmas Nathanailides. A combination of selected indexes for assessing the environmental impact of marine fish farms using long term metadata analysis. International Aquatic Research, 2(3):167–171, December 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677701.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677701_a47472620f33cef6fc445ff3244392e5.
pdf.

Motamedzadegan:2010:OEH

[29] Ali Motamedzadegan, Bahareh Davarniam, Gholamhassan Asadi, Abdolmohammad Abedian, and Mahmoudreza Ovissipour. Optimization of enzymatic hydrolysis of yellowfin tuna Thunnus albacares viscera using Neutrase. International Aquatic Research, 2(3):173–181, December 2010. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677702.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677702_0b4158b6b3c461135a9fefa46d55c5c2.
pdf.

Vlamis:2010:MPH

[30] Aristidis Vlamis, Panagiota Katikou, and Dimitrios Petridis. Management of the production and harvesting in Greek mussel culture areas, according to the occurrence of lipophilic toxic episodes. International Aquatic Research, 2(3):183–192, December 2010. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677703.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677703_802eb32e79b99775517fff5c337181c9.
pdf.

Singh:2011:ESP

[31] Prabjeet Singh, Sajid Maqsood, Munir Hassan Samoon, Vikas Phulia, Mohd Danish, and Ramesh Singh Chalal. Exogenous supplementation of papain as growth promoter in diet of fingerlings of Cyprinus carpio. International Aquatic Research, 3(1):1–9, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677704.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677704_b89c3e9ba8c6ea1e29d88f1391c2380e.
pdf.

Ogunji:2011:EHM

[32] Johnny O. Ogunji, Jorge Nimptsch, Claudia Wiegand, Carsten Schulz, and Bernhard Rennert. Effect of housefly maggot meal (magmeal) diets on catalase, and glutathione S-transferase in the liver and gills of carp Cyprinus carpio fingerling. International Aquatic Research, 3(1):11–20, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_677705.html;
https://intelaquares.tonekabon.iau.ir/article_677705_481716285a12bfd3997f4a1fbce55ecc.
pdf.

Tiwari:2011:MME

[33] Mohit Tiwari, Naresh Sahebrao Nagpure, Wazir Singh Lakra, and Devendra Nath Saksena. Metallothionein mRNA expression in freshwater teleost, Channa punctata (Bloch) under the influence of heavy metal, cadmium — a dose kinetic study. International Aquatic Research, 3(1): 21–29, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
677706.html;   https://intelaquares.tonekabon.iau.ir/article_
677706_983dc041018ec2f3072dd604fe8dcb61.pdf.

Williot:2011:SIT

[34] Patrick Williot, Sylvain Comte, and Françoise Le Menn. Stress indicators throughout the reproduction of farmed Siberian sturgeon Acipenser baerii (Brandt) females. International Aquatic Research, 3(1):31–43, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_677707.html;
https://intelaquares.tonekabon.iau.ir/article_677707_939c05d469c381254f4baecc41e2ff17.
pdf.

Nwani:2011:FFH

[35] Christopher Didigwu Nwani, Gregory Ejike Odoh, Emmanuel Fame Ude, and Okechukwu Iduma Okogwu. Food and feeding habits of Gnathonemus petersii (Osteichthyes: Mormyridae) in Anambra River, Nigeria. International Aquatic Research, 3(1):45–51, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677709.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677709_c909190ddcc3c0557a1cdead1c5b1263.
pdf.

Benjakul:2011:AAA

[36] Soottawat Benjakul, Wonnop Visessanguan, Tanong Aewsiri, Munehiko Tanaka, and Mehdi Nikoo. ATPase activities and autolysis of kuruma prawn Penaeus japonicus muscle proteins. International Aquatic Research, 3(1):53–61, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677711.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677711_602f02dfb1e456bd625ac8a0a1e3f22b.
pdf.

Kharyeki:2011:EPC

[37] Mina Esmaeili Kharyeki, Masoud Rezaei, and Ali Motamedzadegan. The effect of processing conditions on physico-chemical properties of whitecheek shark (Carcharhinus dussumieri) skin gelatin. International Aquatic Research, 3(1):63–69, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677712.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677712_bc3594319ee67c2d58218f2e6fbe7f0b.
pdf.

Ochang:2011:ERC

[38] Stephen N. Ochang. Effect of replacing cod liver oil with soybean oil as dietary lipid on carcass composition, haematology and sensory properties of the Nile tilapia Oreochromis niloticus. International Aquatic Research, 3(1):71–77, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677713.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677713_1114fc1ee7647fdaf26ebb0d931f0ec7.
pdf.

Castritsi-Catharios:2011:MPP

[39] Jeanne Castritsi-Catharios, Vassiliki Syriou, Konstantina Mandrou, and Nektarios Ladas. Measurement of prolactin (PRL), growth hormone (GH), insulin-like growth factor-I (IGF-I), cortisol and insulin in gilthead seabream Sparus auratus. International Aquatic Research, 3(1):79–82, April 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
677714.html;   https://intelaquares.tonekabon.iau.ir/article_
677714_0c401e4c97b5508ba151a3ee7c622a94.pdf.

Duan:2011:CSS

[40] Ming Duan, Yi Qu, Juan Yan, Longzhen Zhang, and Ping Zhuang. Critical swimming speed associated with body shape of Chinese sturgeon Acipenser sinensis under different rearing conditions. International Aquatic Research, 3(2):83–91, August 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677719.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677719_893ebe5743b17f01cf86eb405b76e848.
pdf.

Molla:2011:OEH

[41] Abbas Esmaeili Molla and Hrachya G. Hovannisyan. Optimization of enzymatic hydrolysis of visceral waste proteins of beluga Huso huso using Protamex. International Aquatic Research, 3(2):93–99, August 2011. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677720.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677720_efbfaf00d6e5a9331a078eade6202d05.
pdf.

Ostaszewska:2011:MCD

[42] Teresa Ostaszewska, Maciej Kamaszewski, Ryszard Kolman, Grzegorz Wiszniewski, Dobrochna Adamek, and Arkadiusz Duda. Morphological changes in digestive tract of Atlantic sturgeon Acipenser oxyrinchus during organogenesis. International Aquatic Research, 3(2):101–105, August 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_677733.html;
https://intelaquares.tonekabon.iau.ir/article_677733_306d6d4dc9642e446f28a8ec92b07843.
pdf.

Duan:2011:ETS

[43] Ming Duan, Jianyi Liu, Liang Qu, Longzhen Zhang, Guangpeng Feng, Tao Zhang, and Ping Zhuang. Effects of temperature, salinity, illumination and Cu^{2+} on oxygen consumption of juvenile Chinese sturgeon Acipenser sinensis. International Aquatic Research, 3(2):107–115, August 2011. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677734.html;   https:/

/intelaquares.tonekabon.iau.ir/article_677734_edbfed5e41ae3c9a97b74971e52090ef.
pdf.

Ercan:2011:GSF

[44] Ertan Ercan. A glance on sturgeon farming potential of Turkey. International Aquatic Research, 3(2):117–124, August 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677735.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677735_a6ad81af4d56cfeb54c4b43323f12e70.
pdf.

Fakharzadeh:2011:EHT

[45] Seyed Mohammad Esmael Fakharzadeh, Mehrdad Farhangi, Bagher Mojazi Amiri, Mohammad Ahmadi, and Nastaran Mazloumi. The effect of hydrocortisone treatment by bathing and daphnia enrichment on the salinity stress in Persian sturgeon Acipenser persicus juvenile. International Aquatic Research, 3(2):125–133, August 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677736.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677736_89701c2c19ced6cb51398a750a569780.
pdf.

Xu:2011:CGS

[46] Xueming Xu, Mehdi Nikoo, Soottawat Benjakul, Gangchun Xu, Juan Carlos Ramirez-Suarez, Ali Ehsani, Ladislaus Manaku Kasankala, Xiang Duan, and Shabbar Abbas. Characterization of gelatin from the skin of farmed Amur sturgeon Acipenser schrenckii. International Aquatic Research, 3(2):135–145, August 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677737.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677737_9d6108801567acabd3a3e5a868c6e661.
pdf.

Maqsood:2011:ERI

[47] Sajid Maqsood, Prabjeet Singh, Munir Hassan Samoon, and Khusheeba Munir. Emerging role of immunostimulants in combating the disease outbreak in aquaculture. International Aquatic Research, 3(3):147–163, December 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
677738.html;   https://intelaquares.tonekabon.iau.ir/article_
677738_a1095f73caaeba5d0c4557dd5c1b783b.pdf.

Hoque:2011:CSP

[48] Md. Sazedul Hoque, Soottawat Benjakul, and Thummanoon Prodpran. Cuttlefish Sepia pharaonis skin gelatin-based film: storage stability and its effectiveness for shelf-life extension of chicken meat powder. International Aquatic Research, 3(3):165–179, December 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677739.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677739_2ed0effc88230a9e28c38e8624adf095.
pdf.

Lenas:2011:HCF

[49] Dimitrios Lenas and Cosmas Nathanailides. High content of n − 6 fatty acids in the flesh of farmed fish sea bass Dicentrarchus labrax l. and gilthead sea bream Sparus aurata l.: implications for the human health. International Aquatic Research, 3(3):181–187, December 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677741.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677741_bd7bb511dd566ed82cdf779833a4ac9b.
pdf.

Ridha:2011:EMC

[50] Mohammad T. Ridha. Evaluation of monosex culture of GIFT and nonimproved strains of Nile tilapia Oreochromis niloticus in recirculating tanks. International Aquatic Research, 3(3):189–195, December 2011. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677742.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677742_fa7a462b000600d8fa3c30506238fab1.
pdf.

Jamandre:2011:GVS

[51] Brian Wade Jamandre, Kang-Ning Shen, and Wann-Nian Tzeng. Is the geographic variation in size of American eel Anguilla rostrata elvers due to genetic differentiation? International Aquatic Research, 3(3):197–203, December 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
677743.html;   https://intelaquares.tonekabon.iau.ir/article_
677743_ac91e7d5b82c21a70c5f2bfd21fa4dd3.pdf.

Saravi:2011:MAW

[52] Hassan Nasrollahzadeh Saravi, Asieh Makhlough, Reza Pourgholam, Zubir Bin Din, and Swee Yeok Foong. Multivariate analysis of water quality parameters and phytoplankton composition in the southern of Caspian Sea. International Aquatic Research, 3(3):205–216, December 2011. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677744.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677744_f69ea5f9f5162f66e0826f2fa48fc0fb.
pdf.

Torre:2011:GBD

[53] Michele Torre, Argyris Kallianiotis, Benedetto Sicuro, and Vassiliki Tsavalou. Geographical and bathymetric distribution of silver scabbardfish Lepidopus caudatus in North Aegean Sea. International Aquatic Research, 3(3):217–226, December 2011. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677745.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677745_0ef9f285506c3e8279df4511676ff4e8.
pdf.

Penick:2012:ABA

[54] Mary D. Penick, Scott A. Grubbs, and Albert J. Meier. Algal biomass accrual in relation to nutrient availability and limitation along a longitudinal gradient of a karst riverine system. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-4-20.

Jantakoson:2012:EHP

[55] Thitima Jantakoson, Kongkarn Kijroongrojana, and Soottawat Benjakul. Effect of high pressure and heat treatments on black tiger shrimp (Penaeus monodon Fabricius) muscle protein. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-4-19.

Nanba:2012:HOI

[56] Takeshi Nanba, Hideya Takahashi, Tsukasa Abe, Waichirou Godo, Maho Ogoshi, Hirotaka Sakamoto, Naoaki Tsutsui, and Tatsuya Sakamoto. Hemolymph osmotic, ionic status, and branchial Na^{+} /K^{+}-ATPase activity under varying environmental conditions in the intertidal grapsid crab, Gaetice depressus. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1186/2008-6970-4-18.

NogueiraVieira-Girao:2012:DIH

[57] Patrícia Raquel Nogueira Vieira-Girão, Ítalo Régis Castelo Branco Rocha, Francisco Hiran Farias Costa, and Gandhi Rádis-Baptista. Differential induction of HSP-70 expression in response to IHHNV in white shrimp Litopenaeus vannamei naturally co-infected with IHHNV and IMNV. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1186/2008-6970-4-17.

Kushwaha:2012:LSA

[58] Basdeo Kushwaha, Sanjay Pandey, Shilpi Sharma, Rashmi Srivastava, Ravindra Kumar, Naresh Nagpure, Anurag Dabas, and Satish Srivastava. In situ assessment of genotoxic and mutagenic potential of polluted river water in Channa punctatus and Mystus vittatus. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-4-16.

Manganaro:2012:APF

[59] Antonio Manganaro, Marilena Sanfilippo, Gianluca Fortino, Franco Daprà, Giovanni Battista Palmegiano, Francesco Gai, Elvira Lembo, Alessandra Reale, and Marisa Ziino. Artificial pigmentation and flesh quality in red porgy (Pagrus pagrus). International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-4-15.

Ramakrishnan:2012:CPL

[60] Vrinda Ramakrishnan, Bijinu Balakrishnan, Amit Kumar Rai, Bhaskar Narayan, and Prakash M. Halami. Concomitant production of lipase, protease and enterocin by Enterococcus faecium NCIM5363 and Enterococcus durans NCIM5427 isolated from fish processing waste. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1186/2008-6970-4-14.

Bondareva:2012:SLS

[61] Lydia Bondareva, Olga Mogilnaya, and Irina Vlasova. Subcellular localization of ^{241}Am in structural components of submerged macrophyte of the River Yenisei Elodea canadensis. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1186/2008-
6970-4-13.

Kaewudom:2012:EBF

[62] Pimchanok Kaewudom, Soottawat Benjakul, and Kongkarn Kijroongrojana. Effect of bovine and fish gelatin in combination with microbial transglutaminase on gel properties of threadfin bream surimi. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1186/2008-6970-4-12.

Abdel-Tawwab:2012:CEA

[63] Mohsen Abdel-Tawwab. Chronic effect after acute exposure to commercial petroleum fuels on physiological status of Nile tilapia, Oreochromis niloticus (L.). International Aquatic Research, 4(1):??, December 2012. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1186/2008-6970-4-11.

Uddin:2012:BFP

[64] Naim Uddin and Ahmed H. Al-Harbi. Bacterial flora of polycultured common carp (Cyprinus carpio) and African catfish (Clarias gariepinus). International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1186/2008-6970-4-10.

Khangembam:2012:PCT

[65] Bronson Kumar Khangembam, Kameshwar Sharma YVR, and Rina Chakrabarti. Purification and characterization of trypsin from the digestive system of carp Catla catla (Hamilton). International Aquatic Research, 4(1):1–12, December 2012. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-4-9.

Zhang:2012:EAF

[66] Hui Zhang, Qi W. Wei, Chong Li, Xiang S. Wu, and Wen G. Liao. Effects of annual flow characteristics on the freshwater life history of Chinese sturgeon: concern inferred from the number of seaward migrating juveniles. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1186/2008-6970-4-8.

Yarnpakdee:2012:EPC

[67] Suthasinee Yarnpakdee, Soottawat Benjakul, and Hordur G. Kristinsson. Effect of pretreatments on chemical compositions of mince from Nile tilapia (Oreochromis niloticus) and fishy odor development in protein hydrolysate. International Aquatic Research, 4(1):??, December 2012. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1186/2008-6970-4-7.

Nirmal:2012:BPP

[68] Nilesh Prakash Nirmal and Soottawat Benjakul. Biochemical properties of polyphenoloxidase from the cephalothorax of Pacific white shrimp (Litopenaeus vannamei). International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic).

URL https://link.springer.com/article/10.1186/2008-6970-4-6.

Arfat:2012:GCS

[69] Yasir Ali Arfat and Soottawat Benjakul. Gelling characteristics of surimi from yellow stripe trevally (Selaroides leptolepis). International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-4-5.

Christopher:2012:OUT

[70] Johnsamuel Godwin Christopher, Arunachalam Ganesan Murugesan, and Natarajapillai Sukumaran. Optimization of UV treatment to induce haploid androgenesis in the stinging catfish, Heteropneustes fossilis. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1186/2008-6970-4-4.

Abdel-Tawwab:2012:EDP

[71] Mohsen Abdel-Tawwab. Effects of dietary protein levels and rearing density on growth performance and stress response of Nile tilapia, oreochromis niloticus (L.). International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1186/2008-6970-4-3.

Zilli:2012:MMR

[72] Loredana Zilli, Roberta Schiavone, Carlo Storelli, and Sebastiano Vilella. Molecular mechanism regulating axoneme activation in marine fish: a review. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1186/2008-6970-4-2.

Ghomi:2012:E

[73] Mohammad Ghomi, Elena Tricarico, Sajid Maqsood, and Hamid Jamalzadeh. Editorial. International Aquatic Research, 4(1):??, December 2012. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1186/2008-6970-4-1.

Takeungwongtrakul:2013:OSS

[74] Sirima Takeungwongtrakul and Soottawat Benjakul. Oxidative stability of shrimp oil-in-water emulsions as affected by antioxidant incorporation. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1186/2008-6970-5-14.

Khouadja:2013:CVP

[75] Sadok Khouadja, Faouzi Lamari, and Amina Bakhrouf. Characterization of Vibrio parahaemolyticus isolated from farmed sea bass (Dicentrarchus labrax) during disease outbreaks. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-5-13.

Sakai:2013:IPC

[76] Hitomi Sakai, Kozue Edo, Hideyuki Nakagawa, Mitsuko Shinohara, Rie Nishiitsutsuji, and Kiyoshi Ohura. Isolation and partial characterization of a L-rhamnose-binding lectin from the globiferous pedicellariae of the toxopneustid sea urchin, Toxopneustes pileolus. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-5-12.

Abdel-Tawwab:2013:CGB

[77] Mohsen Abdel-Tawwab, Mohamed N. M. Mousaad, Khaled M. Sharafeldin, and Nahla E. M. Ismaiel. Changes in growth and biochemical status of common carp, Cyprinus carpio L. exposed to water-born zinc toxicity for different periods. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1186/2008-6970-5-11.

Lertwittayanon:2013:EDS

[78] Kosol Lertwittayanon, Soottawat Benjakul, Sajid Maqsood, and Angel B. Encarnacion. Effect of different salts on dewatering and properties of yellowtail barracuda surimi. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1186/2008-
6970-5-10.

Vate:2013:AAM

[79] Naveen Kumar Vate and Soottawat Benjakul. Antioxidative activity of melanin-free ink from splendid squid (Loligo formosana). International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1186/2008-6970-5-9.

Mitsoura:2013:PMF

[80] Amalia Mitsoura, Ifigenia Kagalou, Nikolaos Papaioannou, Panagiotis Berillis, Eleni Mente, and Theodoti Papadimitriou. The presence of microcystins in fish Cyprinus carpio tissues: a histopathological study. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1186/2008-6970-5-8.

Rakhi:2013:AHF

[81] Sharmin Ferdewsi Rakhi, Abdul Hakim Mohammad Mohsinul Reza, Mohammad Shafaet Hossen, and Zakir Hossain. Alterations in histopathological features and brain acetylcholinesterase activity in stinging catfish Heteropneustes fossilis exposed to polluted river water. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-5-7.

Ghomi:2013:EAW

[82] Mohammad Reza Ghomi, Mehdi Nikoo, and Mehdi Sohrabnezhad. Effect of alive weight on body composition and fatty acid content of farmed beluga sturgeon (Huso huso). International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1186/2008-
6970-5-6.

Burr:2013:ECP

[83] Gary S. Burr, William R. Wolters, Frederic T. Barrows, and Alan W. Donkin. Evaluation of a canola protein concentrate as a replacement for fishmeal and poultry by-product meal in a commercial production diet for Atlantic salmon (Salmo salar). International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1186/2008-
6970-5-5.

Srivastav:2013:ELP

[84] Ajai K. Srivastav, Rubi Rai, Nobuo Suzuki, Diwakar Mishra, and Sunil K. Srivastav. Effects of lead on the plasma electrolytes of a freshwater fish, Heteropneustes fossilis. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic).

URL https://link.springer.com/article/10.1186/2008-6970-5-4.

Barbosa:2013:MMP

[85] Tatiana S. Barbosa, Carina M. Pereyra, Carla A. Soleiro, Erica O. Dias, Aguida A. Oliveira, Kelly M. Keller, Pedro P. O. Silva, Lilia R. Cavaglieri, and Carlos A. R. Rosa. Mycobiota and mycotoxins present in finished fish feeds from farms in the Rio de Janeiro State, Brazil. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-5-3.

Talbert:2013:ARN

[86] Eric Talbert, Andrea Bodnar, Dorothy M. Morré, and D. James Morré. Age-related NADH oxidase (arNOX) activity is significantly reduced in coelomic fluid of long-lived sea urchins. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1186/2008-6970-5-2.

Manheem:2013:EPC

[87] Kusaimah Manheem, Soottawat Benjakul, Kongkarn Kijroongrojana, Nandhsha Faithong, and Wonnop Visessanguan. Effect of pre-cooking times on enzymes, properties, and melanosis of Pacific white shrimp during refrigerated storage. International Aquatic Research, 5(1):??, December 2013. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic).

URL https://link.springer.com/article/10.1186/2008-6970-5-1.

Srijaya:2014:OCT

[88] Thekkeparambil Chandrabose Srijaya, Padmaja Jayaprasad Pradeep, Anuar Hassan, Anil Chatterji, Faizah Shaharom, and Andrew Jeffs. Oxygen consumption in trilobite larvae of the mangrove horseshoe crab (Carcinoscorpius rotundicauda; Latreille, 1802): effect of temperature, salinity, pH, and light–dark cycle. International Aquatic Research, 6(1):??, April 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-014-0060-z.

Sai-Ut:2014:OGE

[89] Samart Sai-Ut, Soottawat Benjakul, Punnanee Sumpavapol, and Hideki Kishimura. Optimization of gelatinolytic enzyme production by B. amyloliquefaciens sp. H11 through Plackett–Burman design and response surface methodology. International Aquatic Research, 6(1):??, April 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-014-0059-5.

Hassaan:2014:PED

[90] Mohamed S. Hassaan, Ashraf M. A.-S. Goda, Soaad A. Mahmoud, and Safaa I. Tayel. Protective effect of dietary vitamin E against fungicide copperoxychloride stress on Nile tilapia, Oreochromis niloticus (L.) fingerlings. International Aquatic Research, 6(1):??, April 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0058-6.

Mamat:2014:EMF

[91] Nawwar Z. Mamat and Andrea C. Alfaro. Evaluation of microalgal and formulated diets for the culture of the New Zealand pipi clam Paphies australis. International Aquatic Research, 6(1):??, April 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0057-7.

Shoko:2014:CDD

[92] Amon P. Shoko, Samwel M. Limbu, Hillary D. J. Mrosso, and Yunus D. Mgaya. A comparison of diurnal dynamics of water quality parameters in Nile tilapia (Oreochromis niloticus, Linnaeus, 1758) monoculture and polyculture with African sharp tooth catfish (Clarias gariepinus, Burchell, 1822) in earthen ponds. International Aquatic Research, 6(1):??, April 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-014-0056-8.

Shitole:2014:USS

[93] Snehal Subhash Shitole, Amjad Khansaheb Balange, and Shardul Sham Gangan. Use of seaweed (Sargassum tenerrimum) extract as gel enhancer for lesser sardine (Sardinella brachiosoma) surimi. International Aquatic Research, 6(1):??, April 2014. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-014-0055-9.

Aramli:2014:PCE

[94] Mohammad Sadegh Aramli, Mohammad Reza Kalbassi, and Rajab Mohammad Nazari. Protein concentration and enzyme activities of fresh and frozen–thawed Persian sturgeon, Acipenser persicus (Borodin, 1897) semen. International Aquatic Research, 6(1):??, April 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0054-x.

Aketch:2014:ESD

[95] Ben Ochieng’ Aketch, Paul Oyieng’ Ang’ienda, John Odero Radull, and Eliud Nyandago Waindi. Effect of stocking density on the expression of glucose transporter protein 1 and other physiological factors in the Lake Victoria Nile tilapia, Oreochromis niloticus (L.). International Aquatic Research, 6(2):??, June 2014. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-014-0069-3.

Kobayashi:2014:HOD

[96] Yasuhisa Kobayashi, Ryo Nozu, Ryo Horiguchi, and Masaru Nakamura. Histological observation of doublesex-mab 3-related transcription factor 1 (DMRT1) localization in the adult testis of three-spot wrasse. International Aquatic Research, 6(2):??, June 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-014-0068-4.

Moutopoulos:2014:DDF

[97] Dimitrios K. Moutopoulos, Alexis Ramfos, Catherine Moukas, and George Katselis. Description of a daily fishing activity from a small-scale fisherman in Central Greece (Korinthiakos Gulf). International Aquatic Research, 6(2):??, June 2014. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-014-0067-5.

Mohammadzadeh:2014:GPT

[98] Sedigheh Mohammadzadeh, Hosein Ouraji, Fatemeh Hasantabar, and Mohammad K. Khalesi. Growth performance and thyroid hormones of Caspian kutum, Rutilus frissi, juveniles in response to dietary carbohydrate levels. International Aquatic Research, 6(2):??, June 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0066-6.

Hossen:2014:EPF

[99] Md. Shafaet Hossen, A. H. M. Mohsinul Reza, Sharmin Ferdewsi Rakhi, Koretaro Takahashi, and Zakir Hossain. Effects of polyunsaturated fatty acids (PUFAs) on gonadal maturation and spawning of striped gourami, Colisa fasciatus. International Aquatic Research, 6(2):??, June 2014. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1007/s40071-014-0065-7.

Koupaei:2014:MDC

[100] Atoosa Noori Koupaei, Pargol Ghavam Mostafavi, Jalil Fallah Mehrabadi, and Seyed Mohammad Reza Fatemi. Molecular diversity of coral reefassociated zoanthids off Qeshm Island, northern Persian Gulf. International Aquatic Research, 6(2):??, June 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-014-0064-8.

Abro:2014:MIG

[101] Rani Abro, Ali Ata Moazzami, Jan Erik Lindberg, and Torbjörn Lundh. Metabolic insights in Arctic charr (Salvelinus alpinus) fed with zygomycetes and fish meal diets as assessed in liver using nuclear magnetic resonance (NMR) spectroscopy. International Aquatic Research, 6 (2):??, June 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
014-0063-9.

Tongnuanchan:2014:CSP

[102] Phakawat Tongnuanchan, Soottawat Benjakul, and Thummanoon Prodpran. Comparative studies on properties and antioxidative activity of fish skin gelatin films incorporated with essential oils from various sources. International Aquatic Research, 6(2):??, June 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-014-0062-x.

Mohindra:2014:ICR

[103] Vindhya Mohindra, Ratnesh Kumar Tripathi, Akanksha Singh, Rajeev Kumar Singh, and Kuldeep Kumar Lal. Identification of candidate reference genes for quantitative expression analysis by real-time PCR for hypoxic stress in Indian catfish, Clarias batrachus (Linnaeus, 1758). International Aquatic Research, 6(2):??, June 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-014-0061-y.

Baron:2014:CCP

[104] Ron Baron, Alexander Davie, Alec Gaines, Darren Grant, Oya Okay, and Emin Ozsoy. Continuous cultures of phytoplankton. International Aquatic Research, 6(3):95–111, September 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-014-0070-x.

Chatterjee:2014:HEZ

[105] Nilanjana Chatterjee, Baibaswata Bhattacharjee, and Chung-Hsin Lu. Hazardous effect of ZnS nanoparticles on the feeding behaviour, growth and maturation process of the Asian striped catfish, Mystus vittatus (Bloch, 1794). International Aquatic Research, 6(3):113–125, September 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-014-0071-9.

Glamuzina:2014:OIW

[106] Branko Glamuzina, Ana Pešić, Aleksandar Joksimović, Luka Glamuzina, Sanja Matić-Skoko, Alexis Conides, Dimitris Klaoudatos, and Paraskevi Zacharaki. Observations on the increase of wild gilthead seabream, Sparus aurata abundance, in the eastern Adriatic Sea: problems and opportunities. International Aquatic Research, 6(3):127–134, September 2014. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1007/s40071-014-0073-7.

Kathiravan:2014:DDB

[107] V. Kathiravan and K. K. Krishnani. Diversity of denitrifying bacteria in the greenwater system of coastal aquaculture. International Aquatic Research, 6(3):135–145, September 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-014-0074-6.

Sarasiab:2014:MMM

[108] Abdolah Raeisi Sarasiab, Mehdi Hosseini, and Fatemeh Tadi Beni. Mercury and methyl mercury concentration in sediment, benthic, Barbus grypus and pelagic, Barbus esocinus fish species, from Musa estuary, Iran. International Aquatic Research, 6(3):147–153, September 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0075-5.

Pongsetkul:2014:CCP

[109] Jaksuma Pongsetkul, Soottawat Benjakul, Punnanee Sampavapol, Kazufumi Osako, and Nandhsha Faithong. Chemical composition and physical properties of salted shrimp paste (Kapi) produced in Thailand. International Aquatic Research, 6(3):155–166, September 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0076-4.

Dash:2014:IEG

[110] Pujarini Dash, Swati Yadav, and P. K. Sahoo. Immunoadjuvant effect of garlic (Allium sativum)–mineral oil suspension on immunity and resistance to Aeromonas hydrophila infection in rohu, Labeo rohita. International Aquatic Research, 6(3):167–173, September 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0072-8.

Valenti:2014:TGS

[111] Wagner C. Valenti. Toward a good scientific writing. International Aquatic Research, 6(4):175–176, December 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-014-0080-8.

Yurimoto:2014:MME

[112] Tatsuya Yurimoto, Faizul Mohd Kassim, Reiko Fuseya, and Alias Man. Mass mortality event of the blood cockle, Anadara granosa, in aquaculture ground along Selangor coast, Peninsular Malaysia. International Aquatic Research, 6(4):177–186, December 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-014-0077-3.

Arimoro:2014:FSL

[113] Francis O. Arimoro, Robert B. Ikomi, Francis O. Nwadukwe, Ovie D. Eruotor, and Augustine O. Edegbene. Fluctuating salinity levels and an increasing pollution gradient on fish community structure and trophic levels in a small creek in the Niger delta, Nigeria. International Aquatic Research, 6(4):187–202, December 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-014-0078-2.

Flakova:2014:WQG

[114] Renáta Fľaková, Milan Seman, Hana Drahovská, Zlatica Ženišová, and Alexandra Ďuričková. The water quality of the Danube River and gravel pits in the Bratislava area (Slovakia). International Aquatic Research, 6(4):203–210, December 2014. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-014-0079-1.

Patel:2014:PPI

[115] Chanda Patel, Neelima Raipuria, Monica Agrawal, and Manish K. Agrawal. Is production of protease inhibitors from cyanobacteria nutrient dependent? Comparison of protease inhibitory activities in three species of Oscillatoria isolated from Central India. International Aquatic Research, 6(4):211–220, December 2014. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-014-0081-7.

Shakir:2014:EMC

[116] Hafiz Abdullah Shakir, Javed Iqbal Qazi, and Abdul Shakoor Chaudhry. Examining muscles of Cirrhinus mrigala for biochemical parameters as a bio-indicator of water pollution by municipal and industrial effluents into River Ravi, Pakistan. International Aquatic Research, 6(4):221–228, December 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
014-0082-6.

Agbohessi:2014:EAP

[117] Tachégnon Prudencio Agbohessi, Ibrahim Imorou Toko, Innocent N’tcha, Florian Geay, S. N. M. Mandiki, and Patrick Kestemont. Exposure to agricultural pesticides impairs growth, feed utilization and energy budget in African catfish Clarias gariepinus (Burchell, 1822) fingerlings. International Aquatic Research, 6(4):229–243, December 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0083-5.

Jiwyam:2014:SDD

[118] Wirat Jiwyam and Nudtha Nithikulworawong. Stocking densitydependent growth and survival of Asian red-tailed catfish (Hemibagrus wyckioides) fries: early nursing in cages. International Aquatic Research, 6(4):245–250, December 2014. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-014-0084-4.

Soberon:2014:HPC

[119] Luis Soberon, Patrick Mathews, and Antonio Malherios. Hematological parameters of Colossoma macropomum naturally parasitized by Anacanthorus spathulatus (Monogenea: Dactylogiridae) in fish farm in the Peruvian Amazon. International Aquatic Research, 6(4):251–255, December 2014. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-014-0087-1.

Rattaya:2015:EAA

[120] Saowapa Rattaya, Soottawat Benjakul, and Thummanoon Prodpran. Extraction, antioxidative, and antimicrobial activities of brown seaweed extracts, Turbinaria ornata and Sargassum polycystum, grown in Thailand. International Aquatic Research, 7(1):1–16, March 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1007/s40071-014-0085-3.

Vaseem:2015:EPT

[121] Huma Vaseem and T. K. Banerjee. Efficacy of phytoremediation technology in decontaminating the toxic effluent released during recovery of metals from polymetallic sea nodules. International Aquatic Research, 7(1):17–26, March 2015. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-014-0089-z.

Mohammad:2015:OEH

[122] Abdul Wahab Mohammad, Abd Ghani Kumar, and Roseliza Kadir Basha. Optimization of enzymatic hydrolysis of tilapia (Oreochromis spp.) scale gelatine. International Aquatic Research, 7(1):27–39, March 2015. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1007/s40071-014-0090-6.

Siddik:2015:PAG

[123] Muhammad Abu Bakar Siddik and Nguyen Thi Ngoc Anh. Preliminary assessment of the gut weed Ulva intestinalis as food for herbivorous fish. International Aquatic Research, 7(1):41–46, March 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1007/s40071-014-0091-5.

Tsutsui:2015:HTC

[124] Isao Tsutsui, Tatsuo Miyoshi, Dusit Aue-umneoy, Jaruwan Songphatkaew, Chonlada Meeanan, Sirimas Klomkling, Halethichanok Sukchai, Piyarat Pinphoo, Iori Yamaguchi, Monthon Ganmanee, Yukio Maeno, and Kaoru Hamano. High tolerance of Chaetomorpha sp. to salinity and water temperature enables survival and growth in stagnant waters of central Thailand. International Aquatic Research, 7(1):47–62, March 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0092-4.

Limbu:2015:EIS

[125] Samwel Mchele Limbu, Amon Paul Shoko, Hieromin Amon Lamtane, Evantuce Dominick Shirima, Mary Alphonce Kishe-Machumu, Huruma Fredrick Mgana, and Yunus Daud Mgaya. Effect of initial stocking size of the predatory African sharptooth catfish (Clarias gariepinus) on recruits, growth performance, survival and yield of mixed-sex Nile tilapia (Oreochromis niloticus) in concrete tank culture system. International Aquatic Research, 7(1):63–73, March 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-014-0093-3.

Caipang:2015:CLE

[126] Christopher Marlowe A. Caipang, Fernand F. Fagutao, Effrosyni Fatira, Carlo C. Lazado, and Michail Pavlidis. Cortisol levels and expression of selected stress- and apoptosis-related genes in the embryos of Atlantic cod, Gadus morhua following short-term exposure to air. International Aquatic Research, 7(1):75–84, March 2015. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-015-0094-x.

Joseph:2015:VSP

[127] Toms C. Joseph, L. Anbu Rajan, Roswin James, K. V. Lalitha, and P. K. Surendran. Variations of structural protein sequences among geographical isolates of white spot syndrome virus. International Aquatic Research, 7(1):85–91, March 2015. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-014-0086-2.

Sicuro:2015:FBR

[128] Benedetto Sicuro. Freshwater bivalves rearing: a brief overview. International Aquatic Research, 7(2):93–100, June 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-015-0098-6.

Karnjanapratum:2015:AGH

[129] Supatra Karnjanapratum and Soottawat Benjakul. Antioxidative gelatin hydrolysate from unicorn leatherjacket skin as affected by prior autolysis. International Aquatic Research, 7(2):101–114, June 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-014-0088-0.

Jamali:2015:EGS

[130] Hadi Jamali, Nasrollah Ahmadifard, and Daruosh Abdollahi. Evaluation of growth, survival and body composition of larval white shrimp (Litopenaeus vannamei) fed the combination of three types of algae. International Aquatic Research, 7(2):115–122, June 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-015-0095-9.

Kotova:2015:MDB

[131] Anastasiya V. Kotova, Oleg A. Kovtun, Olga I. Podgornaya, and Leonid S. Adonin. Mesoglein differences between two jellyfish populations of the genus Aurelia. International Aquatic Research, 7(2):123–131, June 2015.

CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-015-0096-8.

Yurimoto:2015:BTD

[132] Tatsuya Yurimoto, Dusit Aue-umneoy, Chonlada Meeanan, and Isao Tsutsui. Bloom of the two dinoflagellates Ceratium furca and Diplopsalis lenticula in a mangrove estuary of Thailand. International Aquatic Research, 7(2):133–141, June 2015. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-015-0099-5.

Nakagawa:2015:INL

[133] Hideyuki Nakagawa, Kuniko Nagasaka, Hitomi Sakai, Kozue Edo, Mitsuko Shinohara, and Kiyoshi Ohura. Isolation of a novel lectin from the dorsal spines of the devil stinger, Inimicus japonicus. International Aquatic Research, 7(2):143–150, June 2015. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-015-0101-2.

Tavakoli-Kolour:2015:OGA

[134] Pavriz Tavakoli-Kolour, Javid Kavousi, and Hamid Rezai. Outbreak of growth anomalies in coral communities of Qeshm Island, Persian Gulf. International Aquatic Research, 7(2):151–156, June 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1007/s40071-015-0100-3.

Noor:2015:DVG

[135] Rashed Noor, Md. Faqrul Hasan, Md. Sakil Munna, and M. Majibur Rahman. Demonstration of virulent genes within Listeria and Klebsiella isolates contaminating the export quality frozen shrimps. International Aquatic Research, 7(2):157–161, June 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-015-0097-7.

Prusty:2015:SPT

[136] A. K. Prusty, D. K. Meena, S. Mohapatra, P. Panikkar, P. Das, S. K. Gupta, and B. K. Behera. Synthetic pyrethroids (Type II) and freshwater fish culture: Perils and mitigations. International Aquatic Research, 7 (3):163–191, September 2015. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-015-0106-x. See erratum [151].

Tsutsui:2015:CCC

[137] Isao Tsutsui, Jaruwan Songphatkaew, Chonlada Meeanan, Dusit Aueumneoy, Halethichanok Sukchai, Piyarat Pinphoo, Sirimas Klomkling, Monthon Ganmanee, Hiroyuki Sudo, and Kaoru Hamano. Co-culture with Chaetomorpha sp. enhanced growth performance and reduced feed conversion ratio of the giant tiger prawn, Penaeus monodon. International Aquatic Research, 7(3):193–199, September 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-015-0103-0.

Karami:2015:CPF

[138] Ali Karami, Peyman Eghtesadi Araghi, Mohd Arif Syed, and Scott P. Wilson. Chromosome preparation in fish: effects of fish species and larval age. International Aquatic Research, 7(3):201–210, September 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-015-0104-z.

Lee:2015:USI

[139] Jaeyong Lee, Jae-Ok Kim, Jeffrey S. Owen, and Bomchul Kim. Using stable isotopes to investigate movement of fish in rice paddy fields. International Aquatic Research, 7(3):211–220, September 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-015-0105-y.

Winckelmann:2015:OPC

[140] Dominik Winckelmann, Franziska Bleeke, Bimiya Thomas, Clemens Elle, and Gerd Klöck. Open pond cultures of indigenous algae grown on nonarable land in an arid desert using wastewater. International Aquatic Research, 7(3):221–233, September 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-015-0107-9.

Bleeke:2015:OFM

[141] Franziska Bleeke, Malgorzata Milas, Dominik Winckelmann, and Gerd Klöck. Optimization of freshwater microalgal biomass harvest using polymeric flocculants. International Aquatic Research, 7(3):235–244, September 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
015-0108-8.

Chow:2015:UPE

[142] Seinen Chow, Takashi Yanagimoto, and Yoji Nakamura. Universal primers for exon-priming intron-crossing (EPIC) PCR on ribosomal protein genes in marine animals. International Aquatic Research, 7(3):245–250, September 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
015-0109-7.

Uc-Peraza:2015:ATR

[143] R. G. Uc-Peraza and V. H. Delgado-Blas. Acute toxicity and risk assessment of three commercial detergents using the polychaete Capitella sp. C from Chetumal Bay, Quintana Roo, Mexico. International Aquatic Research, 7(4):251–261, December 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-015-0112-z.

Ovissipour:2015:EWS

[144] Mahmoudreza Ovissipour, Hamzah M. Al-Qadiri, Xiaonan Lu, Yaxi Hu, Carolyn F. Ross, Joel P. Van Eenennaam, Serge I. Doroshov, and Barbara Rasco. The effect of white sturgeon (Acipenser transmontanus) ovarian fat deposition on caviar yield and nutritional quality: introducing image processing method for sturgeon ovary fat determination. International Aquatic Research, 7(4):263–272, December 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-015-0111-0.

Pandey:2015:ADC

[145] Jitendra Pandey, Anand V. Singh, Rachna Singh, Pooja Kaushik, and Usha Pandey. Atmospheric deposition coupled terrestrial export of organic carbon in Ganga River (India): linking cross-domain carbon transfer to river DOC. International Aquatic Research, 7(4):273–285, December 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-015-0110-1.

Sriyasak:2015:EWS

[146] Patcharawalai Sriyasak, Chanagun Chitmanat, Niwooti Whangchai, Jongkon Promya, and Louis Lebel. Effect of water de-stratification on dissolved oxygen and ammonia in tilapia ponds in Northern Thailand. International Aquatic Research, 7(4):287–299, December 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-015-0113-y.

Piontkovski:2015:MGO

[147] Sergey A. Piontkovski, Asyla Al-Mawali, Ahlam Al-Kharusi, WardalMuna Al-Manthri, Sharon Smith, and Elena Popova. Mesozooplankton of the Omani shelf: taxonomy, seasonality, and spatial distribution. International Aquatic Research, 7(4):301–314, December 2015. CODEN ????

ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1007/s40071-015-0114-x.

Khaliullina:2015:SDP

[148] L. Yu. Khaliullina and G. V. Demina. Seasonal dynamics of phytoplankton communities residing in different types of shallow waters in the Kuibyshev Reservoir (Russia). International Aquatic Research, 7(4):315–328, December 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
015-0116-8.

Caipang:2015:VSP

[149] Christopher Marlowe A. Caipang, Han Xuan Choo, Zemin Bai, Huilin Huang, and Clara M. Lay-yag. Viability of sweet potato flour as carbon source for the production of biofloc in freshwater culture of tilapia, Oreochromis sp. International Aquatic Research, 7(4):329–336, December 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-015-0117-7.

Ansari:2015:SDT

[150] Abid Ali Ansari, Sulaiman Al Ghanim, Subrata Trivedi, Hasibur Rehman, Zahid Khorshid Abbas, and Shalini Saggu. Seasonal dynamics in the trophic status of water, floral and faunal density along some selected coastal areas of the Red Sea, Tabuk, Saudi Arabia. International Aquatic Research, 7(4):337–348, December 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-015-0118-6.

Prusty:2015:ESP

[151] A. K. Prusty, D. K. Meena, S. Mohapatra, P. Panikkar, P. Das, S. K. Gupta, and B. K. Behera. Erratum to: Synthetic pyrethroids (Type II) and freshwater fish culture: Perils and mitigations. International Aquatic Research, 7(4):349–351, December 2015. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-015-0115-9. See [136].

Joyce:2016:USI

[152] Melissa K. Joyce and Igor Pirozzi. Using stable isotope analysis to determine the contribution of naturally occurring pond biota and supplementary feed to the diet of farmed Australian freshwater crayfish, redclaw (Cherax quadricarinatus). International Aquatic Research, 8(1):1– 13, March 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
015-0119-5.

Takeungwongtrakul:2016:ADL

[153] Sirima Takeungwongtrakul and Soottawat Benjakul. Astaxanthin degradation and lipid oxidation of Pacific white shrimp oil: kinetics study and stability as affected by storage conditions. International Aquatic Research, 8(1):15–27, March 2016. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-015-0120-z.

Chow:2016:UPP

[154] Seinen Chow and Takashi Yanagimoto. Universal PCR primers for ribosomal protein gene introns of fish. International Aquatic Research, 8(1):29– 36, March 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
016-0122-5.

Dias-Neto:2016:TSF

[155] José Dias-Neto, Gustavo Moraes Ramos Valladão, Pedro Henrique de Oliveira Viadanna, and Fabiana Pilarski. Three strategic feeding during hatchering of Nile tilapia: effects on organs integrity, parasitism and performance parameters. International Aquatic Research, 8(1):37–48, March 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-016-0123-4.

Piontkovski:2016:DCG

[156] Sergey A. Piontkovski and Bastien Y. Queste. Decadal changes of the Western Arabian Sea ecosystem. International Aquatic Research, 8(1):49– 64, March 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
016-0124-3.

DAmbrosio:2016:ZDP

[157] D. Sabina D’Ambrosio, María C. Claps, and Adriana García. Zooplankton diversity of a protected and vulnerable wetland system in southern South America (Llancanelo area, Argentina). International Aquatic Research, 8(1):65–80, March 2016. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-016-0125-2.

Winckelmann:2016:DWA

[158] Dominik Winckelmann, Franziska Bleeke, Peter Bergmann, Clemens Elle, and Gerd Klöck. Detection of weed algae in open pond cultures of Cyanobacterium aponinum using PAM. International Aquatic Research, 8(1):81–90, March 2016. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-016-0126-1.

Angel:2016:EPS

[159] Alejandro Angel, Irma Vila, and Venecia Herrera. Extremophiles: photosynthetic systems in a high-altitude saline basin (Altiplano, Chile). International Aquatic Research, 8(2):91–108, June 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-016-0121-6.

Arechavala-Lopez:2016:VCF

[160] Pablo Arechavala-Lopez, Marius Berg, Ingebrigt Uglem, Pål-Arne Bjørn, and Bengt Finstad. Variations in coastal fish species composition captured by traps in Romsdalsfjord, Western Norway. International Aquatic Research, 8(2):109–119, June 2016. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-016-0127-0.

Simide:2016:AAP

[161] Rémy Simide, Simone Richard, Nathalie Prévot-D’Alvise, Thomas Miard, and Sandrine Gaillard. Assessment of the accuracy of physiological blood indicators for the evaluation of stress, health status and welfare in Siberian sturgeon (Acipenser baerii) subject to chronic heat stress and dietary supplementation. International Aquatic Research, 8(2):121–135, June 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-016-0128-z.

Mansourkiaei:2016:PRS

[162] Ana Mansourkiaei, Pargol Ghavam Mostafavi, Seyed Mohammad Reza Fatemi, Farhad Kaymaram, and Ali Nazemi. Phylogenetic relationships of Scomberomorus commerson using sequence analysis of the mtDNA Dloop region in the Persian Gulf, Oman Sea and Arabian Sea. International Aquatic Research, 8(2):137–148, June 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-016-0129-y.

Tanyaros:2016:EFC

[163] Suwat Tanyaros, Woraporn Tarangkoon, and Tossaporn Klomkleing. Evaluation of flocculated concentrates from intensive shrimp pond water as a substitute for microalgal concentrates in the nursery culture of juvenile oyster (Crassostrea belcheri). International Aquatic Research, 8(2):149– 160, June 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
016-0130-5.

Lim:2016:EPB

[164] Leong-Seng Lim, Sian-Kang Jason Lai, Annita Seok-Kian Yong, Rossita Shapawi, and Gunzo Kawamura. Evaluation on the potential of betaine, taurine, nucleotide and nucleoside as feeding stimulant for juvenile marble goby Oxyeleotris marmoratus through behavioural assays. International Aquatic Research, 8(2):161–167, June 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-016-0131-4.

Okamoto:2016:SIA

[165] Yuki Okamoto, Nozomu Muto, Koetsu Kon, Kazuya Watanabe, Takashi Yoshikawa, Jintana Salaenoi, and Satoshi Ishikawa. Stable isotope analysis suggests the existence of multiple populations of streaked spinefoot (Siganus javus L.) in Bandon Bay, Southern Thailand. International Aquatic Research, 8(2):169–178, June 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-016-0132-3.

Barakat:2016:PES

[166] Khouloud M. Barakat, Heba S. El-Sayed, and Yousry M. Gohar. Protective effect of squilla chitosan–silver nanoparticles for Dicentrarchus labrax larvae infected with Vibrio anguillarum. International Aquatic Research, 8(2):179–189, June 2016. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-016-0133-2.

Schaefer:2016:IKL

[167] F. J. Schaefer and S. Wuertz. Insights into kisspeptin- and leptin-signalling on GnRH mRNA expression in hypothalamic organ cultures of immature pikeperch Sander lucioperca. International Aquatic Research, 8(2):191– 196, June 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
016-0134-1.

Abdel-Tawwab:2016:GBV

[168] Mohsen Abdel-Tawwab, Gamal O. El-Sayed, and Sherien H. Shady. Growth, biochemical variables, and zinc bioaccumulation in Nile tilapia, Oreochromis niloticus (L.) as affected by water-born zinc toxicity and exposure period. International Aquatic Research, 8(3):197–206, September 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-016-0135-0.

Stabnikov:2016:SSU

[169] Viktor Stabnikov, Volodymyr Ivanov, and Jian Chu. Sealing of sand using spraying and percolating biogrouts for the construction of model aquaculture pond in arid desert. International Aquatic Research, 8(3):207–216, September 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
016-0136-z.

Kung:2016:HFF

[170] Hsien-Feng Kung, Yi-Chen Lee, Ya-Ling Huang, Chun-Yung Huang, Chiu-Chu Hwang, and Yung-Hsiang Tsai. Histamine formation in flying fish contaminated with Staphylococcus xylosus. International Aquatic Research, 8(3):217–226, September 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-016-0137-y.

Nilsuwan:2016:QCS

[171] Krisana Nilsuwan, Soottawat Benjakul, and Thummanoon Prodpran. Quality changes of shrimp cracker covered with fish gelatin film without and with palm oil incorporated during storage. International Aquatic Research, 8(3):227–238, September 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-016-0138-x.

Kim:2016:EDP

[172] Kang-Woong Kim, Mohammad Moniruzzaman, Kyoung-Duck Kim, Hyon Sob Han, Hyeonho Yun, Seunghan Lee, and Sungchul C. Bai. Effects of dietary protein levels on growth performance and body composition of juvenile parrot fish, Oplegnathus fasciatus. International Aquatic Research, 8(3):239–245, September 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-016-0139-9.

Soltanian:2016:EEH

[173] Siyavash Soltanian and Mohammad Saeid Fereidouni. Effect of Henna (Lawsonia inermis) extract on the immunity and survival of common carp, Cyprinus carpio infected with Aeromonas hydrophila. International Aquatic Research, 8(3):247–261, September 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-016-0141-2.

Arechavala-Lopez:2016:UTE

[174] P. Arechavala-Lopez, M. Milošević-González, and P. Sanchez-Jerez. Using trace elements in otoliths to discriminate between wild and farmed European sea bass (Dicentrarchus labrax L.) and gilthead sea bream (Sparus aurata L.). International Aquatic Research, 8(3):263–273, September 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-016-0142-1.

Theodorou:2016:OBE

[175] John A. Theodorou, Costas Perdikaris, and Barbara Venou. Origin of broodstock and effects on the deformities of gilthead sea bream (Sparus aurata L. 1758) in a Mediterranean commercial hatchery. International Aquatic Research, 8(3):275–282, September 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-016-0140-3.

Dalzochio:2016:UBA

[176] Thaís Dalzochio, Gabriela Zimmermann Prado Rodrigues, Ismael Evandro Petry, Günther Gehlen, and Luciano Basso da Silva. The use of biomarkers to assess the health of aquatic ecosystems in Brazil: a review. International Aquatic Research, 8(4):283–298, December 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-016-0147-9.

Kazemi:2016:AHE

[177] Soodabeh Kazemi, Behrooz Heidari, and Mehdi Rassa. Antibacterial and hemolytic effects of aqueous and organic extracts from different tissues of sea urchin Echinometra mathaei on pathogenic streptococci. International Aquatic Research, 8(4):299–308, December 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-016-0143-0.

Da:2016:GPF

[178] Chau Thi Da, Torbjörn Lundh, Jan Erik Lindberg, and Håkan Berg. Growth performance, feed utilisation and biological indices of tra catfish (Pangasianodon hypophthalmus) cultured in net cages in pond fed diets based on locally available feed resources. International Aquatic Research, 8(4):309–321, December 2016. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-016-0144-z.

Hua:2016:DPC

[179] Nhan Thai Hua and Harry Ako. Dietary protein and carbohydrate requirement of juvenile Hawaiian limpet (Cellana sandwicensis Pease, 1861) fed practical diet. International Aquatic Research, 8(4):323–332, December 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-016-0145-y.

Djissou:2016:CRF

[180] Arnauld S. M. Djissou, Dogbè C. Adjahouinou, Shunsuke Koshio, and Emile D. Fiogbe. Complete replacement of fish meal by other animal protein sources on growth performance of Clarias gariepinus fingerlings. International Aquatic Research, 8(4):333–341, December 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-016-0146-x.

Medeiros-Silva:2016:PGA

[181] Renata Maria Medeiros-Silva and Humber Agrelli Andrade. Proportions of Atlantic sailfish (Istiophorus platypterus) and spearfish (Tetrapturus pfluegeri) caught in tuna longline fisheries. International Aquatic Research, 8(4):343–352, December 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-016-0149-7.

Sicuro:2016:UFD

[182] Benedetto Sicuro. The use of formaldehyde for the disinfection of maternally incubated eggs of noble crayfish (Astacus astacus). International Aquatic Research, 8(4):353–359, December 2016. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-016-0148-8.

Crisp:2017:ETS

[183] Jason A. Crisp, Gavin J. Partridge, Frances M. L. D’Souza, James R. Tweedley, and Navid R. Moheimani. Effects of temperature and salinity on larval survival and development of the western school prawn Metapenaeus dalli. International Aquatic Research, 9(1):1–10, March 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-016-0151-0.

Yun:2017:EDS

[184] Hyeonho Yun, Erfan Shahkar, Ali Hamidoghli, Seunghan Lee, Seonghun Won, and Sungchul C. Bai. Evaluation of dietary soybean meal as fish meal replacer for juvenile whiteleg shrimp, Litopenaeus vannamei reared in biofloc system. International Aquatic Research, 9(1):11–24, March 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0152-7.

Moazenzadeh:2017:DZR

[185] Kadusa Moazenzadeh, Houman Rajabi Islami, Abasali Zamini, and Mehdi Soltani. Dietary zinc requirement of Siberian sturgeon (Acipenser baerii, Brandt 1869) juveniles, based on the growth performance and blood parameters. International Aquatic Research, 9(1):25–35, March 2017. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1007/s40071-017-0153-6.

Seki:2017:AMD

[186] Hiroko Seki, Kaori Nakazato, and Naoko Hamada-Sato. Adenosine monophosphate degradation and inosinic acid accumulation in the shrimp Penaeus japonicus. International Aquatic Research, 9(1):37–52, March 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0154-5.

Chow:2017:LSE

[187] Seinen Chow, Nobuaki Suzuki, Yoji Nakamura, Motoshige Yasuike, Kenji Saitoh, and Takashi Yanagimoto. Locus-specific EPIC–PCR primers for four distinct calmodulin genes of the Pacific bluefin tuna Thunnus orientalis (Temminck & Schlegel, 1844). International Aquatic Research, 9(1):53–59, March 2017. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-017-0155-4.

Kusmita:2017:CCP

[188] Lia Kusmita, Erlita Verdia Mutiara, Handung Nuryadi, Petrick Ariska Pratama, Awang Surya Wiguna, and Ocky Karna Radjasa. Characterization of carotenoid pigments from bacterial symbionts of soft-coral Sarcophyton sp. from North Java Sea. International Aquatic Research, 9(1):61– 69, March 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
017-0157-2.

Wang:2017:DPR

[189] Jingjing Wang, Wei Zhao, Yajie Liang, Yifan Han, Xingbiao Wang, Xiaoxia Zhang, Xiaomei Ma, Demin Zhang, and Zhiyong Huang. Development of a prototype 16S rRNA gene-based microarray for monitoring planktonic actinobacteria in shrimp ponds. International Aquatic Research, 9(1):71–80, March 2017. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-017-0158-1.

Rios-Villamizar:2017:SWQ

[190] Eduardo Antonio Ríos-Villamizar, Maria T. F. Piedade, Wolfgang J. Junk, and Andréa Viviana Waichman. Surface water quality and deforestation of the Purus river basin, Brazilian Amazon. International Aquatic Research, 9(1):81–88, March 2017. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-016-0150-1.

Azpeitia:2017:SNV

[191] K. Azpeitia, Y. Ríos, I. Garcia, J. Pagaldai, and D. Mendiola. A sensory and nutritional validation of open ocean mussels (Mytilus galloprovincialis Lmk.) cultured in SE Bay of Biscay (Basque Country) compared to their commercial counterparts from Galician Rías (Spain). International Aquatic Research, 9(2):89–106, June 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-017-0159-0.

Sulaeman:2017:YUG

[192] Sulaeman and Ravi Fotedar. Yolk utilization and growth during the early larval life of the silver perch, Bidyanus bidyanus (Mitchell, 1838). International Aquatic Research, 9(2):107–116, June 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-017-0160-7.

Majumder:2017:CIS

[193] Rajib Majumder and Anilava Kaviraj. Cypermethrin induced stress and changes in growth of freshwater fish Oreochromis niloticus. International Aquatic Research, 9(2):117–128, June 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-017-0161-6.

Anvo:2017:PUC

[194] Morgane Paul Magouana Anvo, Rokyatou Sissao, Benié Rose Danielle Aboua, Chantal Yvette Zoungrana-Kaboré, Athanase Kraidy Otchoumou, Essetchi Paul Kouamelan, and Aboubacar Toguyéni. Preliminary use of cashew kernel oil in Clarias gariepinus fingerlings diet: comparison with fish oil and palm oil. International Aquatic Research, 9(2):129–139, June 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0162-5.

Ouyang:2017:CNP

[195] Xu Ouyang, Jiancao Gao, Bojian Chen, Zaizhao Wang, Hong Ji, and Martin Plath. Characterizing a novel predator–prey relationship between native Diplonychus esakii (Heteroptera: Belostomatidae) and invasive Gambusia affinis (Teleostei: Poeciliidae) in central China. International Aquatic Research, 9(2):141–151, June 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-017-0163-4.

Abdel-Tawwab:2017:CSA

[196] Mohsen Abdel-Tawwab, Gamal O. El-Sayed, and Sherien H. H. Shady. Capability of some agricultural wastes for removing some heavy metals from polluted water stocked in combination with Nile tilapia, Oreochromis niloticus (L.). International Aquatic Research, 9(2):153–160, June 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0166-1.

Yu:2017:AFG

[197] Xinxiu Yu, Changwen Wu, and Hans Magnus Gjøen. Artificial fertilization and generating families for a selective breeding programme of large yellow croaker (Larimichthys crocea). International Aquatic Research, 9(2):161– 167, June 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
017-0164-3.

Bhaduri:2017:STR

[198] Ritindra N. Bhaduri and Maya Wolf. Saccate thallus of the red alga Halosaccion glandiforme harbor diverse invertebrate fauna. International Aquatic Research, 9(2):169–176, June 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-017-0165-2.

Maizatul:2017:OUM

[199] A. Y. Maizatul, Radin Maya Saphira Radin Mohamed, Adel A. Al-Gheethi, and M. K. Amir Hashim. An overview of the utilisation of microalgae biomass derived from nutrient recycling of wet market wastewater and slaughterhouse wastewater. International Aquatic Research, 9 (3):177–193, September 2017. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-017-0168-z.

Wu:2017:MSH

[200] Hao Wu, Jianlong Zhang, Yongxing He, Junfang Zhou, Jizhou Yan, and Min Jiang. A metabolic study in hepatopancreas of Litopenaeus vannamei response to white spot syndrome virus. International Aquatic Research, 9(3):195–201, September 2017. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-017-0167-0.

Pikle:2017:FDI

[201] R. P. Pikle, R. M. Jatiger, and C. B. Ganesh. Food-deprivation-induced suppression of pituitary–testicular-axis in the tilapia Oreochromis mossambicus. International Aquatic Research, 9(3):203–213, September 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0169-y.

Yasemi:2017:DGM

[202] Mehran Yasemi, Alireza Nazari Bajgan, and Mehran Parsa. Determining the growth and mortality parameters of longtail tuna (Thunnus tonggol Bleeker, 1851) using length frequency data in coastal waters of the northern Persian Gulf and Oman Sea, Iran. International Aquatic Research, 9(3):215–224, September 2017. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-017-0170-5.

Trigueiro:2017:CTC

[203] Thiago Gaban Trigueiro, Dinaelza Castelo Pereira, Aline P. Martins, Pio Colepicolo, and Eliane Marinho-Soriano. Cultivation of three color strains of Gracilaria domingensis in an integrated organic system. International Aquatic Research, 9(3):225–233, September 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-017-0171-4.

Rahman:2017:TCB

[204] Norazira Abdu Rahman, Helena Khatoon, Norhayati Yusuf, Sanjoy Banerjee, Noor Azilah Haris, Fathurrahman Lananan, and Katayama Tomoyo. Tetraselmis chuii biomass as a potential feed additive to improve survival and oxidative stress status of Pacific white-leg shrimp Litopenaeus vannamei postlarvae. International Aquatic Research, 9(3):235–247, September 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
017-0173-2.

Morozov:2017:AVH

[205] Alexey A. Morozov, Grigorii M. Chuiko, and Victoria V. Yurchenko. Annual variations in hepatic antioxidant defenses and lipid peroxidation in a temperate fish, common bream Abramis brama (L.). International Aquatic Research, 9(3):249–257, September 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-017-0176-z.

Tweedley:2017:BSF

[206] James R. Tweedley, Chris S. Hallett, and Stephen J. Beatty. Baseline survey of the fish fauna of a highly eutrophic estuary and evidence for its colonisation by goldfish (Carassius auratus). International Aquatic Research, 9(3):259–270, September 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-017-0174-1.

Sohrabnezhad:2017:EDS

[207] M. Sohrabnezhad, M. Sudagar, and M. Mazandarani. Effect of dietary soybean meal and multienzyme on intestine histology of beluga sturgeon (Huso huso). International Aquatic Research, 9(4):271–280, December 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0172-3.

Song:2017:PCS

[208] Weiwei Song, Chunlin Wang, Ronghua Li, and Changkao Mu. Peroxiredoxin 6 from cuttlefish (Sepiella maindroni): molecular characterisation during development and the response to thermal stress. International Aquatic Research, 9(4):281–292, December 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-017-0175-0.

Ng:2017:DSC

[209] Wing-Keong Ng, Chia-Ling Lim, Nicholas Romano, and Beng-Chu Kua. Dietary short-chain organic acids enhanced resistance to bacterial infection and hepatopancreatic structural integrity of the giant freshwater prawn, Macrobrachium rosenbergii. International Aquatic Research, 9 (4):293–302, December 2017. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-017-0177-y.

Katya:2017:EIL

[210] Kumar Katya, M. Z. S. Borsra, Dev Ganesan, Giva Kuppusamy, Max Herriman, Andrew Salter, and Sayed Azam Ali. Efficacy of insect larval meal to replace fish meal in juvenile barramundi, Lates calcarifer reared in freshwater. International Aquatic Research, 9(4):303–312, December 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0178-x.

Dogan:2017:DAA

[211] Güntekin Doğan and Ömer Osman Ertan. Determination of amino acid and fatty acid composition of goldband goatfish [ Upeneus moluccensis (Bleeker, 1855)] fishing from the Gulf of Antalya (Turkey). International Aquatic Research, 9(4):313–327, December 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-017-0179-9.

Azpeitia:2017:VRC

[212] K. Azpeitia, M. Ortiz-Zarragoitia, M. Revilla, and D. Mendiola. Variability of the reproductive cycle in estuarine and coastal populations of the mussel Mytilus galloprovincialis Lmk. from the SE Bay of Biscay (Basque Country). International Aquatic Research, 9(4):329–350, December 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0180-3.

Amoah:2017:EDD

[213] Yaa Tiwaah Amoah, Mohammad Moniruzzaman, Seunghan Lee, Jinho Bae, Seonghun Won, Minji Seong, and Sungchul C. Bai. Evaluation of different dietary additives based on growth performance, innate immunity and disease resistance in juvenile Amur catfish, Silurus asotus. International Aquatic Research, 9(4):351–360, December 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-017-0181-2.

Nebo:2017:AEA

[214] Caroline Nebo, Ken Overturf, Andreas Brezas, Maeli Dal-Pai-Silva, and Maria Célia Portella. Alteration in expression of atrogenes and IGF-1 induced by fasting in Nile tilapia Oreochromis niloticus juveniles. International Aquatic Research, 9(4):361–372, December 2017. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-017-0182-1.

Kurihara:2018:STJ

[215] Takeo Kurihara, Shoji Nakano, Yukihiko Matsuyama, Kazumasa Hashimoto, Katsumasa Yamada, Atsushi Ito, and Masaei Kanematsu. Survival time of juvenile pen shell Atrina pectinata (Bivalvia: Pinnidae) in hyposaline water. International Aquatic Research, 10(1):1–11, March 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-017-0183-0.

Baamstedt:2018:IPM

[216] Ulf Båmstedt and Henrik Larsson. An indoor pelagic mesocosm facility to simulate multiple water-column characteristics. International Aquatic Research, 10(1):13–29, March 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-017-0185-y.

Ghadiri:2018:BAO

[217] Motahereh Ghadiri, Soodabeh Kazemi, Behrooz Heidari, and Mehdi Rassa. Bioactivity of aqueous and organic extracts of sea cucumber Holothuria leucospilota (Brandt 1835) on pathogenic Candida and Streptococci. International Aquatic Research, 10(1):31–43, March 2018. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1007/s40071-017-0186-x.

Mozanzadeh:2018:EDP

[218] Mansour Torfi Mozanzadeh, Morteza Yaghoubi, Jasem G. Marammazi, Omid Safari, and Enric Gisbert. Effects of dietary protein and essential amino acid deficiencies on growth, body composition, and digestive enzyme activities of silvery-black porgy (Sparidentex hasta). International Aquatic Research, 10(1):45–55, March 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-017-0187-9.

Ende:2018:GPH

[219] Stephan S. W. Ende, Vanessa Fuchs, Annabel Schuhn, Christiane von der Marwitz, Andrea Wirtz, Joachim Henjes, and Matthew Slater. Growth performance of hybrid striped bass (Morone chrysops × M. saxatilis) fed with commercial pike perch and trout diets. International Aquatic Research, 10(1):57–63, March 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0188-3.

Li:2018:GER

[220] Yaqin Li, Shannon L. Meseck, Mark S. Dixon, and Gary H. Wikfors. The East River tidal strait, New York City, New York, a high-nutrient, lowchlorophyll coastal system. International Aquatic Research, 10(1):65–77, March 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
018-0189-2.

daSilva:2018:CLA

[221] Carolina Vieira da Silva and Raoul Henry. Chironomidae larvae associated with Eichhornia azurea leaf detritus: decomposition community structure and colonization dynamics. International Aquatic Research, 10(1):79–93, March 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
018-0190-9.

Teoh:2018:RHI

[222] Chui-Fen Teoh, Leong-Seng Lim, and Gunzo Kawamura. Remarkably high ingestion ratio of acidic food in juvenile marble goby, Oxyeleotris marmorata. International Aquatic Research, 10(1):95–100, March 2018. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1007/s40071-017-0184-z.

Irungu:2018:MCE

[223] F. G. Irungu, C. M. Mutungi, A. K. Faraj, H. Affognon, C. Tanga, S. Ekesi, D. Nakimbugwe, and K. K. M. Fiaboe. Minerals content of extruded fish feeds containing cricket (Acheta domesticus) and black soldier fly larvae (Hermetia illucens) fractions. International Aquatic Research, 10(2):101–113, June 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0191-8.

Gora:2018:EDS

[224] Adnan Hussain Gora, Narottam Prasad Sahu, Sujata Sahoo, Saima Rehman, Showkat Ahmad Dar, Irshad Ahmad, and Deepak Agarwal. Effect of dietary Sargassum wightii and its fucoidan-rich extract on growth, immunity, disease resistance and antimicrobial peptide gene expression in Labeo rohita. International Aquatic Research, 10(2):115–131, June 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-018-0193-6.

Goncalves:2018:OEC

[225] José Fernando M. Gonçalves, Mariana Hinzmann, Jorge Machado, Sandra Araújo, José Carlos Oliveira, and Rodrigo O. A. Ozório. Oxidative effect of l-carnitine on energy metabolism in diploid and triploid rainbow trout (Oncorhynchus mykiss): impact on metabolites. International Aquatic Research, 10(2):133–143, June 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0194-5.

Alghabshi:2018:ASI

[226] Aliya Alghabshi, Brian Austin, and Margaret Crumlish. Aeromonas salmonicida isolated from wild and farmed fish and invertebrates in Oman. International Aquatic Research, 10(2):145–152, June 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1007/s40071-018-0195-4.

Namayandeh:2018:DBI

[227] Armin Namayandeh, David V. Beresford, Keith M. Somers, and Peter J. Dillon. Difference in benthic invertebrate communities of headwater streams can be detected using a short elevation gradient. International Aquatic Research, 10(2):153–164, June 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-018-0197-2.

Agostini:2018:IBB

[228] Vanessa Ochi Agostini, Alexandre José Macedo, and Erik Muxagata. Inhibition of biofilm bacteria and adherent fungi from marine plankton cultures using an antimicrobial combination. International Aquatic Research, 10(2):165–177, June 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0198-1.

Creencia:2018:ESG

[229] Lota Alcantara Creencia and Tadahide Noro. Effects of salinity on the growth and mucous cells of the abalone Haliotis diversicolor Reeve, 1846. International Aquatic Research, 10(2):179–189, June 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1007/s40071-018-0199-0. See correction

[239].

Afandi:2018:TMD

[230] Imane Afandi, Sophia Talba, Ali Benhra, Samir Benbrahim, Rachid Chfiri, Maylis Labonne, Hicham Masski, Raymond Laë, Luis Tito De Morais, Mohammed Bekkali, and Fatima Zohra Bouthir. Trace metal distribution in pelagic fish species from the north-west African coast (Morocco). International Aquatic Research, 10(2):191–205, June 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-018-0192-7.

Leppanen:2018:OCS

[231] Jaakko Johannes Leppänen. An overview of Cladoceran studies conducted in mine water impacted lakes. International Aquatic Research, 10(3):207–221, September 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0204-7.

Petrova:2018:PFP

[232] Inna Petrova, Ignat Tolstorebrov, and Trygve Magne Eikevik. Production of fish protein hydrolysates step by step: technological aspects, equipment used, major energy costs and methods of their minimizing. International Aquatic Research, 10(3):223–241, September 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-018-0207-4.

DCosta:2018:IDD

[233] Avelyno H. D’Costa, S. K. Shyama, M. K. Praveen Kumar, and Trisha M. Fernandes. Induction of DNA damage in the peripheral blood of zebrafish (Danio rerio) by an agricultural organophosphate pesticide, monocrotophos. International Aquatic Research, 10(3):243–251, September 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-018-0201-x.

Yenmak:2018:PFY

[234] Suthajaree Yenmak, Weerakit Joerakate, and Supawadee Poompuang. Prediction of fillet yield in hatchery populations of Asian sea bass, Lates calcarifer (Bloch, 1790) using body weight and measurements. International Aquatic Research, 10(3):253–261, September 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1007/s40071-018-0202-9.

Amaninejad:2018:EDE

[235] Parisa Amaninejad, Homayoun Hosseinzadeh Sahafi, Mehdi Soltani, and Seyed Pezhman Hosseini Shekarabi. Endocrine disrupting effects of 4nonylphenol on plasma vitellogenin, reproductive system and histology in koi carp (Cyprinus carpio). International Aquatic Research, 10(3): 263–274, September 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0203-8.

Yang:2018:ABC

[236] Yi Yang, Yuchun Li, and Zhenlin Liang. Assessment of bacterial communities in skin ulceration in cultured sea cucumber Apostichopus japonicus (Selenka). International Aquatic Research, 10(3):275–282, September 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-018-0205-6.

Ikari:2018:MSH

[237] Takahiro Ikari, Yuki Kobayashi, Yoichiro Kitani, Toshio Sekiguchi, Masato Endo, Akira Kambegawa, Kiyoshi Asahina, Atsuhiko Hattori, Yoshiaki Tabuchi, Thumronk Amornsakun, Kanta Mizusawa, Akiyoshi Takahashi, and Nobuo Suzuki. α-melanocyte-stimulating hormone directly increases the plasma calcitonin level and involves calcium metabolism in goldfish. International Aquatic Research, 10(3):283–292, September 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-018-0206-5.

Muller:2018:AAI

[238] T. Müller, I. Ittzés, Zs. Szőke, Á. Hegyi, E. Mészáros, K. K. Lefler, Z. Bokor, B. Urbányi, and B. Kucska. Attempts on artificial induction of sexual maturation of sterlet (Acipenser ruthenus) and identification of late spermatogenesis stage in hermaphroditic fish. International Aquatic Research, 10(3):293–297, September 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-018-0196-3.

Creencia:2018:CES

[239] Lota Alcantara Creencia and Tadahide Noro. Correction to: Effects of salinity on the growth and mucous cells of the abalone Haliotis diversicolor Reeve, 1846. International Aquatic Research, 10(3):299–301, September 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
018-0200-y. See [229].

Ergonul:2018:TGD

[240] Mehmet Borga Ergönül, Jan Breine, and Erika Van den Bergh. A technical guide to develop a statistically valid fish-based index in compliance with the water framework directive: an evaluation for Turkish freshwaters. International Aquatic Research, 10(4):303–319, December 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-018-0209-2.

Lobo:2018:EDS

[241] Gaspar Lobo, Luís F. Pereira, José F. M. Gonçalves, Maria J. Peixoto, and Rodrigo O. A. Ozório. Effect of dietary seaweed supplementation on growth performance, antioxidant and immune responses in European seabass (Dicentrarchus labrax) subjected to rearing temperature and salinity oscillations. International Aquatic Research, 10(4):321–331, December 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-018-0208-3.

Garaffo:2018:AFD

[242] G. V. Garaffo, M. L. Jaubet, E. N. Llanos, M. A. Saracho Bottero, and R. Elías. Assessing functional diversity of macrobenthic assemblages in sewage-affected intertidal shores. International Aquatic Research, 10 (4):333–347, December 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0211-8.

Kurihara:2018:HVV

[243] Takeo Kurihara, Kazumasa Hashimoto, Shoji Nakano, Yukihiko Matsuyama, Atsushi Ito, Masaei Kanematsu, and Kazumaro Okamura. Horizontal and vertical variations in survival rates of juvenile bivalves, Anadara kagoshimensis (Tokunaga, 1906), Ruditapes philippinarum (Adams & Reeve, 1850) and Atrina sp., experimentally caged in the inner area of the Sea of Ariake, Japan. International Aquatic Research, 10(4):349–359, December 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
018-0212-7.

Gonzalez-Wanguemert:2018:NII

[244] Mercedes González-Wangüemert, Christina C. Roggatz, Maria João Rodrigues, Luísa Barreira, Manuela Moreira da Silva, and Luísa Custódio. A new insight into the influence of habitat on the biochemical properties of three commercial sea cucumber species. International Aquatic Research, 10(4):361–373, December 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0210-9.

Heydari:2018:SSD

[245] N. Heydari, S. M. R. Fatemi, A. Mashinchian, R. Musavi Nadushan, and B. Raeisi. Seasonal species diversity and abundance of phytoplankton from the southwestern Caspian Sea. International Aquatic Research, 10 (4):375–390, December 2018. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-018-0213-6.

Ching:2018:NSE

[246] Fui Fui Ching, Norazmi Othman, Azaharie Anuar, Rossita Shapawi, and Shigeharu Senoo. Natural spawning, embryonic and larval development of F2 hybrid grouper, tiger grouper Epinephelus fuscoguttatus × giant grouper E. lanceolatus. International Aquatic Research, 10(4):391–402, December 2018. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
018-0214-5.

SobrinhoVentura:2019:PPR

[247] Arlene Sobrinho Ventura, Tarcila Sousa de Castro Silva, Ricardo Basso Zanon, Luis Antonio Kioshi Aoki Inoue, and Claudia Andrea Lima Cardoso. Physiological and pharmacokinetic responses in neotropical Piaractus mesopotamicus to the essential oil from Lippia sidoides (Verbenaceae) as an anesthetic. International Aquatic Research, 11(1):1–12, March 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-019-0215-z.

Nuntapong:2019:DEM

[248] Nutt Nuntapong, Wutiporn Phromkunthong, Sunee Wanlem, and Mali Boonyaratpalin. Dietary exposure to melamine and cyanuric acid induced growth reduction, oxidative stress and pathological changes of hepatopancreas in Pacific white shrimp. International Aquatic Research, 11(1):13–31, March 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
019-0216-y.

Suchocki:2019:RPI

[249] Christopher R. Suchocki and Osvaldo J. Sepulveda-Villet. The role of phototaxis in the initial swim bladder inflation of larval yellow perch (Perca flavescens). International Aquatic Research, 11(1):33–42, March 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-019-0217-x.

Gallardo-Colli:2019:RWB

[250] Alfredo Gallardo-Collí, Carlos Iván Pérez-Rostro, and Martha Patricia Hernández-Vergara. Reuse of water from biofloc technology for intensive culture of Nile tilapia (Oreochromis niloticus): effects on productive performance, organosomatic indices and body composition. International Aquatic Research, 11(1):43–55, March 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-019-0218-9.

Gabriel:2019:EDA

[251] Ndakalimwe Naftal Gabriel, Margit R. Wilhelm, Habte-Michael Habte-Tsion, Percy Chimwamurombe, Edosa Omoregie, Linda N. Iipinge, and Kaspar Shimooshili. Effect of dietary Aloe vera polysaccharides supplementation on growth performance, feed utilization, hemato-biochemical parameters, and survival at low pH in African catfish (Clarias gariepinus) fingerlings. International Aquatic Research, 11(1):57–72, March 2019. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https:

//link.springer.com/article/10.1007/s40071-019-0219-8.

Nunes:2019:EAB

[252] Etelvina M. C. G. Nunes, Maria M. G. Pereira, Amilton P. R. Costa, María N. B. A. Araripe, Rodrigo M. Calvet, Carina M. Pereyra, Maria L. X. Azevedo, Raizza E. E. Pinheiro, Verbena C. Alves, Airton M. Conde Júnior, Lidiane S. N. Ramos, Joao B. Lopes, Camila G. Marinho, and María C. S. Muratori. Effects of aflatoxin B_{1} on performance and health of tambaqui fingerlings (Colossoma macropomum). International Aquatic Research, 11(1):73–83, March 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-0220-2.

Hamayeli:2019:AAA

[253] Homa Hamayeli, Mehdi Hassanshahian, and Majid Askari Hesni. The antibacterial and antibiofilm activity of sea anemone (Stichodactyla haddoni) against antibiotic-resistant bacteria and characterization of bioactive metabolites. International Aquatic Research, 11(1):85–97, March 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-019-0221-1.

eSousa:2019:PGG

[254] João M. C. e Sousa, Vitor A. de Oliveira, Ana P. Peron, Ataíde M. V. Lima, Ila B. S. Sales, Felipe C. C. da Silva, Leonardo H. G. M. Lima, Leomá A. Matos, Marcus V. O. B. de Alencar, Luzia C. Rodrigues, Manik C. Shill, Muhammad T. Islam, Ana Amélia C. Melo-Cavalcante, Cláudia C. Bonecker, and Horácio F. J. Junior. Pollutants of the Guaribas river water and their toxicogenic effects. International Aquatic Research, 11(1):99–112, March 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-0222-0.

Daskalova:2019:FFW

[255] Aleksandra Daskalova. Farmed fish welfare: stress, post-mortem muscle metabolism, and stress-related meat quality changes. International Aquatic Research, 11(2):113–124, June 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-019-0230-0.

Sipauba-Tavares:2019:MAG

[256] Lúcia H. Sipaúba-Tavares, João B. K. Fernandes, Gustavo L. Melo-Santos, and Bruno Scardoeli-Truzzi. Microalgae Ankistrodesmus gracilis as feed ingredient for ornamental fish Xiphophorus maculatus. International Aquatic Research, 11(2):125–134, June 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-019-0223-z.

Bitalebi:2019:EAP

[257] Sedighe Bitalebi, Mehdi Nikoo, Kaveh Rahmanifarah, Farzaneh Noori, and Hassan Ahmadi Gavlighi. Effect of apple peel extract as natural antioxidant on lipid and protein oxidation of rainbow trout (Oncorhynchus mykiss) mince. International Aquatic Research, 11(2):135–146, June 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-019-0224-y.

Rudneva:2019:UFE

[258] Irina Ivanovna Rudneva. Use of fish embryo biomarkers for the evaluation of mazut toxicity in marine environment. International Aquatic Research, 11(2):147–157, June 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-0225-x.

Mero:2019:IWM

[259] Fedelia Flor C. Mero, Fiona. L. Pedroso, Mary Jane S. Apines-Amar, Josel F. Cadangin, Donna C. Rendaje, Christine S. Verde, Jean Rose H. Maquirang, and Lily Anne G. Piňosa. Influence of water management, photoperiod and aeration on growth, survival, and early spat settlement of the hatchery-reared green mussel, Perna viridis. International Aquatic Research, 11(2):159–172, June 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-0226-9.

Sibero:2019:SAF

[260] Mada Triandala Sibero, Yasuhiro Igarashi, Ocky Karna Radjasa, Agus Sabdono, Agus Trianto, Dewi Seswita Zilda, and Yusuf Jati Wijaya. Sponge-associated fungi from a mangrove habitat in Indonesia: species composition, antimicrobial activity, enzyme screening and bioactive profiling. International Aquatic Research, 11(2):173–186, June 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-019-0227-8.

Velasco-Blanco:2019:TPT

[261] Gabriela Velasco-Blanco, Ana Denise Re, Fernando Díaz, Leonardo Ibarra-Castro, Maria Isabel Abdo de la Parra, Luz Estela Rodríguez-Ibarra, and Carlos Rosas. Thermal preference, tolerance, and thermal aerobic scope in clownfish Amphiprion ocellaris (Cuvier, 1830) predict its aquaculture potential across tropical regions. International Aquatic Research, 11(2):187–197, June 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-0228-7.

Nathanailides:2019:MDD

[262] Cosmas Nathanailides, Dimitris Klaoudatos, Costas Perdikaris, Spyros Klaoudatos, Markos Kolygas, and Fotini Athanassopoulou. Metabolic differentiation of diploid and triploid European sea bass juveniles. International Aquatic Research, 11(2):199–206, June 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-019-0229-6.

Sodre:2019:HPM

[263] Elder de Oliveira Sodré and Reinaldo Luiz Bozelli. How planktonic microcrustaceans respond to environment and affect ecosystem: a functional trait perspective. International Aquatic Research, 11(3):207–223, September 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
019-0233-x.

Kumari:2019:EAA

[264] Sunil Kumari, Anil Kumar Tyor, and Anita Bhatnagar. Evaluation of the antibacterial activity of skin mucus of three carp species. International Aquatic Research, 11(3):225–239, September 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-019-0231-z.

Diodato:2019:ATS

[265] Soledad L. Diodato, Oscar A. Amin, and Laura I. Comoglio. Ammonia toxicity in southern king crab (Lithodes santolla, Molina 1742) larvae. International Aquatic Research, 11(3):241–251, September 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-019-0232-y.

Tariang:2019:EAL

[266] Kermi u Tariang, Sunkam N. Ramanujam, and Bidyadhar Das. Effect of arsenic (As) and lead (Pb) on glycogen content and on the activities of selected enzymes involved in carbohydrate metabolism in freshwater catfish, Heteropneustes fossilis. International Aquatic Research, 11(3):253–266, September 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
019-00234-2.

Takagi:2019:DSM

[267] Takashi Takagi, Toshio Sekiguchi, Yuichi Sasayama, Atsuhiko Hattori, Takushi X. Watanabe, Yoichiro Kitani, Yoshiaki Tabuchi, Hajime Matsubara, Ajai K. Srivastav, Tran Ngoc Hai, Thumronk Amornsakun, and Nobuo Suzuki. Development of a system for measuring calcitonin in the stingray Dasyatis akajei (a cartilaginous fish): the possible involvement of stingray calcitonin in gonadal development. International Aquatic Research, 11(3):267–276, September 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-019-00236-0.

Hassanin:2019:NCG

[268] Abeer A. I. Hassanin and Yoshio Kaminishi. A novel cytochrome P450 1D1 gene in Nile tilapia fish (Oreochromis niloticus): partial cDNA cloning and expression following benzo-a-pyrene exposure. International Aquatic Research, 11(3):277–285, September 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-019-00237-z.

Avillanosa:2019:USB

[269] Arlene L. Avillanosa and Christopher Marlowe A. Caipang. Use of sodium bicarbonate as an inexpensive general anesthetic for juvenile red tilapia hybrids. International Aquatic Research, 11(3):287–294, September 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://link.springer.com/article/10.1007/s40071-019-00235-1.

Rutkovska:2019:STM

[270] Ilze Rutkovska, Ruta Medne, and Santa Purvin.a. Seawater tolerance and morphological assessment of yearling sea trout (Salmo trutta L.). International Aquatic Research, 11(3):295–302, September 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

link.springer.com/article/10.1007/s40071-019-00238-y.

Thang:2019:MCF

[271] Nguyen Dinh Thang, Le Dinh Tu, Nguyen Thi Le Na, Ngo Thi Trang, and Phan Tuan Nghia. Melanin-containing feedstuffs protect Litopenaeus vannamei from white spot syndrome virus. International Aquatic Research, 11(3):303–310, September 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-00240-4.

Ariyati:2019:ETT

[272] Restiana Wisnu Ariyati, Sri Rejeki, Lestari L. Widowati, Tita Elfitasari, and Roel H. Bosma. Effect of three types of liquid compost combined with Avicennia marina leaves on growth and survival of tiger prawns (Penaeus monodon). International Aquatic Research, 11(4):311–321, December 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-

019-00239-x. See correction [273].

Ariyati:2019:CET

[273] Restiana Wisnu Ariyati, Sri Rejeki, Lestari L. Widowati, Tita Elfitasari, and Roel H. Bosma. Correction to: Effect of three types of liquid compost combined with Avicennia marina leaves on growth and survival of tiger prawns (Penaeus monodon). International Aquatic Research, 11 (4):323–324, December 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-00248-w. See [272].

Doughty:2019:EDF

[274] Katarina H. Doughty, Shawn R. Garner, Mark A. Bernards, John W. Heath, and Bryan D. Neff. Effects of dietary fishmeal substitution with corn gluten meal and poultry meal on growth rate and flesh characteristics of Chinook salmon (Oncorhynchus tshawytscha). International Aquatic Research, 11(4):325–334, December 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.com/

article/10.1007/s40071-019-00241-3.

Hisano:2019:EGN

[275] Hamilton Hisano, Phillipe T. L. Barbosa, Liliam A. Hayd, and Cristiano C. Mattioli. Evaluation of Nile tilapia in monoculture and polyculture with giant freshwater prawn in biofloc technology system and in recirculation aquaculture system. International Aquatic Research, 11 (4):335–346, December 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-00242-2.

Natsumeda:2019:HSE

[276] Takaharu Natsumeda and Kei’ichiro Iguchi. Habitat-specific effects of interstitial space between stream substrate particles on the colonization of aquatic organisms. International Aquatic Research, 11(4):347–358, December 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://link.springer.com/article/10.1007/s40071-
019-00243-1.

Nagasoe:2019:PUG

[277] Sou Nagasoe, Tetsuji Hisada, Goh Nishitani, Hiroyuki Takeuchi, and Kazutaka Sakiyama. Phytoplankton uptake and growth rate in the Japanese egg cockle Fulvia mutica. International Aquatic Research, 11 (4):359–375, December 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-00244-0.

Sharma:2019:EMB

[278] Shaishav Sharma, Ekta Shah, Deepti Davla, Gaurav Dixit, Aesha Patel, and Adepu Kiran Kumar. Effect of microalga-based diet on oxidative stress enzymes of African catfish, Clarias gariepinus. International Aquatic Research, 11(4):377–387, December 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.springer.

com/article/10.1007/s40071-019-00245-z.

Cocci:2019:IPC

[279] Paolo Cocci, Francesco Alessandro Palermo, Stefania Pucciarelli, Antonino Miano, Massimiliano Cuccioloni, Mauro Angeletti, Alessandra Roncarati, and Gilberto Mosconi. Identification, partial characterization, and use of grey mullet (Mugil cephalus) vitellogenins for the development of ELISA and biosensor immunoassays. International Aquatic Research, 11(4):389–399, December 2019. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://link.springer.com/article/10.
1007/s40071-019-00246-y.

Chappell:2019:ECF

[280] Jessica Chappell, Stefanie Whitmire, David Sotomayor-Ramírez, and Gustavo Martínez-Rodríguez. Effects of Corbicula fluminea on the nutrient concentration and phytoplankton biomass of tropical reservoirs. International Aquatic Research, 11(4):401–415, December 2019. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://link.

springer.com/article/10.1007/s40071-019-00247-x.

Lee:2020:AAF

[281] Yang Lee, Lesley Maurice Bilung, Badiozaman Sulaiman, and Yee Ling Chong. The antibacterial activity of fish skin mucus with various extraction solvents and their in-vitro evaluation methods. International Aquatic Research, 12(1):1–21, March 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_670998.html;   https:/
/intelaquares.tonekabon.iau.ir/article_670998_cf6c6b206e46dc16d8922f539ad70094.
pdf.

Gharaei:2020:ETT

[282] Ahmad Gharaei, Hossein Ebrahimi Jorjani, Javad Mirdar Harijani, and Hamed Kolangi Miandare. Effects of Tribullus terrestris extract on masculinization, growth indices, sex determination reversal and steroid hormones level in zebra fish (Danio rerio). International Aquatic Research, 12(1):22–29, March 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_670992.html;   https:/
/intelaquares.tonekabon.iau.ir/article_670992_d8a03069c7d7a064e9d5a52a9267e370.
pdf.

Cejudo:2020:EAU

[283] Eduardo Cejudo, William Taylor, and Sherry Schiff. Epilithic algae from an urban river preferentially use ammonium over nitrate. International Aquatic Research, 12(1):30–39, March 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_671068.html;   https:/
/intelaquares.tonekabon.iau.ir/article_671068_cc359abf08662c0515cd89fb84274200.
pdf.

Saleh:2020:BFW

[284] Norhan Saleh, Elham Wassef, Kholoud Barakat, and Heba Helmy. Bioconversion of fish-wastes biomass into a fish meal alternate for European seabass (Dicentrarchus labrax) diets. International Aquatic Research, 12 (1):40–52, March 2020. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_671145.html;   https://intelaquares.tonekabon.iau.ir/
article_671145_03be305c3eaaca0a378a0beca102511a.pdf.

Laurenz:2020:ENE

[285] Jan Laurenz, Arne Georg, Heinz Brendelberger, and Kai Lehmann. Effects of nitrate on early life stages of Astacus astacus (Linnaeus, 1758) and Procambarus virginalis (Lyko, 2017). International Aquatic Research, 12(1):53–62, March 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_671232.html;   https:/
/intelaquares.tonekabon.iau.ir/article_671232_50aab81f9ee2dd847267b3c21ed6b115.
pdf.

Kantserova:2020:DSD

[286] Nadezhda Kantserova, Liudmila A. Lysenko, Maria V. Churova, Ekaterina D. Tushina, Irina Sukhovskaya, and Nina Nemova. Dietary supplement with dihydroquercetin and arabinogalactan affects growth performance, intracellular protease activities and muscle-specific gene expression in bacterially infected Oncorhynchus mykiss. International Aquatic Research, 12(1):63–73, March 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_671431.html;   https:/
/intelaquares.tonekabon.iau.ir/article_671431_c936e950bdcfce07eb1548e60405d6c8.
pdf.

Jusoh:2020:IFM

[287] Malinna Jusoh, Nor Azman Kasan, Farah Shahirah Hashim, Noorazilah Haris, Muhammad Faiz Zakaria, Nor Najwa Mohamed, Nadiah W. Rasdi, Mohd Effendy Abd Wahid, Tomoyo Katayama, and Kazutaka Takahashi. Isolation of freshwater and marine indigenous microalgae species from Terengganu water bodies for potential uses as live feeds in aquaculture industry. International Aquatic Research, 12(1):74–83, March 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_671730.html;   https:/
/intelaquares.tonekabon.iau.ir/article_671730_7778816a3f54f6072b0404e771cf02c6.
pdf.

Setubal:2020:ZSP

[288] Rayanne Setubal, Rodrigo Araújo do Nascimento, and Reinaldo Bozelli. Zooplankton secondary production: main methods, overview and perspectives from Brazilian studies. International Aquatic Research, 12(2):85–99, June 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
672861.html;   https://intelaquares.tonekabon.iau.ir/article_
672861_baf0e201888a0b8b5ce68d7dc5b8081a.pdf. See correction [302].

Mousavi:2020:RIM

[289] Seyedehlaleh Mousavi, Somayeh Zahedinezhad, and Jiun Yan Loh. A review on insect meals in aquaculture: the immunomodulatory and physiological effects. International Aquatic Research, 12(2):100–115, June 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_673050.html;
https://intelaquares.tonekabon.iau.ir/article_673050_29aa6de2b304d7e1baef56d4f672e902.
pdf.

Dara:2020:CAS

[290] Pavan Kumar Dara, Krishnamoorthy Elavarasan, and Bangalore Aswathnarayan Shamasundar. Characterization of antioxidant and surface active properties of gelatin hydrolysates obtained from croaker fish skin. International Aquatic Research, 12(2):116–126, June 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_674048.html;   https:/
/intelaquares.tonekabon.iau.ir/article_674048_9933a663143581533083afcac3b936ac.
pdf.

Fujimoto:2020:EOS

[291] Rodrigo Fujimoto, Gabriela R. Gomes, Roberson Sakabe, Claudinei de Cruz, Natalino Costa Sousa, Márcia Valéria Silva Couto, Peterson Emmanuel Guimarães Paixão, Juliana Meneses, Paulo Carneiro, and Alexandre Maria. Effects of organic selenium on thrombocytes and parasite infestation in Nile tilapia juveniles. International Aquatic Research, 12(2):127–132, June 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_672367.html;   https:/
/intelaquares.tonekabon.iau.ir/article_672367_3e0ce5138306734079ac04072334b08c.
pdf.

Eagderi:2020:LWR

[292] Soheil Eagderi, Atta Mouludi-Saleh, and Erdo an Cicek. Length–weight relationship of ten species of Leuciscinae sub-family (Cyprinidae) from Iranian inland waters. International Aquatic Research, 12(2):133–136, June 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_672429.html;
https://intelaquares.tonekabon.iau.ir/article_672429_f801811b70da31ad941cbafc8fe3c0ed.
pdf.

Haoujar:2020:EFD

[293] Imane Haoujar, Jamal Abrini, Housni Chadli, Adil Essafi, Hassan Nhhala, Kamal Chebbaki, Francesco Cacciola, and Nadia Skali senhaji. Effect of four diets based on three microalgae on the growth performance and quality of Mediterranean mussel flesh, Mytilus galloprovincialis. International Aquatic Research, 12(2):137–145, June 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_672732.html;   https:/
/intelaquares.tonekabon.iau.ir/article_672732_2ebed1c1c74b751d6c6ce1442ad542a2.
pdf.

Katayama:2020:CRT

[294] Satoshi Katayama, Masayuki Yamamoto, and Thomas Munroe. Comparison of the reproductive traits of male and female Soleoidei fishes in Japan. International Aquatic Research, 12(2):146–150, June 2020. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_672962.html;   https:/

/intelaquares.tonekabon.iau.ir/article_672962_b38352eb836668c3901bf95b47f72461.
pdf.

Gammone:2020:PAM

[295] Maria Alessandra Gammone and Nicolantonio D’Orazio. Potential applications of marine macrolides: New drugs from the sea? International Aquatic Research, 12(3):??, September 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_674222.html;   https:/
/intelaquares.tonekabon.iau.ir/article_674222_82c1b835018e8337acbef7659fd543aa.
pdf.

Lajus:2020:SHT

[296] Dmitry Lajus, Liudmila Lysenko, Nadezhda Kantserova, Ekaterina Tushina, Tatiana Ivanova, and Nina Nemova. Spatial heterogeneity and temporal dynamics of protein-degrading activity and lifehistory traits in threespine stickleback Gasterosteus aculeatus. International Aquatic Research, 12(3):??, September 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_675340.html;   https:/
/intelaquares.tonekabon.iau.ir/article_675340_9585d919216ec04404e7e3fd08964d39.
pdf.

ElQendouci:2020:SMF

[297] Mouna El Qendouci, Khadija Amenzoui, and Ahmed Yahyaoui. Size at maturity, fecundity and spawning period of anchovy Engraulis encrasicolus (Linnaeus, 1758) in the central area of the Moroccan Atlantic coast. International Aquatic Research, 12(3):??, September 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_675341.html;   https:/
/intelaquares.tonekabon.iau.ir/article_675341_6461384d2180371789cf5b0b3da21db4.
pdf.

Loh:2020:BEA

[298] Jiun Yan Loh, Alicia Hui-Ying Khor, Kok Song Lai, and Hon Jung Liew. Biosorption efficacy of alginate-immobilized live and metal chlorideactivated Azolla microphylla in Pb(II) removal from aqueous solution. International Aquatic Research, 12(3):??, September 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_674703.html;   https:/
/intelaquares.tonekabon.iau.ir/article_674703_2bd08bb075b36362da5da447bb65dd2a.
pdf.

Rojas-Garcia:2020:EPI

[299] Carlos R. Rojas-Garcia, Isabel Jimenez-Garcia, and Edgar Mendoza-Franco. Ecto- parasitic infection in Nile tilapia (Oreochromis niloticus) fry during male reversal in Veracruz, México. International Aquatic Research, 12(3):??, September 2020. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_675342.html;   https://intelaquares.tonekabon.iau.ir/
article_675342_8e4fbfb4593ef8e66500534a734111f3.pdf.

Nur:2020:EWT

[300] Firman Nur, Agung Batubara, Kartini Eriani, Usman Tang, Abdullah A. Muhammada, Mohd N. Siti-Azizah, Martin Wilkes, Nur Fadli, Syamsul Rizal, and Zainal A. Muchlisin. Effect of water temperature on the physiological responses in Betta rubra, Perugia 1893 (Pisces: Osphronemidae). International Aquatic Research, 12(3):??, September 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_674704.html;   https:/
/intelaquares.tonekabon.iau.ir/article_674704_ae3e52b18a696970aa949a9122d017a2.
pdf.

Goncalves:2020:PBA

[301] Bruno Goncalves and Percilia Cardoso Giaquinto. Propofol and benzocaine anesthetics responses profiles in Nile tilapia. International Aquatic Research, 12(3):??, September 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_675419.html;   https:/
/intelaquares.tonekabon.iau.ir/article_675419_22be0d461d447fa85d6711f0ddbcc33f.
pdf.

Setubal:2020:CZS

[302] Rayanne Barros Setubal, Rodrigo Araújo do Nascimento, and Reinaldo Luiz Bozelli. Correction to: Zooplankton secondary production: main methods, overview and perspectives from Brazilian studies. International Aquatic Research, 12(3):??, September 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_675913.html;   https:/
/intelaquares.tonekabon.iau.ir/article_675913_bd180d5e57bb23318d2e5c350ffc67f8.

pdf. See [288].

Oliveira:2020:MLV

[303] Manildo Oliveira, Ana Ventura, Aline Fraga, Maria Helena Baeta Neves, Rafael Dantas, Renato Lopes, Rachel Hauser-Davis, João B.

Salles, Frederico Bastos, Vera Cunha Bastos, and Jayme Bastos. Mugil liza (Valenciennes, 1836) biomarker responses in a hypersaline lagoon in southeastern Brazil before and after a microalgal bloom. International Aquatic Research, 12(4):231–241, December 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_676569.html;   https:/
/intelaquares.tonekabon.iau.ir/article_676569_90da068caa876c5056074cf9cd2fbbf1.
pdf.

Cardoso:2020:PBR

[304] Israel Cardoso, Michelly Soares, Carolina Angelis, Giovanni Ferri, Marcia Ishikawa, Claudio Jonsson, Francisco Rantin, and Fernanda Sampaio. Physiological and biochemical responses of Nile tilapia (Oreochromis niloticus) to acute trichlorfon exposure. International Aquatic Research, 12(4):243–253, December 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_676944.html;   https:/
/intelaquares.tonekabon.iau.ir/article_676944_118dd3a5a4e6a9bc1cb702c7178bfec2.
pdf.

Laurenz:2020:EDE

[305] Jan Laurenz, Heinz Brendelberger, and Kai Lehmann. Effects of Diclofenac on the embryonic development of freshwater crayfish. International Aquatic Research, 12(4):255–265, December 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677597.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677597_a1c2e66fe3764cd2522c28459f4caf25.
pdf.

Meena:2020:PES

[306] Dharmendra Kumar Meena, Amiya Sahoo, Prem Prakash Srivastava, Narottam Prasad Sahu, Bijay Kumar Behera, Simanku Borah, Himanshu Sekhar Swain, and Basanta Kumar Das. Protective effects of selected solvent extracts of Terminalia arjuna against environment mediated parasitic infection in Labeo rohita. International Aquatic Research, 12(4):267–278, December 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677765.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677765_e5720d454335c0f11aca63360b4718b7.
pdf.

Nagai:2020:RSS

[307] Takeshi Nagai, Masataka Saito, Yasuhiro Tanoue, Norihisa Kai, and Nobutaka Suzuki. Reduced-salt shrimp sauces from Alaskan pink shrimp and nonglutinous rice cultivar Tsuyahime koji: Preparation and characterization. International Aquatic Research, 12(4):279–290, December 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_677944.html;
https://intelaquares.tonekabon.iau.ir/article_677944_94b652fa2909797bfea330c961cd3509.
pdf.

Mameloco:2020:SCM

[308] Emelyn Joy Mameloco and Rex Ferdinand Traifalgar. Supplementation of combined mannan oligosaccharide and β-glucan immunostimulants improves immunological responses and enhances resistance of Pacific whiteleg shrimp, Penaeus vannamei, against Vibrio parahaemolyticus infection. International Aquatic Research, 12(4):291–299, December 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677644.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677644_75efbe4bb337d280875b9ff8f40e0c88.
pdf.

Adikesavalu:2020:IEA

[309] Harresh Adikesavalu, Farhana Hoque, and Thangapalam Abraham. Immune effector activities and clinical biochemistry of normal pangas catfish Pangasius pangasius (Hamilton, 1822). International Aquatic Research, 12(4):301–307, December 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677851.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677851_d670ba6e8421619395829e5a98635e1b.
pdf.

Ranjbar:2020:EWS

[310] Milad Ranjbar and Majid Mohammad Nejad. Effect of water salinity on enzymatic and hormonal indices of (Oncorhynchus mykiss) fingerlings. International Aquatic Research, 12(4):309–314, December 2020. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_677945.html;   https:/
/intelaquares.tonekabon.iau.ir/article_677945_8821371f8b71eb4d15f65665d6317336.
pdf.

Cesena:2021:UUY

[311] Carlos Ernesto Ceseña, Fernando Vega-Villasante, Gabriel Aguirre-Guzman, Antonio Luna-Gonzalez, and Angel Campa-Cordova. Update on the use of yeast in shrimp aquaculture: a minireview. International Aquatic Research, 13(1):1–16, March 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_678897.html;   https:/
/intelaquares.tonekabon.iau.ir/article_678897_b5c386a05e2c23c8cf5f0cd51472ab1a.
pdf.

Blasco:2021:ATE

[312] Felipe Blasco, Diana Monteiro, Cleoni Carvalho, and Francisco Rantin. Acute toxicity and effects of the Roundup Transorb(R), a glyphosatebased herbicide, on freshwater teleost matrinxã, Brycon amazonicus. International Aquatic Research, 13(1):17–28, March 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_678848.html;   https:/
/intelaquares.tonekabon.iau.ir/article_678848_679bfc81afe0bdd4012174bae109eeb5.
pdf.

Pourali:2021:SSH

[313] Aflaton Pourali and Mehdi Sohrabnezhad. Study of steroid hormones, biochemical parameters and body size based on different gonad’s quality and gonad’s fat content from sonogrphy method in female cultured Siberian sturgeon (Acipenser baerii). International Aquatic Research, 13(1):29–36, March 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_679355.html;   https:/
/intelaquares.tonekabon.iau.ir/article_679355_577b7059969664096b939f38865c5946.
pdf.

Adeniyi:2021:EDT

[314] Olarinke Victoria Adeniyi, Flora Eyibio Olaifa, Benjamin Obukowho Emikpe, and Samuel Temitope Ogunbanwo. Effects of dietary tamarind (Tamarindus indica L.) leaves extract on growth performance, nutrient utilization, gut physiology, and susceptibility to Aeromonas hydrophila infection in Nile tilapia (Oreochromis niloticus L.). International Aquatic Research, 13(1):37–51, March 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_680525.html;   https:/
/intelaquares.tonekabon.iau.ir/article_680525_148bca9c1057e530e230a032fb24a274.
pdf.

Ardohain:2021:MDS

[315] Diego Martin Ardohain, Nestor Adrian Gabellone, and Maria Cristina Claps. Main drivers in the structure and dynamics of the zooplankton community in a Pampean seepage shallow lake throughout an annual cycle during turbid and clear water regimes. International Aquatic Research, 13(1):53–70, March 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_680688.html;   https:/
/intelaquares.tonekabon.iau.ir/article_680688_e51c17a375ea1b5c29cb2ba881d76073.
pdf.

Mejia-Mejia:2021:EFM

[316] Moises Mejia-Mejia, Elsah Arce, Judith García-Rodríguez, and Luis M. Burciaga. Effect of feeding mosquito larvae on the coloration of fighting fish (Betta splendens) during grow-out. International Aquatic Research, 13(1):71–79, March 2021. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_679754.html;   https://intelaquares.tonekabon.iau.ir/
article_679754_7dae80ff41ab8f4375737b054682a88b.pdf.

Srivastav:2021:ECC

[317] Ajai Kumar Srivastav, Khushbu Agarwal, Abhishek Kumar, ManiRam Prasad, Sunil Kumar Srivastav, and Nobuo Suzuki. Effect of chromium chloride on serum calcium and phosphate levels of stinging catfish Heteropneustes fossilis. International Aquatic Research, 13(1):81–87, March 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_679966.html;
https://intelaquares.tonekabon.iau.ir/article_679966_b8db7e8a79a02452104b92b5048d39c8.
pdf.

Naiel:2021:PPS

[318] Mohammed Naiel, Mahmoud Elnakeeb, Lydia Vasilyeva, Natalia Sudakova, Adelya Anokhina, Ahmed G. A. Gewida, and Mahmoud S. Amer. Paddlefish, Polyodon spathula: Historical, current status and future aquaculture prospects in Russia. International Aquatic Research, 13(2):89–107, June 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_680892.html;   https:/
/intelaquares.tonekabon.iau.ir/article_680892_323041c432ca52e03691cb22b92d6808.
pdf.

Musa:2021:PVR

[319] Nadirah Musa, Chen Fei Low, Hazza Roshada Ramli, Mohammad Tajuddin Abdul Manaf, Kok Leong Lee, Shau Hwai Tan, Alia Syafiqah Aznan, and Najiah Musa. Prominent vulnerability of red hybrid tilapia (Oreochromis spp.) liver to heat stress-induced oxidative damage. International Aquatic Research, 13(2):109–118, June 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_681276.html;   https:/
/intelaquares.tonekabon.iau.ir/article_681276_808358f8ac35db3af3ba2b8fab7c3629.
pdf.

Hong:2021:ESI

[320] Peiqi Hong, Shengjian Jiang, and Satoshi Katayama. Experimental study on the influence of water temperature on the otolith formation of the marbled flounder (Pseudopleuronectes yokohamae). International Aquatic Research, 13(2):119–128, June 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_681575.html;   https:/
/intelaquares.tonekabon.iau.ir/article_681575_9bbdd546795277e5e5cdcf85bc479ef4.
pdf.

Tariang:2021:EAL

[321] Kermi u Tariang, Kitboklang Khongsti, and Bidyadhar Das. Effect of arsenic and lead on glucose level and expression of ERK in air-breathing catfish, Heteropneustes fossilis. International Aquatic Research, 13(2):129– 137, June 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
681776.html;   https://intelaquares.tonekabon.iau.ir/article_
681776_11db97b503bb6150e0c291bc195e627c.pdf.

Hassan:2021:MAS

[322] Md Mahbubul Hassan, Imran Hoshan, Mst Antara Masuda, Imran Parvez, Nadiah W. Rasdi, and Chanchal Roy. Motility activation and short-term storage of silver carp, Hypopthalmichthys molitrix, sperm. International Aquatic Research, 13(2):139–145, June 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_682429.html;   https:/
/intelaquares.tonekabon.iau.ir/article_682429_983eb4b90d39b8ad22718af07d607748.
pdf.

Bozkurt:2021:ESD

[323] Yusuf Bozkurt and lker Yavas. Effect of supplementations of docosahexaenoic acid (DHA) into a tris-glucose based extender on the post-thaw sperm quality, fertility and hatching rates in brown trout (Salmo trutta macrostigma) following cryopreservation. International Aquatic Research, 13(2):147–154, June 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_682473.html;   https:/
/intelaquares.tonekabon.iau.ir/article_682473_698a61dc17998d6c4e2ef902febd191f.
pdf.

Haghi:2021:IIS

[324] Maryam Haghi, Mahmoud Mohseni, and Masoud Ghane. Intra- and interspecific variability of IGF-I expression in farmed cold and warm-water species in Iran. International Aquatic Research, 13(2):155–162, June 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_682480.html;
https://intelaquares.tonekabon.iau.ir/article_682480_6a17126052612abd7d40d9accc8e107d.
pdf.

Kobayashi:2021:ESA

[325] Yasuhisa Kobayashi, Takamasa Morishita, and Hisashi Chuda. Effect of synthetic androgen treatment on the gonad of juvenile longtooth grouper (Epinephelus bruneus). International Aquatic Research, 13(3):163–171, September 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
683077.html;   https://intelaquares.tonekabon.iau.ir/article_
683077_a34b0919bfaf03907b73c3c815f2bb70.pdf.

Hoque:2021:QAM

[326] Md. Sazedul Hoque, Md. Abu Sayed Ripon, Md. Belal Hossain, Md. Arifur Rahman, Md. Mahmudul Hasan, Biplob Dey Mithun, Avirup Saha, Suprakash Chakma, Mst. Niloy Jaman, Akhinur Rahman, and Muhammad Manjurul Karim. Quality assessment and molecular identification of bacteria across the black tiger shrimp (Penaeus monodon) supply chains in Southern Bangladesh. International Aquatic Research, 13(3):173–184, September 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_683284.html;   https:/
/intelaquares.tonekabon.iau.ir/article_683284_f15732042087fdb07e49d8c59e5ef6a1.
pdf.

Tarallo:2021:TER

[327] Andrea Tarallo, Christyn Bailey, Claudio Agnisola, and Giuseppe D’Onofrio. A theoretical evaluation of the respiration rate partition in the Gasterosteus aculeatus-Schistocephalus solidus host-parasite system. International Aquatic Research, 13(3):185–194, September 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_683464.html;   https:/

/intelaquares.tonekabon.iau.ir/article_683464_4b791304c3aeb83809deb83a84bc3356.
pdf.

Kamaruding:2021:EHS

[328] Noor Azlina Azlina Kamaruding and Mhd Ikhwanuddin Abdullah. Effects of handling stress and moulting cycle on reproductive performance and digestive capacity of female giant freshwater prawn, Macrobrachium rosenbergii. International Aquatic Research, 13(3):195–207, September 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_683711.html;
https://intelaquares.tonekabon.iau.ir/article_683711_97943eed11b96d8485495d7a51f2e01f.
pdf.

Santos:2021:CSF

[329] Ítala Gabriela S. Santos, Glauber Pereira C. Santos, Carlos Yure B. Oliveira, Clarissa Vilela F. S. Campos, Luis Otávio Brito, and Alfredo Olivera Gálvez. Can shrimp farming wastewater negatively affect water quality and zooplankton community structure of a neotropical estuary? A case study during a productive cycle of Litopenaeus vannamei. International Aquatic Research, 13(3):209–217, September 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_685355.html;   https:/
/intelaquares.tonekabon.iau.ir/article_685355_1bf67fcf3088a9baf2dbdc22fa55e492.
pdf.

Kurosaki:2021:PGM

[330] Mikihei Kurosaki, Mayu Kunimoto, Shiro Itoi, Nobuhiko Akiyama, and Haruo Sugita. Predominant gut microbiota in the early life stages of red seabream (Pagrus major) raised in indoor tanks. International Aquatic Research, 13(3):219–226, September 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_683322.html;   https:/
/intelaquares.tonekabon.iau.ir/article_683322_8202b9c410fcff6b4ccd564f7173bedc.
pdf.

Kembenya:2021:GPM

[331] Elijah Migiro Kembenya and Robert Nyakwama Ondiba. Growth performance of male monosex and mixed sex Nile tilapia (Oreochromis niloticus L.) reared in cages, Lake Victoria, Kenya. International Aquatic Research, 13(3):227–232, September 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_684658.html;   https:/

/intelaquares.tonekabon.iau.ir/article_684658_9032280466dad93a2926f45ec7256174.
pdf.

Tran-Van:2021:PPP

[332] Hieu Tran-Van, Phuong Thao Thi Nguyen, Quoc-Gia Mai, and Thuoc Linh Tran. Production and purification of polyclonal antibody against shrimp lectin LvLTLC1. International Aquatic Research, 13(4): 233–240, December 2021. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_685618.html;   https://intelaquares.tonekabon.iau.ir/
article_685618_47f6de6850ee0be7935dbc7b16e97f00.pdf.

Rodriguez-Forero:2021:VPC

[333] Adriana Rodríguez-Forero, Katrina Medina-Lambraño, and Ernesto Acosta-Ortíz. Variations in the proximate composition of the sea cucumber, Isostichopus sp. aff badionotus. International Aquatic Research, 13 (4):241–252, December 2021. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_686053.html;   https://intelaquares.tonekabon.iau.ir/
article_686053_93f4538f19af87ad8ad23b5214d161c1.pdf.

Do:2021:AEP

[334] Trung Quang Do, Anh The Luu, Thu Thi Tuyet Tran, Tri Trong Nguyen, and Vien Van Dinh. Assessment of enhanced phytoremediation of shrimp aquaculture wastewater by endophytic bacteria-inoculated floating treatment wetlands. International Aquatic Research, 13(4):253–260, December 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_686446.html;
https://intelaquares.tonekabon.iau.ir/article_686446_0048a1522647595bf61a001d2276c2d7.
pdf.

Ventura:2021:TDP

[335] Arlene S. Ventura, Andrea Maria de Araújo Gabriel, Jefferson Rodrigues Gandra, Isabelle Zocolaro Noia, Jayme Aparecido Povh, and Gabriela Tomas Jerônimo. Thermal dynamics and physiological implications in pacu Piaractus mesopotamicus anaesthetised with Ocimum basilicum essential oil. International Aquatic Research, 13(4):261–270, December 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
686478.html;   https://intelaquares.tonekabon.iau.ir/article_
686478_472c42e07101f61e453e48bcf37aab7e.pdf.

Noor:2021:ARG

[336] Mohd Iqbal Mohd Noor, Mohamad Nor Azra, Voon-Ching Lim, Azniza Ahmad Zaini, Fadli Dali, Ina Murni Hashim, Hamnah Che Hamzah, and Muhammad Fuad Abdullah. Aquaculture research in Southeast Asia — a scientometric analysis (1990–2019). International Aquatic Research, 13(4):271–288, December 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_686997.html;   https:/
/intelaquares.tonekabon.iau.ir/article_686997_19a74c65847463422fb5a266061cc3b9.
pdf.

Ugya:2021:EAR

[337] Adamu Yunusa Ugya. The efficiency and antioxidant response of microalgae biofilm in the phycoremediation of wastewater resulting from tannery, textile, and dyeing activities. International Aquatic Research, 13(4):289–300, December 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_687285.html;   https:/
/intelaquares.tonekabon.iau.ir/article_687285_bf7073d4199ee2bcd600860f13db523c.
pdf.

Saldana-Serrano:2021:TFI

[338] Miguel Saldaña-Serrano, Dante Espinoza-Morriberon, Gabriel Enrique De la Torre, Guillermo Odilón Álvarez-Béjar, and Alberto López. Toxicity of fipronil insecticide on the early life stages of Colossoma macropomum (gamitana). International Aquatic Research, 13(4):301–307, December 2021. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL

https://intelaquares.tonekabon.iau.ir/article_687481.html;
https://intelaquares.tonekabon.iau.ir/article_687481_83e8b02371b39a7fa1a679daa8f76ebc.
pdf.

Hossain:2022:EIA

[339] Mohammod Kamruzzaman Hossain, Sairatul Dalian Ishak, Abol Munafi Ambok-Bolong, NoorDiyana Maat Noordin, Shumpei Iehata, and Md. Abdul Kader. Effect of intestinal autochthonous Enterococcus faecalis on the growth performance, gut morphology of Malaysian mahseer (Tor tambroides) and protection against Aeromonas hydrophila. International Aquatic Research, 14(1):1–12, March 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_688224.html;   https:/
/intelaquares.tonekabon.iau.ir/article_688224_baf952d2946c5259fb9bfb8ecfbcb9ae.
pdf.

Nabi:2022:OSS

[340] Shafaquat Nabi, Syed Tanveer, Showkat Ahmad Ganie, and Khalid Bashir. Oxidative stress, serum biochemistry and DNA damage of Cyprinus carpio communis naturally infected with helminths. International Aquatic Research, 14(1):13–22, March 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_688534.html;   https:/
/intelaquares.tonekabon.iau.ir/article_688534_2327742d05ef373697dcefb05117ec14.
pdf.

Izar:2022:CCB

[341] Gabriel Mendes Izar, Giovanna Teixeira Gimiliani, Caio Rodrigues Nobre, Hideshige Takada, Roberto Fioravanti Carelli Fontes, and Denis Moledo de Souza Abessa. Can the colors of beach-stranded plastic pellets in beaches provide additional information for the environmental monitoring? A case study around the Santos Port, Brazil. International Aquatic Research, 14(1):23–40, March 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_689369.html;   https:/
/intelaquares.tonekabon.iau.ir/article_689369_bbe750bc151f521b58350a8d5c7fdcda.
pdf.

Chen:2022:AVP

[342] Chia-Hui Chen, Kazuyuki Sagara, Ryosuke Minamishima, Yuka Yoshino, Daisuke Yamaguchi, Shiro Itoi, and Haruo Sugita. Abundance of Vibrio populations in the gut of Japanese coastal fishes. International Aquatic Research, 14(1):41–50, March 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_689360.html;   https:/
/intelaquares.tonekabon.iau.ir/article_689360_601f75bb604fe74ff9beed090d79caaa.
pdf.

Pakingking:2022:PCA

[343] Rolando Pakingking, Jr., Roselyn Usero, Evelyn Grace de Jesus-Ayson, Dan Joseph Logronio, and Christopher Marlowe Caipang. Phytochemical composition, antioxidant, and antibacterial activity of the Philippine marine green alga (Ulva pertusa). International Aquatic Research, 14(1):51– 62, March 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
689757.html;   https://intelaquares.tonekabon.iau.ir/article_
689757_e8bb4bc10dff9d5f176441ee7959687c.pdf.

Hussein:2022:FMU

[344] Mamdouh Hussein, Ayman Zaki, Neveen Abdel-Raouf, Khawla Alsamhary, Wael Fathy, Mohamed Abdelhameed, and Khaled Elsayed. Flocculation of microalgae using calcium oxide nanoparticles; process optimization and characterization. International Aquatic Research, 14(1):63– 70, March 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
689976.html;   https://intelaquares.tonekabon.iau.ir/article_
689976_0954afbee2463e1ea05475a5fc425590.pdf.

Bavali:2022:FRE

[345] Shima Bavali, Mahsa Haghi, Mohammad Zakeri, and Preeta Kochanian. Feeding and reproduction ecological patterns of Coptodon zillii in Shadegan wetland: an aggressive potential species for local aquaculture. International Aquatic Research, 14(1):71–79, March 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_690166.html;   https:/
/intelaquares.tonekabon.iau.ir/article_690166_e0953031ec132af2e910890af292e159.
pdf.

Fathy:2022:OKE

[346] Wael A. Fathy, Ehab Essawy, Eman Tawfik, Rania Omar, Mohamed S. Abdelhameed, Ola Hammouda, Neveen Abdel-Raouf, and Khaled N. M. Elsayed. Overexpressing key enzymes in lipogenesis to boost microalgae cellular oil content for biofuel production: a mini-review. International Aquatic Research, 14(2):81–94, June 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_691885.html;   https:/
/intelaquares.tonekabon.iau.ir/article_691885_34be5a1bad2c9218b1c0b9cae60c225b.
pdf.

Thome:2022:SAA

[347] Patricia E. Thomé and Jacqueline Rivera-Ortega. Specific antibacterial activity against potential pathogens and restraining of larvae settlement from a pigmented Pseudoalteromonas strain isolated from the jellyfish Cassiopea xamachana. International Aquatic Research, 14(2):95–105, June 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
690387.html;   https://intelaquares.tonekabon.iau.ir/article_
690387_f0bc3cf36f56522242276152f799c1da.pdf.

daSilva:2022:ESD

[348] Weverson Ailton da Silva, Jorge Luiz da Silva, Carlos Yure B. Oliveira, Ana Paula Mariane de Moraes, Renata Akemi Shinozaki-Mendes, and Ugo Lima Silva. Effect of stocking density on water quality, plankton community structure, and growth performance of Litopenaeus vannamei (Boone, 1931) post-larvae cultured in low-salinity biofloc system. International Aquatic Research, 14(2):107–116, June 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_690569.html;   https:/
/intelaquares.tonekabon.iau.ir/article_690569_7b067dadfa6a4c3785926a232fdaa205.
pdf.

Saygili:2022:DMW

[349] Eyup Ilker Saygili, Mehmet Naz, Emine Sukran Okudan, Zafer Cetin, Necla Benlier, Eda Ogut, Meltem Gungor, Seda Busra Bakir, P. nar Gunel Karadeniz, Salih Veziroglu, Ayd n Gulses, Tolga Depci, Oral Cenk Aktas, and Selin Sayin. The determination of the molecular weight profiles and biochemical compositions eight macroalgae species from Turkey. International Aquatic Research, 14(2):117–125, June 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_691409.html;   https:/
/intelaquares.tonekabon.iau.ir/article_691409_2bcb1017f02308ca374f5078cbdbf6a9.
pdf.

Erbil:2022:EFC

[350] Gokhun Cagatay Erbil, Mahmut Elp, and Ya ar Durmaz. Effect of ferric chloride (FeCl_{3}) concentration on pigment production of Porphyridium cruentum. International Aquatic Research, 14(2):127–137, June 2022. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_691568.html;   https:/
/intelaquares.tonekabon.iau.ir/article_691568_4c5cd75e5a99e8ecd0a1cdeab57c4736.
pdf.

doValeFigueiredo:2022:SCG

[351] Julianna Paula do Vale Figueiredo, Ana Paula Mariane De Morais, Weverson Ailton da Silva, Ramires Eloise Queiroz Rafael, Ivanilson De lima Santos, Mauricio Laterça Martins, Felipe do Nascimento Vieira, Frank Belettini, and Edemar Roberto Andreatta. Salinity in the culture of Nile tilapia in a biofloc system: Influence on growth and hematological parameters. International Aquatic Research, 14(2):139–146, June 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_691891.html;   https:/

/intelaquares.tonekabon.iau.ir/article_691891_6a72a948f0a186414534f388d52fdb08.
pdf.

Sagara:2022:DVP

[352] Kazuyuki Sagara, Ryosuke Minamishima, Yusuke Takahashi, Shiro Itoi, and Haruo Sugita. Dynamics of Vibrio populations in the rearing water of tiger puffer (Takifugu rubripes). International Aquatic Research, 14 (2):147–156, June 2022. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_692201.html;   https://intelaquares.tonekabon.iau.ir/
article_692201_ee3db704a54f824e0cbf20b1e25df5f3.pdf.

Haoujar:2022:NSS

[353] Imane Haoujar, Mohammed Haoujar, Ammar B. Altemimi, Adil Essafi, and Francesco Cacciola. Nutritional, sustainable source of aqua feed and food from microalgae: a mini review. International Aquatic Research, 14(3):157–167, September 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_693834.html;   https:/
/intelaquares.tonekabon.iau.ir/article_693834_0de15f72ba58d831b557eeb16c7cddaa.
pdf.

Gyamfi:2022:WQD

[354] Stephen Gyamfi, Regina Esi Edziyie, Kwasi Adu Obirikorang, Daniel Adjei-Boateng, and Peter Vilhelm Skov. Water quality dynamics in earthen ponds with and without fish. International Aquatic Research, 14(3):169–180, September 2022. CODEN ???? ISSN 2008-4935 (print),

2008-6970 (electronic). URL https://intelaquares.tonekabon.iau.
ir/article_693289.html;   https://intelaquares.tonekabon.iau.
ir/article_693289_e8366646164c509610378964a61ad086.pdf.

Victoriano:2022:SPP

[355] Precious Victoriano, Emelyn Joy Gallego Mameloco, Rowena Cadiz, and Rex Ferdinand M. Traifalgar. Sargassum polycystum polysaccharide extract improved immunological responses and enhanced resistance of Penaeus monodon against Vibrio harveyi infection. International Aquatic Research, 14(3):181–192, September 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_695341.html;   https:/
/intelaquares.tonekabon.iau.ir/article_695341_be9b1dac38e123d8681c2d19bd6c3ea7.
pdf.

Azani:2022:DDE

[356] Nizalmie Azani, Abol Munafi Ambok Bolong, Hon Jung Liew, Salleh Kamarudin, Aziz Arshad, Md Mahbubul Hassan, Abu Hena Mustafa Kamal, and Nadiah W. Rasdi. Different dietary effects on growth and reproduction of freshwater zooplankton Ceriodaphnia cornuta (Sars, 1885) and its potential use in Pangasius nasutus larval rearing. International Aquatic Research, 14(3):193–202, September 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_693253.html;   https:/
/intelaquares.tonekabon.iau.ir/article_693253_279261a5e6fa2e0cd10736bc750e08b1.
pdf.

Kuragodage:2022:ELC

[357] Ayesha U. Kuragodage, Palawinnege R. T. Cumaranatunga, and Kariyawasam H. M. A. Deepananda. Effect of light colour on Gracilaria edulis growth and agar. International Aquatic Research, 14(3):203–212, September 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
694414.html;   https://intelaquares.tonekabon.iau.ir/article_
694414_4a20db97682433594a787ba130af0796.pdf.

Fatahi:2022:DCE

[358] Abdolreza Fatahi, Hamid Faghani Langarudi, Majid Mohammadnejad, and hamed mousavi sabet. Dietary cinnamon effects on growth performance, hematological parameters, blood biochemical and immunological indices of rainbow trout (Oncorhynchus mykiss) fingerlings. International Aquatic Research, 14(3):213–221, September 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_695340.html;   https:/
/intelaquares.tonekabon.iau.ir/article_695340_3ad56f837c340a85af0c99dbf96b71a4.
pdf.

Mirmazloomi:2022:MMC

[359] Sina Mirmazloomi, Maryam Ghiasi, Seyedeh Ainaz Shirangi, Saltanat Najjar Lashgari, and Ali Reza Khosravi. Morphological and molecular characterization of two species of Saprolegnia isolated from a rainbow trout (Oncorhynchus mykiss) hatchery in Iran. International Aquatic Research, 14(3):223–230, September 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_695134.html;   https:/
/intelaquares.tonekabon.iau.ir/article_695134_4164f313c77e76157faa6556b78ab55b.
pdf.

Marcuello:2022:PFO

[360] Carlos Marcuello. Present and future opportunities in the use of atomic force microscopy to address the physico-chemical properties of aquatic ecosystems at the nanoscale level. International Aquatic Research, 14(4):231–240, December 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_697391.html;   https:/
/intelaquares.tonekabon.iau.ir/article_697391_962877a743a4a5dba304818072c2e9aa.
pdf.

Hatano:2022:DMS

[361] Kaito Hatano, Akihiro Sakatoku, Daisuke Tanaka, Shoki Tanaka, Tadashi Isshiki, and Nobuo Suzuki. Development of a method of sensitively and specifically detecting a Vibrio sp. strain MA3 associated with mass mortalities of the pearl oyster Pinctada fucata martensii using quantitative PCR. International Aquatic Research, 14(4):241–250, December 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_697209.html;   https:/
/intelaquares.tonekabon.iau.ir/article_697209_a9743dad123fda2cbb3136dd2d79af9a.
pdf.

GoncalvesJunior:2022:UAS

[362] Genes Fernando Gonçalves Junior, Rudã Fernandes Brandão Santos, Carlos Yure Barbosa Oliveira, Ângelo Raphael Alexandre da Silva, Elizabeth Pereira Santos, Ranilson de Souza Bezerra, and Alfredo Olivera Gálvez. The use of Artemia sp. conserved on larval performance of the Pacific white shrimp Penaeus vannamei. International Aquatic Research, 14(4):251–261, December 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_697831.html;   https:/
/intelaquares.tonekabon.iau.ir/article_697831_5ef1edef824a74f887a89aa953591a81.
pdf.

Amir:2022:EIW

[363] Fachrurozi Amir, Zainal A. Muchlisin, Firman M. Nur, Nur Fadli, Moh N. Siti-Azizah, Martin Wilkes, Usman Tang, Bustari Hasan, Agung S. Batubara, Filiz K. Kocabas, and Kasi Marimuthu. Effect of increasing water temperature on the physiology and gill histology of Barramundi, Lates calcarifer (Pisces, Perciformes) fingerlings. International Aquatic Research, 14(4):263–273, December 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_698144.html;   https:/

/intelaquares.tonekabon.iau.ir/article_698144_2b951d725f1bf1bb91ba514c69aa3258.
pdf.

Vo:2022:IOS

[364] Minh Thi Tuyet Vo and Cheng-Hui Lin. Inhibition of orange-spotted grouper nervous necrosis virus replication by short hairpin RNAs in grouper GF-1 cells. International Aquatic Research, 14(4):275–283, December 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (elec-

tronic). URL https://intelaquares.tonekabon.iau.ir/article_
698245.html;   https://intelaquares.tonekabon.iau.ir/article_
698245_2faf8e9efd39bf04efaf22d42a3bf7d5.pdf.

Celis:2022:TCE

[365] José E. Celis, Winfred Espejo, Karen Kidd, Daryl McGoldrick, Mandi Clark, Daiki Kitamura, and Shosaku Kashiwada. Technology-critical elements in fishes from two of the Laurentian Great Lakes. International Aquatic Research, 14(4):285–292, December 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_696849.html;   https:/
/intelaquares.tonekabon.iau.ir/article_696849_6aa7986c5f1a0060794c7ee45d7cabe4.
pdf.

Vital:2022:CRD

[366] Zoltán Vitál, Dorottya Lilla Fazekas, Béla Halasi-Kovács, Zsolt Udvari, Zoltán Ugrai, Flórián Tóth, and Attila Mozsar. Capture rate and distribution patterns of newly-stocked common carps (Cyprinus carpio) in a put and take lotic fishery: a multi-year case study. International Aquatic Research, 14(4):293–302, December 2022. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_697225.html;   https:/
/intelaquares.tonekabon.iau.ir/article_697225_77a7e33114e6140013ce2bf9896ab8b9.
pdf.

Nugraha:2023:RAS

[367] Muhammad Ar Rozzaaq Nugraha, Novi Rosmala Dewi, Muhammad Awaluddin, Ari Widodo, Md Afsar Ahmed Sumon, Mamdoh T. Jamal, and Muhammad Browijoyo Santanumurti. Recirculating Aquaculture System (RAS) towards emerging whiteleg shrimp (Penaeus vannamei) aquaculture. International Aquatic Research, 15(1):1–14, March 2023. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_699267.html;   https:/
/intelaquares.tonekabon.iau.ir/article_699267_e8f59776d7f41fa1fa375f2bd6cb4947.
pdf.

Kuroda:2023:PIC

[368] Kohei Kuroda, Kaito Hatano, Ryoya Kawamura, Ayaka Fukushima, Yuichi Sasayama, Yoshiaki Tabuchi, Yukihiro Furusawa, Mika Ikegame, Atsuhiko Hattori, Jun Hirayama, Tatsuya Fukuda, Shinji Uekura, Hajime Matsubara, Umi Kawago, Toshio Sekiguchi, Ajai Kumar Srivastav, and Nobuo Suzuki. Possible involvement of Calcitonin I and II in calcium metabolism of the female reproductive physiology of goldfish (Carassius auratus). International Aquatic Research, 15(1):15–26, March 2023. CO-DEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_698985.html;   https:/
/intelaquares.tonekabon.iau.ir/article_698985_0ed4df4bb89ab71542c5dbdfbe13d968.
pdf.

Sanpapao:2023:EPS

[369] Pathipat Sanpapao, Wittaya Tawong, Piyawat Pongpadung, Pattareeya Ponza, and Supat Ponza. Effect of phosphorus and sodium chloride levels on growth performance, carotenoid accumulation and isomerization to 9-cis-carotene in Thai Dunaliella salina NUAC09. International Aquatic Research, 15(1):27–38, March 2023. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_699253.html;   https:/
/intelaquares.tonekabon.iau.ir/article_699253_28e6d2e3c19e8a14b2f1b74030018a22.
pdf.

Kocaba:2023:MPS

[370] Filiz Kutluyer Kocaba, Mehmet Kocaba, Aykut Çanakçi, and A. Hasan Karabacak. Mechanical property and structural-elemental analysis of marine bivalve mollusc shells: Cerastoderma edule, Chamelea gallina, Donax trunculus, Ruditapes decussatus. International Aquatic Research, 15(1):39–50, March 2023. CODEN ???? ISSN 2008-4935 (print), 2008-

6970 (electronic). URL https://intelaquares.tonekabon.iau.ir/
article_700058.html;   https://intelaquares.tonekabon.iau.ir/
article_700058_d49b29e3afc9e2ae90c5d468bbebd886.pdf.

Kalteh:2023:EPF

[371] Abdolvahab Kalteh, Afshin Ghelichi, Sara Jorjani, Reza Akrami, and Fariborz Ghojoghi. The effect of partial and full fish oil replacement with poultry fat on growth performance, hematological and serum biochemical parameters, antioxidant capacity, and intestine histology of rainbow trout (Oncorhynchus mykiss). International Aquatic Research, 15(1):51–64, March 2023. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_700618.html;   https:/
/intelaquares.tonekabon.iau.ir/article_700618_88ef3b60ab89eff09c05b393e628be71.
pdf.

Kishore:2023:CGS

[372] Thomas D. Kishore and Alvin Lal. Coastal groundwater salinization due to saltwater intrusion — new insights from a preliminary in-situ assessment. International Aquatic Research, 15(1):65–74, March 2023. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_700635.html;   https:/
/intelaquares.tonekabon.iau.ir/article_700635_9216725c61502e90e46314f24f0e4020.
pdf.

Kordrostami:2023:IGI

[373] Shahin Kordrostami, Masoud Hedayatifard, Mojtaba Keshavarz, Seyed Rohollah Javadian, and Kayvan Hazaei. Improvement of growth, immunological and hematological parameters of juvenile common carp (Cyprinus carpio) treated with zinc sulfate and Vitamin E nanoparticles. International Aquatic Research, 15(1):75–83, March 2023. CODEN ???? ISSN 2008-4935 (print), 2008-6970 (electronic). URL https://

intelaquares.tonekabon.iau.ir/article_700636.html;   https:/
/intelaquares.tonekabon.iau.ir/article_700636_96d1a6734f7b93c39c89f5dcb20edc7d.
pdf.